### v1

# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

Abstract.
A linkage $(G,\ell)$ is a pair containing a finite simple undirected graph $G$ and a squared edge-length map $\ell$ that assigns squared Euclidean lengths to the edges of $G$. A $d$-realization of $(G,\ell)$ is an assignment of points in $\mathbb{R}^{d}$ to the vertices of $G$ for which pair-wise distances between points agree with $\ell$. For $d\leq 3$, we characterize pairs $(G,f)$, where $f$ is a nonedge of $G$, such that, for any squared edge-length map $\ell$, there is a single interval of attained distance values between the endpoints of $f$ over all $d$-realizations of $(G,\ell)$, answering a question posed in [22] a decade ago, which gave an equivalent characterization for $d\leq 2$ that does not generalize to $d\geq 3$. Two notable byproducts of this characterization are a new tool for partial $3$-tree completion, a well-studied problem, and a deeper understanding of the realization space of partial $3$-tree linkages through the lens of Cayley realization spaces. Although related to the minor closed class of $3$-flattenable graphs, the class of pairs $(G,f)$ with the above property is not minor closed, has no obvious well quasi-ordering, and there are infinitely many minimal graphs-nonedge pairs - w.r.t. edge contractions - in the complement class. Our characterization overcomes these obstacles, is based on the forbidden minors of the class of $3$-flattenable graphs, and contributes to the theory of Cayley configurations used for analyzing distance-constrained configuration spaces in a range of application domains. Generalizations to higher dimensions and efficient algorithmic characterizations are conjectured.
Key words and phrases:
Configuration spaces, Graph flattenability, Cayley configuration spaces, Forbidden minor characterization, Distance geometry, Convex geometry
1CISE Department, University of Florida, partially supported by NSF DMS 1564480, NSF DMS 1563234, and DARPA HR00111720031.
2Affiliate in Department of Mathematics
1. Introduction
In this paper, all sets are closed and compact, all maps are closed, and connected implies path-connected unless otherwise specified. A linkage $(G,\ell)$ is a pair containing a (finite simple undirected) graph $G$ and a squared edge-length map $\ell$ that sends each edge in the edge set $E(G)$ to a nonnegative real number. For any integer $d\geq 0$, a (Euclidean) $d$-realization $p$ of $(G,\ell)$ is a map that sends each vertex in the vertex set $V(G)$ to a point in $\mathbb{R}^{d}$ such that $\ell(uv)=\|p(u)-p(v)\|^{2}$ for each edge $uv\in E(G)$. We refer to $p$ as a realization of $(G,\ell)$ when $d$ is clear from context. See Figures 1 and 2 for examples. The configuration space (CS) $\mathcal{C}^{d}(G,\ell)$ is the set of all $d$-realizations of $(G,\ell)$. Traversing and sampling $\mathcal{C}^{d}(G,\ell)$ are difficult and well-studied problems [10, 11, 14, 21, 23, 26, 30] with applications in many domains [16, 18, 29].
Figure 1. (A) A linkage $(G,\ell)$, where $u_{1}u_{2}$ is a nonedge, and (B) a $2$-realization of $(G,\ell)$, where $p_{i}=p(v_{i})$. (A), (C), and (D) show graph-nonedge pairs. See the discussion below.
Figure 2. Left: linkages $(G,\ell)$ and sets $F$ of nonedges (dashed line-segments). Top-right: the CCS $\Omega^{3}_{F}(G,\ell)$ for the top-left linkage; credit to [27]. Bottom-right: the CCS $\Omega^{2}_{F}(G,\ell)$ for the bottom-left linkage. Credit for both bottom figures to [22].
The following concepts were introduced by Sitharam and Gao [22] to study these problems. Letting $F=\{u_{1}v_{1},\dots,u_{m}v_{m}\}$ be an indexed set of nonedges - i.e., pairs of distinct vertices not in $E(G)$ of $G$, the Cayley map $\phi_{F}$ sends a realization $p$ in $\mathcal{C}^{d}(G,\ell)$ to the vector $(\|p(u_{1})-p(v_{1})\|^{2},\dots,\|p(u_{m})-p(v_{m})\|^{2})$ of squared lengths attained by $F$ under $p$. The Cayley configuration space (CCS) $\Omega^{d}_{F}(G,\ell)$ is the image of $\mathcal{C}^{d}(G,\ell)$ under $\phi_{F}$. Note that the preimage of any point in $\Omega^{d}_{F}(G,\ell)$ under $\phi_{F}$ is a non-empty subset of $\mathcal{C}^{d}(G,\ell)$. See Figure 2 for examples. We call $(G,F)$ a graph-nonedge-set pair. If $F$ contains exactly one nonedge $f$, then we call $(G,f)$ a graph-nonedge pair. The pair $(G,F)$ is $d$-convex (resp. $d$-Cayley-connected) if $\Omega^{d}_{F}(G,\ell)$ is convex (resp. connected) for any squared edge-length map $\ell$. Observe that $(G,F)$ is $d$-convex and $d$-Cayley-connected if $d=0$. Hence, we assume that $d\geq 1$ unless stated otherwise. We study the following property.
Definition 1 ($d$-SIP).
Given an integer $d\geq 0$, a graph-nonedge pair $(G,f)$ has the $d$-single interval property (SIP) if it is $d$-convex (equivalently, $d$-Cayley-connected).
1.0.1. Example
Refer to the graph-nonedge pairs $(G,f)$ in Figure 1. For the linkage $(G,\ell)$ in Figure 1(a), observe that the CS $\mathcal{C}^{1}(G,\ell)$ contains exactly two realizations up to translation and overall reflection, and the CCS $\Omega^{1}_{f}(G,\ell)$ is the set $\{1,9\}$. Hence, $(G,f)$ does not have the $1$-SIP. On the other hand, for any $\ell$ and integer $d\geq 2$, the triangle-inequality shows that the CCS $\Omega^{d}_{f}(G,\ell)$ is a single interval, and hence $(G,f)$ has the $d$-SIP. The $2$-realization in Figure 1(b) provides intuition for why $\Omega^{2}_{f}(G,\ell)$ is a single interval. For similar reasons, in Figure 1(c), $(G,f)$ has the $d$-SIP for any $d\geq 3$ but not for any $d<3$ and, in Figure 1(d), $(G,f)$ has the $d$-SIP for any $d\geq 4$ but not for any $d<4$.
Next, refer to the graph-nonedge-sets $(G,F)$ in Figure 2. For the linkage $(G,\ell)$ in the bottom-left image, the triangle-inequality can be used to compute the boundaries of the CCS $\Omega^{2}_{F}(G,\ell)$ shown in the bottom-right image, which is convex. The convexity of this CCS depends only on the fact that it is the intersection of half-spaces, and not on the precise map $\ell$, and so $(G,F)$ is $d$-convex for any $d\geq 2$. By choosing $\ell$ so that it assigns one edge 00 and the rest $1$, it is easy to see that $\Omega^{1}_{f}(G,\ell)$ has cardinality two, and so $(G,F)$ is not $1$-convex. In the top-left image, similar argument using the tetrahedral-inequality shows that $(G,F)$ is $d$-convex for any $d\geq 3$ but not for any $d<3$. The top-right image shows the convex CCS $\Omega^{3}_{F}(G,\ell)$ for the linkage $(G,\ell)$ in the top-left image. It is easy to see that the CS $\mathcal{C}^{2}(G,\ell)$ is connected for any $\ell$, and so $\Omega^{3}_{F}(G,\ell)$ is connected for any $\ell$ since $\phi_{F}$ is continuous. Therefore, $(G,F)$ is $2$-Cayley-connected. However, an argument similar to those above shows that $(G,F)$ is not $1$-Cayley-connected.
Figure 3. Three $d$-convex graph-nonedge-set pairs (nonedges are shown as dashed line-segments) for any $d\geq 3$. The left and middle graphs together with their non-edge-sets are the graphs $C_{5}\times C_{2}$ and $V_{8}$, respectively. See Examples 1.0.1 and 2.0.1 and Section 1.2.
Finally, the graph-nonedge-set pairs $(G,F)$ in Figure 3 are $d$-convex for any $d\geq 3$ but not for any $d<3$. The former fact can be verified for the first two pairs using Theorem 3 in Section 1.2 and for the third pair using Theorem 4 in Section 2 - our main result. The latter fact can be verified using Theorems 1 and 2 in Section 1.2.
1.1. Organization
Previous results on $d$-conxevity, $d$-Cayley-connectedness, and the $d$-SIP are collected in Section 1.2. We state our main results in Section 2. The majority of this paper is spent proving Theorem 4, our main result, which gives a characterization of graph-nonedge pairs that have the $3$-SIP. The primary proof obstacles are detailed in Section 2.1. The proof is given in Sections 3 (converse direction) and 4 (forward direction). Finally, we state several open problems in Section 5, including a conjectured characterization of graph-nonedge pairs that have the $d$-SIP for any $d\geq 1$.
1.2. Related Works
Theorems 1 and 2, below, are characterizations of graph-nonedge-set pairs that are $d$-convex and $d$-SIP for any $d\leq 2$, given by Sitharam and Gao [22] and restated using our terminology. The proof of the former strongly relies on the latter. See Example 1.0.1, in Section 1, and Example 1.2.1, below, for example applications of these theorems. We require the following concepts to state these theorems.
Figure 4. (A) a graph-nonedge pair $(G,f)$ with a clique separator $C$, shown in purple. (B) the $C$-components of $G\cup f$. (C) the atoms of $G\cup f$. See below, Example 2.0.1 in Section 2, and Example 4.0.1 in Section 4.
Given a graph $G$ and a set of its edges or nonedges $F$, $G\cup F$ is the graph obtained from $G$ by adding $F$ to $E(G)$. For any subset $U\subseteq V(G)$, $G\setminus U$ is the graph obtained from $G$ by removing $U$ from $V(G)$. If either $F$ or $U$ contains exactly one element $f$ or $u$, respectively, then we replace $F$ with $f$ and $U$ with $u$ in this notation. A subgraph $H$ of $G$ is induced if, for any $u,v\in V(H)$, $uv\in E(H)$ if and only if $uv\in E(G)$. We say that $H$ is induced by $V(H)$. $U$ is a separator of $G$ if $G\setminus U$ is disconnected, a clique of $G$ if its induces a clique subgraph, and a clique separator (cs) of $G$ if it is both a clique and a separator. An atom is a graph with no cs. An atom of $G$ is a vertex-maximal induced subgraph that has no cs. See Figure 4 for examples. A graph $M$ is a minor of $G$ if it is isomorphic to some graph that can be obtained from $G$ via some sequence of vertex deletions, edge deletions, and edge contractions - i.e., deleting two vertices connected by an edge and adding a new vertex whose neighborhood is the union of the neighborhoods of the deleted vertices.
Theorem 1 ([22] Theorem 5.11).
For any integer $d\leq 2$, a graph-nonedge-set pair $(G,F)$ is $d$-convex if and only if no atom of $G\cup F$ that contains any non-empty subset of $F$ has a $K_{d+2}$ minor.
Theorem 2 ([22] Theorem 5.1).
For any integer $d\leq 2$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has a $K_{d+2}$ minor.
Figure 5. The forbidden minors for $3$-flattenability: $K_{5}$ (left) and $K_{2,2,2}$ (right). See below.
There are several known connections between $d$-convexity and a property of graphs called $d$-flattenability, which was introduced by Belk and Connelly [5, 4]. A graph $G$ is $d$-flattenable if, for any squared edge-length map $\ell$ and any integer $n\geq d$, the CS $\mathcal{C}^{d}(G,\ell)$ is non-empty whenever the CS $\mathcal{C}^{n}(G,\ell)$ is non-empty. Intuitively, this means that any high-dimensional realization of any linkage $(G,\ell)$ can be flattened to live in $d$-dimensions while preserving squared edge-lengths. For example, $K_{3}$ is $2$-flattenable but not $1$-flattenable, since choosing a squared edge-length map $\ell$ that assigns $1$ to each edge of $K_{3}$ ensures the CS $\mathcal{C}^{1}(K_{3},\ell)$ is empty while the CS $\mathcal{C}^{2}(K_{3},\ell)$ is not.
A property of graphs is minor closed if the family of graphs that have this property is closed under edge deletions and contractions. It is easy to see that $d$-flattenability is minor closed. Alternatively, if a graph is $d$-flattenable, then so are all of its minors. Hence, Robertson and Seymour’s famous Graph Minor Theorem [17] states that there exists a finite set of forbidden minors such that a graph is $d$-flattenable if and only if it has no minor in this set. Belk and Connelly [5, 4] showed that this set contains only $K_{3}$ for $d=2$, $K_{4}$ for $d=3$, and $K_{5}$ and $K_{2,2,2}$ for $d=3$ (see Figure 5). Finite forbidden minor characterizations for $d$-flattenability have also been studied for normed spaces other than Euclidean [20, 9, 8].
With these characterizations, the right-hand side of Theorem 2 can be replaced with “no atom of $G\cup f$ that contains $f$ has a $d$-flattenability forbidden minor”. We will see in Example 2.0.1, in Section 2, that this restatement does not extend to $d=3$. Sitharam and Gao [22] established the following relationship between $d$-convexity and $d$-flattenability for $d=3$, and Sitharam and Willoughby [25] later extended it to any $d\geq 0$ and to other norms. This relationship makes use of previous work on understanding the Euclidean and non-Euclidean distance cones [3, 19]. Given a graph $G$ and a set $F$ of its edges and nonedges, $G\setminus F$ is the graph obtained from $G$ by removing $F$ from $E(G)$. If $F$ contains exactly one element $f$, then we replace $F$ with $f$ in this notation.
Theorem 3 ([22] Theorem 5.15, [25] Theorem 1).
Consider a graph $G$ and an integer $d\geq 0$. For any subset $F\subseteq E(G)$, the pair $(G\setminus F,F)$ is $d$-convex if and only if $G$ is $d$-flattenable.
1.2.1. Example
Figure 6. Graphs with one nonedge shown as a dashed-line segment. The graph-nonedge pair on the left has the $2$-SIP, while the one on the right does not. This illustrates that $d$-convexity is not closed under edge deletions. See below and also Section 2.
We examine applications of Theorems 1 - 3 to the graph-nonedge pairs $(G,f)$ (left) and $(G^{\prime},f)$ (right) in Figure 6. The only atom of $G\cup f$ that contains $f$ is $K_{3}$, which does not have a $K_{4}$ minor. On the other hand, $G^{\prime}\cup f$ is an atom with a $K_{4}$ minor. Hence, Theorem 2 shows that $(G,f)$ has the $2$-SIP while $(G^{\prime},f)$ does not. This theorem also shows that neither of these pairs has the $1$-SIP. These statements can be verified using the techniques in Example 1.0.1.
Next, $G\cup f$ is not $2$-flattenable since it has a $K_{4}$ minor. Therefore, Theorem 3 says that there exists some subset $F\subseteq E(G\cup f)$ such that $((G\cup f)\setminus F,F)$ is not $2$-convex. To verify this, let $f^{\prime}$ be the edge of $G$ deleted to obtain $G^{\prime}$. Since some atom of $G\cup f$ contains $f^{\prime}$ and is $K_{4}$, Theorem 2 shows that $((G\cup f)\setminus f^{\prime},f^{\prime})$ does not have the $2$-SIP.
The following provide motivation to study these properties. If a graph-nonedge-set pair $(G,F)$ is $d$-convex, then the CCS $\Omega^{d}_{F}(G,\ell)$ can be efficiently traversed and sampled using well-known methods. Additionally, if $G$ belongs to certain large and natural classes of graphs, then there are efficiently computable maps from $\Omega^{d}_{F}(G,\ell)$ to $\mathcal{C}^{d}(G,\ell)$ [24]. Furthermore, characterizations of $d$-convexity have led to improvements in molecular and particle assembly modeling and kinematic mechanism analysis and design, respectively, via the opensource software EASAL [15, 16] and CayMos [24, 28], and to efficient algorithms for the core problem of distance constraint graph realization [2].
Finally, some of the concepts in this paper appear to be related to similarly long papers [12, 13] on generically globally-linked pairs in graphs. However, note that our study of $d$-convexity and $d$-Caley-connectedness does not require any genericity conditions.
1.3. Tools for working with graph minors.
Most work with minors is only concerned with proving that a graph either has or does not have a given minor, and the specifics of how this minor is obtained are not important. Hence, minors are often treated as equivalence classes of graphs. However, our results concern minors that are obtained without contracting a given edge. Furthermore, there is a natural minor map, formally defined below, between a graph $G$ and one of its minors $M$, and our proof techniques involve manipulating this map to show either that $M$ can be obtained while ensuring certain vertices of $G$ are mapped to specific vertices of $M$ or that some other graph $M^{\prime}$ is also a minor of $G$. Therefore, we need to introduce notation and terminology for working with minors in this manner, which we do below.
Figure 7. (A) a graph $G$, a subset $X\subseteq E(G)$ of edges contained in the black boxes, and a subset $Y\subseteq E(G)$ of dashed edges. (B) a minor $[G]$ witnessed by contractions $X$ and deletions $Y$. For the following, refer to Section 2.1. $H$ is the induced subgraph of $G$ shown in red and $[H]$ is the subgraph of $[G]$ induced by $[V(H)]$ shown in red. $J$ is the induced subgraph of $[G]$ shown in purple and $J^{-1}$ is the subgraph of $G$ induced by $\bigcup_{v\in V(J)}v^{-1}$ shown in purple. Any edge nonedge of $G$ whose endpoints are both contained in the same set $[u_{i}]^{-1}$ is contracted in $[G]$. The nonedge $u_{1}u_{2}$ is preserved and doubled in $[G]$ since $[u_{1}]\neq[u_{2}]$ and $[u_{1}u_{2}]^{-1}$ contains an edge. The nonedge $u_{2}u_{3}$ is retained in $[G]$ since $[u_{2}]\neq[u_{3}]$ and $[u_{2}u_{3}]^{-1}$ does not contain an edge.
We will only ever work with connected minors of connected graphs. Hence, assume $G$ and $M$ are connected. Then, $M$ can be obtained via some sequence of edge deletions and contractions, and no vertex deletions. From here on, all minors discussed are assumed to be obtained from such a sequence. Consider any sequence $G_{1},\dots,G_{n}$ of graphs such that $G=G_{1}$, $M=G_{n}$, and $G_{i+1}$ is obtained from $G_{i}$ via a single edge deletion or contraction. Also, consider the map from $V(G_{i})\cup E(G_{i})$ to $V(G_{i+1})\cup E(G_{i+1})\cup\{\emptyset\}$ that sends a deleted edge to the empty-set, the endpoints of a contracted edge to the vertex added to $G_{i+1}$ by this contraction, and every other vertex or edge to itself. By composing these maps, we get a map $\pi$ from $V(G)\cup E(G)$ to $V(M)\cup E(M)\cup\{\emptyset\}$, which we refer to as a minor map from $G$ to $M$. An edge $uv\in E(G)$ is contracted under $\pi$ if $\pi(u)=\pi(v)$ and deleted under $\pi$ if $\pi(uv)=\emptyset$. Contractions $X\subseteq E(G)$ and deletions $Y\subseteq E(G)$ witness minor $M$ if there exists a minor map from from $G$ to $M$ under which each edge in $X$ is contracted and each edge in $Y$ is deleted. For example, see to Figure 7. Note that all minors of $G$ witnessed by contractions $X$ and deletions $Y$ are isomorphic.
Next, assume contractions $X$ and deletions $Y$ witness $M$. We define a graph $[G]$ as follows. Let two vertices $u,v\in V(G)$ be equivalent if each edge along some path in $G$ between $u$ and $v$ is contained in $X$. Also, let $[u]^{-1}$ denote the equivalence class containing $u$. The vertex set $V([G])$ is the set of all such equivalence classes. For clarity, if we are referring to $[u]^{-1}$ as a vertex, then we will write it as $[u]$. Additionally, we write the pair $[u][v]$ as $[uv]$ and define the set $[uv]^{-1}$ to be $[u]^{-1}\times[v]^{-1}$ if $[u]\neq[v]$, and $[u]^{-1}$ otherwise. In the latter case, we call $[uv]$ a vertex of $[G]$ and treat it as such. The edge set $E([G])$ contains $[uv]$ if and only if $[u]\neq[v]$ and $[uv]^{-1}$ contains some edge in $E(G)\setminus Y$. Sometimes it will be convenient to refer to a vertex $x$ or an edge or nonedge $f$ of $[G]$ without making use of the bracket notation. The brackets are implied in these instances, and so $x^{-1}$ and $f^{-1}$ denote the sets $[x]^{-1}$ and $[f]^{-1}$, respectively. Note that $[G]$ is a minor of $G$, as it is clearly isomorphic to $M$. We use subscripts to differentiate minors denoted with brackets, e.g., $[G]_{A}$ and $[G]_{B}$ are two minors of $G$.
2. Contributions and proof obstacles
From here on, we shorten “$d$-flattenability forbidden minor” to “$d$-forbidden minor,” and we omit $d$ from this terminology when it is clear from context. Given a minor $[G]$ of a graph $G$ and vertices $u,v\in V(G)$, define $uv$ to be contracted in $[G]$ if $[u]=[v]$ and preserved in $[G]$ otherwise. In the latter case, we call $[G]$ $uv$-preserving.
Theorem 4 (Main: $3$-SIP characterization).
A graph-nonedge pair $(G,f)$ has the $3$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $3$-forbidden minor.
The proof of Theorem 4 is presented in Sections 3 (converse direction) and 4 (forward direction). This theorem appears similar to a finite forbidden minor characterization, but has a few important differences. To discuss these differences, we must first extend the notions of minors and minor closed properties to graph-nonedge-set pairs. Given a graph $G$, a minor $[G]$, and a subset $X$ of vertices, edges, and nonedges of $G$, define the set $[X]=\{[x]:x\in X\}$. Letting $F$ be a set of nonedges of $G$ and $F^{\prime}$ be the set of nonedges in $[F]$, $([G],F^{\prime})$ is the minor pair of $(G,F)$ obtained via the sequence of edge deletions and contractions between $G$ and $[G]$. A property of graph-nonedge-set pairs is minor pair closed if the family of graph-nonedge-set pairs that have this property is closed under edge deletions and contractions.
In Example 1.2.1, it was shown that the family of graph-nonedge pairs with the $2$-SIP is not closed under edge deletions, and hence the $2$-SIP is not minor pair closed. Using a similar example based on $K_{d+2}$, it is easy to see that the $d$-SIP is not minor pair closed for any $d\geq 2$. Therefore, the $d$-SIP on the left-hand side of this theorem is not minor pair closed. On the other hand, the property “no atom of $G\cup f$ has an $f$-preserving $3$-forbidden minor” on the right-hand side of the theorem statement is minor closed. However, when we only consider atoms that contain $f$, this property is no longer closed under edge deletions. The intricacies of these statements and their differences are examined in Section 2.1.
2.0.1. Example
Theorem 4 shows that a graph-nonedge pair $(G,f)$ has the $3$-SIP if it is the pair in Figure 1(a), Figure 1(c), the top-left or bottom-left of Figure 2 such that $f$ is any of the dashed nonedges, Figure 3 such that $f$ is any of the dashed nonedges, Figure 6 left or right, Figure 8(a) or 8(b) such that $f$ is $w_{1}w_{2}$, or Figure 10(b) such that $f$ is any of the dashed nonedges. The same theorem shows that $(G,f)$ has the $3$-SIP if $G\cup f$ is any of the graphs in Figures 9(c), 9(e), 9(h), or 9(j), where $f$ is the top-most green edge. The fact that the pairs in Figure 8(a) or 8(b) have the $3$-SIP demonstrates that Theorem 2 does not extend to $d=3$ even after replacing “$K_{d+2}$ minor” with “$d$-forbidden minor” on the right-hand side. Theorem 4 also shows that $(G,f)$ does not have the $3$-SIP if it is the pair in Figure 1(d) or 4(a) or if $G\cup f$ is any of the graphs in Figures 9(a), 9(b), 9(d), 9(f), 9(g), or 9(i), where $f$ is the green edge.
Corollary 1 ($(d\leq 3)$-SIP characterization).
For any integer $d\leq 3$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbidden minor.
Proof.
For $d=3$, the theorem follows from Theorem 4. For any $d\leq 2$, by Theorem 2, it suffices to show that some atom of $G\cup f$ that contains $f$ has a $d$-forbbiden minor if and only if some atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbbiden minor. The converse direction is immediate and the forward direction follows from Theorem 5.2 in [22]. ∎
In Section 5, we conjecture that Corollary 1 is true for all $d\geq 4$. Corollary 2, below, follows immediately from Corollary 1.
Corollary 2 ($(d\leq 3)$-SIP is preserved in minors pairs of atoms-nonedge pairs).
For any integer $d\leq 3$ and any graph-nonedge pair $(G,f)$ such that $G\cup f$ is an atom, if $(G,f)$ has the $d$-SIP, then so do all of its minor pairs.
As this corollary implies, the obstacle that causes the family $A^{d}$ of graph-nonedge pairs with the $d$-SIP to not be closed under edge deletions vanishes if we consider the subset of $A^{d}$ containing minor pairs of pairs $(G,f)$ such that $G\cup f$ is an atom. As we will see in Lemmas 1 and 2, below, $A^{d}$ is closed under contractions and a restricted type of edge deletion for any $d\geq 0$.
Figure 8. Graphs with $K_{5}$ and $K_{2,2,2}$ minors that can only be reached by contracting the edges labeled $w_{1}w_{2}$. See Proposition 1.
Figure 9. See the above discussion, Examples 2.0.1, 4.0.1, 4.0.2, and 4.0.3, and Sections 2.1 and 4. (A), (B), (D), (F), (G), and (I) show graphs $G\cup f$ where $f$ is the green edge and $(G,f)$ is a minimal graph-nonedge pair (see Definition 3). Dashed green line-segments between vertices are nonedges. The dashed black boxes in (B), (D), (G), and (I) contain expanded $f$-components of top-level expanded $f$-separating cms’s of $G\cup f$, which are the purple and orange subgraphs in (B) and (D), the purple subgraph in (G), and the connected subgraph containing one purple edge and one blue edge in (I). (C), (E), (H), and (J) show minors of the graphs to their left obtained by contracting one purple edge in (C) and (H) and one blue edge in (E) and (J). The purple subgraphs in (C) and (H) and the blue subgraphs in (E) and (J) are $f$-separating cms’s whose $f$-components are the subgraphs in the dashed black boxes. The purple $f$-separating cms’s are in the top-level while the blue $f$-separating cms’s are not.
Figure 10. (A) a $3$-tree $T$. (B) a $3$-connected partial $3$-tree obtained from $T$ by deleting the dashed edges.
Next, Theorems 5 and 6, below, are notable byproducts of Theorem 4. Other notable byproducts of Theorem 4 are general lemmas about the structure of graphs with and without cs’s, which are collected in Appendix E. A $3$-tree is either $K_{4}$ or a graph obtained from a $3$-tree $T$ by adding a single vertex connected to each vertex in a clique in $T$ of size $3$. Equivalently, a $3$-tree is an edge-maximal tree-width $3$ graph. A partial $3$-tree is any subgraph of a $3$-tree. Equivalently, a partial $3$-tree is any graph that avoids having $K_{5}$, $K_{2,2,2}$, $V_{8}$, and $C_{5}\times C_{2}$ as minors [1]; see Figures 3 and 5. See Figure 10 for examples. Considerable effort has gone determining if a given set $F$ of nonedges of a partial $3$-tree $G$ is such that $G\cup F$ is a partial $3$-tree [7]. Theorem 5, which is proved in Appendix D, provides a new tool for this task. See Example 2.0.2, below. We write a countable indexed set $\{X_{1},X_{2},\dots\}$ is written as $\{X_{i}\}$.
Theorem 5 (Byproduct: $3$-connected partial $3$-tree star theorem).
For any $3$-connected partial $3$-tree $G$ and any subset $\{wu_{i}\}$ of its nonedges, if $G\cup wu_{i}$ is a partial $3$-tree for each $wu_{i}$, then $G\cup\{wu_{i}\}$ is a partial $3$-tree.
Proof sketch. First, we show that the theorem is true if $G\cup\{wu_{i}\}$ has no $3$-forbidden minor. Then, if this is not the case, we show that $G\cup wu_{i}$ has a $3$-forbidden minor for some nonedge $wu_{i}$. Since $G\cup wu_{i}$ is a partial $3$-tree, and hence has no $3$-forbidden minor, for each $wu_{i}$, this completes the proof.
The above forbidden minor characterization for partial $3$-trees along with Theorem 3 shows that if a graph-nonedge-set pair $(G,F)$ is such that $G\cup F$ is a partial $3$-tree, then $(G,f)$ is $3$-convex. Theorem 6 shows that if $G\cup F$ is a partial $3$-tree, then $(G,F)$ has an even stronger property, which we now define. $(G,F)$ has the $d$-covering map property if, for any squared edge-length map $\ell$, the image of each connected component of the CS $\mathcal{C}^{d}(G,\ell)$ under the Cayley map $\phi_{F}$ is the entire CCS $\Omega^{d}_{F}(G,\ell)$.
Theorem 6 (Byproduct: partial $3$-tree nonedge-set pairs have the $3$-covering map property).
Any graph-nonedge-set pair $(G,F)$ such that $G\cup F$ is a partial $3$-tree has the $3$-covering map property.
Proof sketch. First, observe that we can assume wlog that $G\cup F$ is a $3$-tree. For any squared edge-length map $\ell$, we show that the preimage of the interior of the CCS $\Omega^{3}_{F}(G,\ell)$ under the Cayley map $\phi_{F}$ can be partitioned into connected sets each of whose image under $\phi_{F}$ is the interior of $\Omega^{3}_{F}(G,\ell)$. Since $\phi_{F}$ is a closed map, the image of the closure of each of these sets under $\phi_{F}$ is $\Omega^{3}_{F}(G,\ell)$. Therefore, since each connected component $X$ of the CS $\mathcal{C}^{3}(G,\ell)$ must contain one of these sets, the image of $X$ under $\phi_{F}$ is $\Omega^{3}_{F}(G,\ell)$.
Next, we sketch how to obtain the above desired partition. Consider the set $A$ of $K_{4}$ subgraphs of $G\cup f$ each of whose points are not all coplanar in some $3$-realization of $(G,\ell)$. Let $\sigma$ be a vector with a coordinate $\sigma_{H}\in\{-1,+1\}$ for each $H\in A$. Consider the subset $\mathcal{C}^{3}_{\sigma}(G,\ell)$ of $\mathcal{C}^{3}(G,\ell)$ of realizations $p$ such that the sign on $\sigma_{H}$ matches the sign on the determinant representing the orientation of the points assigned to $H$ under $p$ for each $H\in A$. $\mathcal{C}^{3}_{\sigma}(G,\ell)$ is one of the connected sets that partition the preimage of the interior of $\Omega^{3}_{F}(G,\ell)$ under $\phi_{F}$, and there is one such set for each case of $\sigma$.
Finally, if $A$ is empty, then we show that $\mathcal{C}^{3}(G,\ell)$ is connected, and so the theorem is immediate. Otherwise, we make the observation that the preimage of each point in the interior of $\Omega^{3}_{F}(G,\ell)$ under $\phi_{F}$ intersects each set $\mathcal{C}^{3}_{\sigma}(G,\ell)$. Thus, showing that $\mathcal{C}^{3}_{\sigma}(G,\ell)$ is connected completes the proof.
2.0.2. Example
Consider the partial $3$-tree $P$ in Figure 10(b) and let $F$ be the set of dashed nonedges. Since $P\cup f$ is a partial $3$-tree for any $f\in F$, Theorem 5 states that $P\cup F$ is a partial $3$-tree. Note that $P\cup F$ is the $3$-tree in Figure 10(a). Consider any squared edge-length map $\ell$ that assigns each edge of $P$ a non-zero value and such that the CS $\mathcal{C}^{3}(P,\ell)$ is non-empty. Also, let $u$ be the left-most vertex in Figure 10(b). For any realization $p$ in this CS, observe that the realization obtained by reflecting $p(u)$ across the plane defined by its neighbors is not in the same connected component of this CS as $p$. Even so, Theorem 6 states that the image of each connected component of this CS is the CCS $\Omega^{3}_{F}(P,\ell)$.
Finally, we present several independently interesting properties of $d$-convexity, $d$-Cayley-connectedness, and the $d$-SIP. In particular, Lemmas 1-3, below, are used throughout the proof of Theorem 4. See Example 2.0.3, below, for example applications of these results. A minor is induced if it can be obtained using only contractions. Note that induced minors and topological minors are not equivalent, i.e., while every topological minor is an induced minor, the converse is not true in general. A property of graph-nonedge-set pairs is induced minor pair closed if the family of graph-nonedge-set pairs with this property is closed under contractions.
Lemma 1 ($d$-Convexity/Cayley-connectedness/SIP is induced minor pair closed).
The $d$-convexity (resp. $d$-Cayley-connectivity) property is induced minor pair closed.
Proof.
Consider any graph-nonedge-set pair $(G,F)$ that is $d$-convex (resp. $d$-Cayley-connected). Also, let $[G]$ be any induced minor and let $F^{\prime}$ be the set of nonedges in $[F]$. Observe that, for any linkage $([G],\ell)$, there exists a linkage $(G,\ell^{\prime})$ such that $\Omega^{d}_{F^{\prime}}([G],\ell)$ is a coordinate projection of $\Omega^{d}_{F}(G,\ell^{\prime})$. Since $\Omega^{d}_{F}(G,\ell^{\prime})$ is convex (resp. connected), by assumption, the lemma follows from the fact that coordinate projections of convex (resp. connected) sets are convex (resp. connected). ∎
Lemma 2, below, is proved in Appendix A.1.
Lemma 2 (Edge deletions that preserve $d$-Cayley-connectivity).
Let $G$ be a graph and $F$ and $F^{\prime}$ be disjoint sets of its nonedges. For any integer $d\geq 0$, if $(G,F)$ and $(G\cup F,F^{\prime})$ are $d$-Cayley-connected, then $(G,F\cup F^{\prime})$ is $d$-Cayley-connected. Consequently, $(G,F^{\prime})$ is $d$-Cayley-connected.
Lastly, Lemma 3, which is proved in Appendix A.2, and its corollary are stated below. Consider a graph $G$, any of its induced subgraphs $H_{1}$ and $H_{2}$, and any subset $U\subseteq V(G)$. $H_{1}\cup U$ and $H_{1}\setminus U$ are the graphs induced by $V(H_{1})\cup U$ and $V(H_{1})\setminus U$, respectively. If $U$ contains exactly one vertex $u$, then we replace $U$ with $u$ in this notation. Letting $\{J_{i}\}$ be the connected components of $G\setminus U$, the $U$-components of $G$ are the graphs $J_{i}\cup U_{i}$, where $U_{i}$ is the minimum cardinality subset of $U$ such that $J_{i}$ is a connected component of $G\setminus U_{i}$. The subgraph $H$ induced by $U$ is endowed with the terminology associated with $U$ e.g., the $H$-components of $G$ are the $U$-components are $G$.
Lemma 3 (Clique-gluing).
Let $(G,f)$ be a graph-nonedge pair, $C$ be a cs of $G\cup f$, and $\{G_{i}\}$ be the set of $C$-components of $G\cup f$ that contain $f$. For any integer $d\geq 0$, $(G,f)$ has the $d$-SIP if and only if $(G_{i}\setminus f,f)$ has the $d$-SIP for each $G_{i}$.
A cs $C$ of a graph $G$ is minimal, in which case we refer to $C$ as a cms of $G$, if for any of its induced proper subgraphs $C^{\prime}$, the set of $C$-components of $G$ is not equal to the set of $C^{\prime}$-components of $G$. Consider any sequence $\mathcal{G}_{0},\dots,\mathcal{G}_{n}$ such that $\mathcal{G}_{0}=\{G\}$, $\mathcal{G}_{n}$ contains only atoms, and $\mathcal{G}_{i+1}=(\mathcal{G}_{i}\setminus\{H\})\cup X$, where $H\in\mathcal{G}_{i}$ has a cms $E$ and $X$ is the set of $E$-components of $H$. It was shown in [6] that all such sequences have the same length and $\mathcal{G}_{n}$ is the set of all atoms of $G$.
Corollary 3 (Atom-gluing).
For any integer $d\geq 0$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if, for each atom $G_{i}$ of $G\cup f$ that contains $f$, $(G_{i}\setminus f,f)$ has the $d$-SIP.
Proof.
Let $\mathcal{G}_{0},\dots,\mathcal{G}_{n}$ be any sequence as defined above such that $\mathcal{G}_{0}=\{G\cup f\}$. Lemma 3 shows that $(G,f)$ has the $d$-SIP if and only if, for each set $\mathcal{G}_{i}$ and each of its graphs $H$ that contain $f$, $(H\setminus f,f)$ has the $3$-SIP. This completes the proof. ∎
2.0.3. Example
Consider the graph-nonedge pairs $(G,f)$ (left) and $(G^{\prime},f)$ (right) in Figure 6, and let $f^{\prime}$ be the edge deleted from $G$ to obtain $G^{\prime}$. We saw in Example 1.2.1 that $(G,f)$ has the $2$-SIP while $(G^{\prime},f)$ does not. Since the endpoints of $f^{\prime}$ form a cs of $G\cup f$ and $G^{\prime}\cup f^{\prime}$, these facts can also be verified using Lemma 3. For any induced minor $[G]$, let $F^{\prime}$ be $\{[f]\}$ if $f$ is preserved in $[G]$, and the empty-set otherwise. Lemma 1 says that $([G],F^{\prime})$ has the $2$-SIP, which can be verified using Theorem 2. The contrapositive of Lemma 2 states that since $(G^{\prime},f)$ does not have the $2$-SIP, either $(G^{\prime}\cup f^{\prime},f)$ or $(G^{\prime},f^{\prime})$ does not have the $2$-SIP. Since some atom of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and is $K_{4}$, Theorem 2 shows that $(G^{\prime},f^{\prime})$ does not have the $2$-SIP, which verifies this statement. Lastly, observe that presence of $f^{\prime}$ separates $f$ from any atom with a $K_{4}$ minor, while the absence of $f^{\prime}$ causes $(G^{\prime},f)$ to inherit the non-$2$-SIP from $(G^{\prime},f^{\prime})$. Corollary 2 states that this situation can only occur if $G\cup f$ is not an atom.
2.1. Proof obstacles
Consider a graph-nonedge pair $(G,f)$. The right-hand side of Theorem 4 appears similar to a finite forbidden minor characterization, but has two key differences. First is that only the atoms of $G\cup f$ that contain $f$ must avoid $3$-forbidden minors, and second is that some such minor must preserve $f$. Using Lemma 3, we can choose $(G,f)$ so that it has the $3$-SIP and some atom of $G\cup f$ that does not contain $f$ has a $3$-forbidden minor. Furthermore, even if some atom of $G\cup f$ that contains $f$ has a $3$-forbidden minor, if no such minor preserves $f$, then $(G,f)$ may still have the $3$-SIP. We exemplify the intricacies of the theorem by examining how the characterization of graph-nonedge pairs with the $d$-SIP changes from $d=1$ to $d=3$. A pair of vertices $uv$ of $G$ is doubled in a minor $[G]$ if it is preserved and $[uv]^{-1}$ contains at least one edge if $uv$ is a nonedge, or at least two edges if $uv$ is an edge. Also, $uv$ is retained in $[G]$, in which case $[G]$ is $uv$-retaining, if it is preserved and not doubled. See Figure 7. For each integer $d\leq 3$, we will note whether edge deletions preserve the $d$-SIP and which of the following statements are equivalent to $(G,f)$ not having the $d$-SIP:
  1. (i)
    Some atom of $G\cup f$ that contains $f$ has a $d$-forbidden minor.
  2. (ii)
    Some atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbidden minor.
  3. (iii)
    Some atom of $G\cup f$ that contains $f$ has an $f$-retaining $d$-forbidden minor.
  4. (iv)
    $G\cup f$ has some $f$-preserving $d$-forbidden minor.
The $d=1$ case is well-behaved. Recall that $K_{3}$ is the only $1$-forbidden minor. Figure 1(a) shows the smallest graph-nonedge pair that does not have the $1$-SIP. This example evidences the following equivalence: $(G,f)$ does not have the $1$-SIP if and only if (v) some cycle in $G$ contains $f$. Using this fact, it is easy to see that the family $A^{d}$ of graph-nonedge pairs that have the $1$-SIP is closed under both edge deletions and contractions. The following cycle of implications shows that (v) is equivalent to (i) - (iv): (v) $\implies$ (i) $\implies$ (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (v). The implication (v) $\implies$ (i) follows easily from Lemma 66 in Appendix E, and the other implications have straightforward proofs.
Once we move to $d=2$, some of the above implications become false. Recall that $K_{4}$ is the only $2$-forbidden minor. Figure 1(c) shows the smallest graph-nonedge pair that does not have the $2$-SIP. Theorem 2 and Theorem 5.2 in [22] show that $(G,f)$ not having the $2$-SIP is equivalent to (i) - (iii). While (i), (ii), or (iii) $\implies$ (iv), the converse is false. For example, see the graph-nonedge pair on the left in Figure 6. Also unlike the $d=1$ case, the above equivalences can be used to show that $A^{d}$ is not closed under edge deletions. See Example 1.2.1.
[…]
To summarize, $(G,f)$ may have the $3$-SIP even though some atom of $G\cup f$ that contains $f$ has a $3$-forbidden minor, so long as $f$ is contracted in any such minor. On the other hand, $(G,f)$ may not have the $3$-SIP even though $f$ cannot be retained in any $3$-forbidden minor. The latter statement results in infinitely many minimal graph-nonedge pairs (Definition 3), with respect to contractions, that do not have the $3$-SIP. This statement is false for any integer $d\leq 2$, and so there are finitely many minimal pairs that do not have the $d$-SIP. Furthermore, the $3$-SIP behaves badly when edges are deleted. For any graph-nonedge pair $(G^{\prime},f^{\prime})$ such that $G^{\prime}$ is obtained from $G$ by deleting some edge, it may be that $(G,f)$ has the $3$-SIP while $(G^{\prime},f^{\prime})$ does not, both $(G,f)$ and $(G^{\prime},f^{\prime})$ have the $3$-SIP, $(G,f)$ does not have the $3$-SIP while $(G^{\prime},f^{\prime})$ does, or both $(G,f)$ and $(G^{\prime},f^{\prime})$ do not have the $3$-SIP.
3. Proof sketch for the converse direction of Theorem 4
From here on, when we refer to a forbidden minor we mean a $3$-forbidden minor. Given a graph-nonedge pair $(G_{\star},f_{\star})$, we show that if no atom of $G_{\star}\cup f_{\star}$ that contains $f_{\star}$ has an $f_{\star}$-preserving forbidden minor, then $(G_{\star},f_{\star})$ has the $3$-SIP. The proof is by induction on the number $n$ of vertices in $G_{\star}$. When $n\leq 5$, Theorem 3 shows that $(G_{\star},f_{\star})$ has the $3$-SIP. In the inductive step, this theorem along with Corollary 3 allow us to assume that $G_{\star}\cup f_{\star}$ is an atom that has a forbidden minor. For the remainder of the argument, refer to Figure 11. Given a graph $G$ and one of its edges $e$, a minor $[G]_{A}$ is an $e$-winged graph minor if it is one of the graphs in Figure 8 such that $[e]$ is $w_{1}w_{2}$. Since atoms are $2$-connected, Proposition 1, stated below and proved in Appendix B.1, shows that $G_{\star}\cup f_{\star}$ has an $f_{\star}$-winged graph minor. See Figure 11(a).
Figure 11. See the discussion below. (A) an $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$. (B) $[G_{\star}\cup f_{\star}]$ has singleton wing tips. (C) pins $\{x,y\}$ and $\{w,z\}$ of $[G_{\star}\cup f_{\star}]$. (D) $[G_{\star}\cup f_{\star}]$ is chained and $(G_{\star},\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected.
Proposition 1 (Graphs with winged graph minors).
If $G$ is a $2$-connected graph and $e$ is one of its edges such that $G$ has a forbidden minor but no $e$-preserving forbidden minor, then $G$ has an $e$-winged graph minor.
A graph-nonedge pair $(G,f)$ is a winged graph pair if $G\cup f$ has no $f$-preserving forbidden minor and has an $f$-winged graph minor. The above discussion demonstrates that $(G_{\star},f_{\star})$ is a winged graph pair. The vertex set of any $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$ is written as an indexed set $\{[a_{i}]\}$, where $[a_{i}]$ is $w_{i}$ in one of the graphs in Figure 8. Letting $f_{\star}=uv$, we assume that $u\in[a_{1}]^{-1}$ and $v\in[a_{2}]^{-1}$. Proposition 2, stated below and proved in Appendix B.2, allows us to assume that $[u]^{-1}=\{u\}$ and $[v]^{-1}=\{v\}$, in which case we say that $[G_{\star}\cup f_{\star}]$ has singleton wing tips. See Figure 11(b).
Proposition 2 (Winged graph minors with singelton wing tips).
Consider a winged graph pair $(G,f)$ such that $G\cup f$ is $2$-connected. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then either $(G,f)$ has the $3$-SIP or some $f$-winged graph minor of $G\cup f$ has singleton wing tips.
For any graph $G$, subset $U\subseteq V(G)$, and vertices $s,t\in V(G)\setminus U$, $U$ is an $st$-separator of $G$ if $s$ and $t$ are not contained in the same connected component of $G\setminus U$. An $st$-separator of $G$ is minimal if none of its proper subsets is an $st$-separator of $G$. A pair of vertices $xy$ of $G_{\star}$ is a pin of $[G_{\star}\cup f_{\star}]$ if $\{x,y\}$ is a minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ such that either $x\in[a_{3}]^{-1}$ and $y\in[a_{4}]^{-1}$ or $x\in[a_{5}]^{-1}$ and $y\in[a_{6}]^{-1}$. Note that the paths in $G_{\star}\setminus[f_{\star}]^{-1}$ implied by the existence of $[G_{\star}\cup f_{\star}]$ show that any minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ that is contained in $V(G_{\star})\setminus\{[u]^{-1},[v]^{-1}\}$ and has size at most two is a pin of $[G_{\star}\cup f_{\star}]$. Proposition 3, stated below and proved in Appendix B.3, shows that $[G_{\star}\cup f_{\star}]$ has pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$. See Figure 11(c).
Proposition 3 (Pins of a winged graph minor).
For any winged graph pair $(G,f)$ and any $f$-winged graph minor $[G\cup f]$ whose vertex set is $\{[a_{i}]\}$, there exist pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$.
Next, Proposition 4, stated below and proved in Appendix B.4, allows us to assume that each pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, in which case we call $[G_{\star}\cup f_{\star}]$ chained. See Figure 11(d).
Proposition 4 (Chained winged-graph minors).
For any winged graph pair $(G,f)$, if the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then there exists a winged graph pair $(G^{\prime},f)$ with a chained $f$-winged graph minor $[G^{\prime}\cup f]$ such that $G$ is a spanning subgraph of $G^{\prime}$ and $(G,f)$ has the $3$-SIP if $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley-connected.
For any pin $xy$ of $[G_{\star}\cup f_{\star}]$, observe that there exist unique $\{x,y\}$-components $H_{u}(xy)$ and $H_{v}(xy)$ of $G_{\star}\setminus[f_{\star}]^{-1}$ that contain $u$ and $v$, respectively. Since every pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, each pin of $[G_{\star}\cup f_{\star}]$ other than $xy$ is contained in either $H_{u}(xy)$ or $H_{v}(xy)$, but not both. Hence, we define the ordered sets of pins $\{x_{i}y_{i}\}$ of $[G_{\star}\cup f_{\star}]$ to be the set of all pins of $[G_{\star}\cup f_{\star}]$ indexed such that $H_{v}(x_{i}y_{i})$ contains $x_{i+1}y_{i+1}$. Note that this set contains at least two pins. Proposition 5, stated below and proved in Appendix B.5, shows that $(G_{\star},\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. See Figure 11(d).
Proposition 5 ($(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected).
Consider a winged graph pair $(G,f)$, with $f=uv$, such that $G\cup f$ is $2$-connected. If $(G,f)$ has a chained $f$-winged graph minor with singleton wing tips and an ordered set of pins $\{x_{i},y_{i}\}$, then either $(G,f)$ has the $3$-SIP or $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected.
We are now ready to prove of the converse direction of Theorem 4.
Proof of the converse direction of Theorem 4.
Let $(G,f)$ be a graph-nonedge pair such that no atom of $G\cup f$ that contains $f$ has an $f$-preserving forbidden minor. We will show that $(G,f)$ has the $3$-SIP by induction of the number $n$ of vertices in $G$. When $n\leq 5$, our assumption implies that $G\cup f$ has no forbidden minor, and so Theorem 3 shows that $(G,f)$ has the $3$-SIP. Next, assume that the converse direction of Theorem 4 is true when $n\leq k$, for any integer $k\geq 5$, and we will prove it when $n=k+1$. If $G\cup f$ has no forbidden minor, then Theorem 3 shows that $(G,f)$ has the $3$-SIP. Otherwise, Corollary 3 allows us to assume wlog that $G\cup f$ is an atom. Since atoms are $2$-connected, Proposition 1 shows that $(G,f)$ is a winged graph pair. Consequently, Propositions 2, 3, and 4 allow us to assume that $(G,f)$ has a chained $f$-winged graph minor $[G\cup f]$ with singleton wing tips. Hence, letting $\{x_{i}y_{i}\}$ be the ordered set of pins of $[G\cup f]$, Proposition 5 shows that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected, where $f=uv$. Therefore, by Lemma 2, it suffices to show that $(G\cup\{ux_{2},uy_{2}\},f)$ has the $3$-SIP.
We complete the proof by showing that $(G\cup\{ux_{2},uy_{2}\},f)$ has the $3$-SIP. Observe that $\{u,x_{2},y_{2}\}$ is a clique $wv$-separator of $G\cup\{f,ux_{2},uy_{2}\}$, for some vertex $w\in\{x_{1},y_{1}\}$. Hence, by Lemma 3, it suffices to consider the $\{u,x_{2},y_{2}\}$-component $H$ of $G\cup\{f,ux_{2},uy_{2}\}$ that contains $f$ and show that $(H\setminus f,f)$ has the $3$-SIP. If $(H\setminus f,f)$ does not have the $3$-SIP, then, since $|V(H)|<|V(G)|$, the inductive hypothesis applies and states that $H$ has an $f$-preserving forbidden minor. However, since both $\{x_{1},y_{1}\}$ and $\{x_{2},y_{2}\}$ are minimal $uv$-separators of $G$, there exist paths in the $\{u,x_{2},y_{2}\}$-component of $G\cup f$ that contains $w$ between $w$ and each vertex in the set $\{u,x_{2},y_{2}\}$ such that each path contains exactly one vertex in $\{u,x_{2},y_{2}\}$. This implies that some minor $[G\cup f]_{A}$ is $H$ such that $[f]_{A}$ is $f$. Consequently, $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus f,f)$ has the $3$-SIP. This completes the proof. ∎
4. Proof sketch for the forward direction of main Theorem 4
Given a graph-nonedge pair $(G_{\star},f_{\star})$, we show that if some atom of $G_{\star}\cup f_{\star}$ that contains $f_{\star}$ has an $f_{\star}$-preserving forbidden minor, then $(G_{\star},f_{\star})$ does not have the $3$-SIP. First, we classify the edges of $G_{\star}$. For any graph-nonedge pair $(G,f)$, we treat any edge of $G$ as belonging to $G\cup f$ and $(G,f)$.
Definition 2 (Types of edges of a graph-nonedge pair).
Given a graph-nonedge pair $(G,f)$, the edges of $G$ can be partitioned into the following four classes. Let $[G\cup f]$ be the minor by contracting any one edge of $G$. Then, the contracted edge is of Type
  1. (1)
    if $[G\cup f]$ has no $[f]$-preserving forbidden minor,
  2. (2)
    if $[G\cup f]$ some an $[f]$-preserving forbidden minor, but no atom of $[G\cup f]$ that contains $[f]$ has such a minor,
  3. (3)
    if some atom of $[G\cup f]$ that contains $[f]$ has an $[f]$-preserving forbidden minor and $f$ is doubled in $[G\cup f]$, or
  4. (4)
    otherwise, and in which case the edge is called reducing.
Definition 3 (Minimal pairs).
A graph-nonedge pair $(G,f)$ is minimal if $G\cup f$ is an atom that has an $f$-preserving forbidden minor and no reducing edge.
4.0.1. Example
A graph-nonedge pair $(G,f)$ is not minimal if it is the pair (i) on the right in Figure 3 since $G\cup f$ has no $f$-preserving forbidden minor, (ii) in Figure 4(a) even though $G\cup f$ has an $f$-preserving forbidden minor, since $G\cup f$ is not an atom, or (iii) on the bottom-right of Figure 4(c) even though $G\cup f$ is an atom that has an $f$-preserving forbidden minor, since the edge whose contraction yields an $f$-retaining $K_{5}$ minor is reducing. $(G,f)$ is minimal if it is any of the pairs in Figure 12 or if $G\cup f$ is any of the graphs in Figures 9(a), 9(b), 9(d), 9(f), 9(g), or 9(i).
Lemmas 1 and 3 allow us to assume wlog that $(G_{\star},f_{\star})$ is minimal. Hence, the forward direction of Theorem 4 follows from Theorem 7, below. A squared edge-length map $\ell$ is a proper non-$d$-SIP map for a graph-nonedge pair $(G,f)$ if $\ell$ maps each edge of $G$ to a positive number and the CCS $\Omega^{3}_{f}(G,\ell)$ contains exactly two distinct positive real numbers. For example, the squared edge-length map in Figure 1(a) is a proper non-$1$-SIP map for the graph-nonedge pair in this figure, as illustrated in Example 1.0.1.
Theorem 7 (Every minimal pair has a proper non-$3$-SIP map).
Every minimal graph-nonedge pair has a proper non-$3$-SIP map, and therefore does not have the $3$-SIP.
We spend the remainder of the section proving Theorem 7. The proof is by induction on the number $n$ of vertices in $G_{\star}$. Since $G_{\star}\cup f_{\star}$ has a forbidden minor, we have $n\geq 5$. When $n=5$, $G_{\star}\cup f_{\star}$ is $K_{5}$, and so Proposition 6, stated below and proved in Appendix C.1, shows that $(G_{\star},f_{\star})$ has a proper non-$3$-SIP map.
Figure 12. All minimal graph-nonedge pairs with no Type (2) edge. See Proposition 6.
Proposition 6 (No Type (2) edge implies non-$3$-SIP map).
Figure 12 shows all minimal graph-nonedge pairs that have no Type (2) edge, each of which has a proper non-$3$-SIP map.
In the inductive step, Proposition 6 allows us to assume wlog that $G_{\star}$ has a Type (2) edge. We now present several definitions that will allow us to exploit the presence of such an edge. Consider a graph $G$, any of its edges $e$, any of its induced subgraphs $H$ and $H^{\prime}$, a minor $[G]_{A}$, and a minor $[H]_{B}$. $H\setminus H^{\prime}$ is the subgraph of $G$ induced by $V(H)\setminus V(H^{\prime})$. $[G]_{A}$ is an extension of $[H]_{B}$ if $|V([G]_{A})|=|V([H]_{B})|$ and, for any two vertices $u,v\in V(H)$, $[u]_{A}=[v]_{A}$ if and only if $[u]_{B}=[v]_{B}$. If $H$ contains $e$, then $e$ being contracted, preserved, or retained in $[H]_{B}$ are well-defined concepts (see Section 2). Otherwise, $e$ is contracted in $[H]_{B}$ if it is contracted in every extension of $[H]_{B}$ and preserved (resp. retained) in $[H]_{B}$ if it is preserved (resp. retained) in at least one extension of $[H]_{B}$. $[G]_{A}$ is $e$-separating if it has an $[e]_{A}$-preserving forbidden minor but none of its atoms that contain $[e]_{A}$ has such a minor. For the following, refer to Figure 7. $[H]_{A}$ is the subgraph of $[G]_{A}$ induced by $[V(H)]$. For any induced subgraph $J$ of $[G]_{A}$, $J^{-1}$ is the subgraph of $G$ induced by $\bigcup_{u\in J}u^{-1}$. Note that $[H]_{A}^{-1}$ is not necessarily $H$.
4.0.2. Example
Let $G\cup f$ be the graph in Figure 9(d), where $f$ is the green edge, and let $[G\cup f]$ be its minor shown in Figure 9(e). Observe that the subgraph $H$ of $G\cup f$ consisting of the bottom five vertices and the vertex incident on the two blue edges has an $f$-preserving $K_{5}$ minor. $[G\cup f]$ is an $f$-separating minor since its $K_{5}$ subgraph is its only atom that has an $f$-preserving forbidden minor. Lastly, letting $J$ be the subgraph of $G\cup f$ induced by the bottom five vertices, note that $[J]^{-1}=H$.
Definition 4 (Expanded edge-separating pairs and cms’s).
Let $G$ be a graph and $e$ be any of its edges. An $e$-separating pair $(C,M)$ of an $e$-separating minor $[G]$ contains an atom $M$ of $[G]$, which has an $[e]$-preserving forbidden minor, and a minimal clique $[ux]$-separator $C$ of $[G]$, for some endpoint $[u]$ of $[e]$ and any vertex $[x]$ of $M\setminus C$. Consider the $C$-components $I$ and $H$ of $[G]$ that contain $M$ and $[e]$, respectively. $C$, $I$, and $H$ are the cms, minor-component, and $e$-component of $(C,M)$, respectively.
Next, the pair $(C^{-1},M^{-1})$ is an expanded $e$-separating pair of $G$, and $C^{-1}$, $I^{-1}$, and $H^{-1}$ are the expanded cms, expanded minor-component, and expanded $e$-component of $(C^{-1},M^{-1})$, respectively. $(C^{-1},M^{-1})$ is a top-level expanded $e$-separating pair of $G$ if no proper subgraph of $H^{-1}$ is the expanded $e$-component of any expanded $e$-separating pair of $G$. $(C,M)$ is a top-level $e$-separating pair of $[G]$ if $(C^{-1},M^{-1})$ is in the top-level of $G$.
Finally, an $e$-separating cms $E$ of $[G]$ is a cms of some $e$-separating pair $(E,M)$ of $[G]$. We treat the minor-component and $e$-component of $(E,M)$ as belonging to $E$. $E^{-1}$ is an expanded $e$-separating cms of $G$, and is in the top-level if $(E^{-1},M^{-1})$ is in the top-level.
4.0.3. Example
Let $G\cup f$ be the graph in Figure 9(b), where $f$ is the green edge, $[G\cup f]$ be the minor shown in Figure 9(c), and $C$ and $M$ be the purple and $K_{5}$ subgraphs of $[G\cup f]$, respectively. Observe that $(C,M)$ is an $f$-separating pair of $[G\cup f]$, and hence $C$ is an $f$-separating cms. The $f$-component $H$ of $C$ is the subgraph in the dashed black box and the minor component is $M$. The subgraph $C^{-1}$ of $G\cup f$ is a top-level expanded $f$-separating cms whose expanded $f$-component and expanded minor component are $H^{-1}$ and $M^{-1}$, respectively. Note that the orange subgraph of $G\cup f$ is also a top-level expanded $f$-separating cms whose expanded $f$-component and expanded minor component are $H^{-1}$ and $M^{-1}$, respectiveely.
Next, the above statements are true if we let $G\cup f$ be the graph in Figure 9(d), where $f$ is the green edge, $[G\cup f]$ be the minor shown in Figure 9(e), and $C$ and $M$ be the blue and $K_{5}$ subgraphs of $[G\cup f]$, respectively, except that $C^{-1}$ is not in the top-level of $G\cup f$. To verify this, note that the purple subgraph $E^{-1}$ of $G\cup f$ is an expanded $f$-separating cms whose expanded $f$-component $J^{-1}$ is the subgraph in the dashed black box, which is a proper subgraph of $H^{-1}$. It is easy to check that $E^{-1}$ is in the top-level of $G\cup f$. Observe that the degree two vertex whose neighbors are the vertices of $C$ is not contained in $H$ or $M$.
Finally, the above statements are true for the pairs of graphs in Figures 9(g) and 9(h) and in Figures 9(i) and 9(j), except, in the latter case, $E^{-1}$ is the connected subgraph consisting of one purple edge and one blue edge.
By definition, contracting any Type (2) edge of $G_{\star}\cup f_{\star}$ yields an $f_{\star}$-separating minor. Hence, Proposition 7, stated below and proved in Appendix C.2, shows that $G_{\star}\cup f_{\star}$ has a top-level expanded $f_{\star}$-separating cms.
Proposition 7 (Graphs with edge-separating minors have an expanded edge-separating cms).
Given a graph $G$ and one of its edges $e$, if $G$ has an $e$-separating minor, then $G$ has a top-level expanded $e$-separating cms.
Next, let $C^{-1}$ be any top-level expanded $f_{\star}$-separating cms of $G_{\star}\cup f_{\star}$, $H^{-1}$ be its expanded $f_{\star}$-component, and $G^{\prime}=G_{\star}\setminus(H^{-1}\setminus C^{-1})$. Using the inductive hypothesis and Proposition 8, stated below and proved in Appendix C.3, it is easy to show that some nonedge $f^{\prime}$ of $C^{-1}$ is such that $(G^{\prime},f^{\prime})$ has a proper non-$3$-SIP map $\ell^{\prime}$.
Proposition 8 (Minimal pair inductive structure).
Let $(G,f)$ be a minimal graph-nonedge pair with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. There exists a nonedge $f^{\prime}=uv$ of $C^{-1}$ and an induced subgraph $J$ of $G^{\prime}$ that contains $f^{\prime}$ such that $(J,f^{\prime})$ is minimal and $G^{\prime}$ can be obtained from $J$ by adding some vertices whose neighborhoods are the set $\{u,v\}$. Furthermore, these properties are true when (i) $f^{\prime}=w_{1}w_{3}$ and $H^{-1}$ is the graph in Figure 15(c), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$, or (ii) $f^{\prime}$ is any nonedge of $C^{-1}$ and $H^{-1}$ is the graph in Figure 15(d), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
The goal is now to use $\ell^{\prime}$ to construct a proper non-$3$-SIP map $\ell$ for $(G_{\star},f_{\star})$. For example, if $G_{\star}\cup f_{\star}$ is the graph in Figure 9(b) (resp. 9(g)), where $f_{\star}$ is the green edge, then we can take $G^{\prime}\cup f^{\prime}$ to be the graph in Figure 9(a) (resp. 9(f)), where $f^{\prime}$ is the green edge. It is not hard to see that $\ell$ can be chosen to agree with $\ell^{\prime}$ on edges of $G^{\prime}$ and to assign squared lengths for the two remaining edges of $G_{\star}$ whose sum and difference are the two values in the CCS $\Omega^{3}_{f^{\prime}}(G^{\prime},\ell^{\prime})$. This ensures that $\ell$ has the desired properties. Similarly, $\ell$ can be extended to obtain a non-$3$-SIP map for the pair in Figure 9(d) (resp. 9(i)). To implement the above strategy in general, Proposition 9, stated below and proved in Appendix C.4, lists all possible graphs for $H^{-1}$. Combining this with Proposition 10, stated below and proved in Appendix C.5, completes the proof by induction.
Proposition 9 (Pruned possible graphs for top-level $H^{-1}$).
Let $H^{-1}$ be the expanded $f$-component of any top-level expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair. Then, $H^{-1}$ is one of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
Proposition 10 (Extending non-$3$-SIP maps for minimal pairs).
Let $G^{\prime}$ be any of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, let $H$ be the subgraph of $G$ induced by the set $\{w_{i}\}$, let $f$ be the red edge, and let $f^{\prime}$ be the nonedge of $H$ that does not share an endpoint with $f$ if possible, or any other nonedge of $H$ otherwise. Also, let $G=(G^{\prime}\setminus f)\cup f^{\prime}$. Lastly, let $\ell_{1}$ and $\ell_{2}$ be squared edge-length maps for $H$ that assign each edge a positive real number and differ only in the numbers assigned to $f^{\prime}$. If the linkages $(H,\ell_{1})$ and $(H,\ell_{2})$ both have at least one $3$-realization, then there exist squared edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for $G$ such that
  1. (i)
    they agree with each other on the edges not in $H$ and assign these edges positive real numbers,
  2. (ii)
    $\ell^{\prime}_{i}$ agrees with $\ell_{i}$ on the edges of $H$, and
  3. (iii)
    the CCSs $\Omega^{3}_{f}(G,\ell^{\prime}_{1})$ and $\Omega^{3}_{f}(G,\ell^{\prime}_{2})$ each contain exactly one positive value, and these values are distinct.
Before proving Theorem 7 and the forward direction of Theorem 4, we sketch the the proofs of Propositions 8 and 9, as they are important and highly technical. Lemmas 4 and 5, stated below and proved in Appendix C.3.1, describe the combinatorial structure of $C^{-1}$ and $H^{-1}$, respectively. For any induced subgraphs $H_{1}$ and $H_{2}$ of $G$, $H_{1}\cup H_{2}$ is the subgraph of $G$ induced by $V(H_{1})\cup V(H_{2})$.
Figure 13. Top-level expanded $f$-separating cms’s $C^{-1}$ of minimal graph-nonedge pairs. See Definition 4 and Lemma 4.
Figure 14. Graphs in Lemma 4. Dashed line-segments may or may not be edges.
Lemma 4 (Unpruned possible graphs for top-level $C^{-1}$).
Let $C^{-1}$ be any top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. Then, the following statements are true:
  1. (1)
    $C^{-1}$ is one of the graphs in Figure 13 and
  2. (2)
    the minor $[H^{-1}\cup I^{-1}]$ obtained by contracting every edge in both $H^{-1}\setminus C^{-1}$ and $I^{-1}\setminus C^{-1}$ is one of the graphs in Figure 14 such that $[C^{-1}]$ is the subgraph induced by the set $\{w_{i}\}$.
Lemma 5 (Structure of top-level $H^{-1}$).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let $H^{-1}$ be its expanded $f$-component. The neighborhood of any vertex $w$ in $H^{-1}\setminus(C^{-1}\cup\{u,v\})$) contains only $u$, $v$, and at least one vertex in $C^{-1}\setminus\{u,v\}$. Furthermore, the vertices $u$ and $v$ may have neighbors in $C^{-1}$.
These two lemmas are used to prove Lemma 6, stated below and proved in Appendix C.3.2, which eliminates the graphs in Figures 13(e) - 13(i) as possibilities for $C^{-1}$.
Lemma 6 (Pruned possible graphs for $C^{-1}$).
If $C^{-1}$ is a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, then $C^{-1}$ is one of the graphs in Figures 13(a)-13(d).
These three lemmas are used to prove Lemma 7, stated below and proved in Appendix C.3.3, which gives the unpruned set of all possible graphs for $H^{-1}$. The proof of this lemma is highly technical and has many cases.
Figure 15. See Lemma 7. All possible expanded $f$-components of an expanded $f$-separating cms that has fewer than five vertices and is in the top-level of a minimal graph-nonedge pair $(G,f)$. Dashed line-segments may be edges or nonedges. Hollow vertices may or may not be present. In 15(h), either $u$ neighbors $w_{1}$ or at least one of the hollow vertices is present. In 15(e), 15(i), or 15(l), either $u$ neighbors $w_{3}$ or at least one of the hollow vertices is present. In 15(k), either $u$ neighbors some vertex in $\{w_{3},w_{4}\}$ or at least one of the hollow vertices is present.
Lemma 7 (Unpruned possible graphs for top-level $H^{-1}$).
If $C^{-1}$ is a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, then the expanded $f$-component of $C^{-1}$ is one of the graphs in Figure 15 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$.
After reducing the possible graphs for $H^{-1}$ to those in Figures 15(a) - 15(m) (Lemma 44 in Appendix C.3.4), proving Proposition 8 is just a matter of case checking (Lemma 47 in Appendix C.3.4). Finally, Proposition 9 is proved by eliminating the graphs in Figures 15(f) - 15(m) as cases for $H^{-1}$ (Lemmas 51 - 54 in Appendix C.4).
Proof of Theorem 7.
We show that a minimal graph-nonedge pair $(G,f)$ has a proper non-$3$-SIP map by induction on the number $n$ of vertices in $G$. Since $G\cup f$ has a forbidden minor, we have $n\geq 5$. When $n=5$, $G\cup f$ is $K_{5}$, and so the theorem follows from Proposition 6. Assume that the theorem is true when $n\leq k$, for any integer $k\geq 5$, and we will show that it is true when $n=k+1$. The same proposition allows us to assume that $(G,f)$ has a Type (2) edge. Hence, $(G,f)$ has a top-level expanded $f$-separating cms $C^{-1}$, by Proposition 7. Let $H^{-1}$ be the expanded $f$-component of $C^{-1}$ and $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Consider any minimal graph-nonedge pair $(J,f^{\prime})$ obtained by applying Proposition 8 to $(G,f)$ and $C^{-1}$. Proposition 9 tells us $H^{-1}$ is one of the graphs in Figures 15(a) - 15(e) such that the dashed line-segment is a nonedge, $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$. Hence, Proposition 8 allows us to choose $f^{\prime}=w_{1}w_{3}$ if $H^{-1}$ is the graph in Figure 15(c), and $f^{\prime}=w_{2}w_{4}$ if $H^{-1}$ is the graph in Figure 15(d).
Next, since $J$ has strictly fewer vertices than $G$, the inductive hypothesis applies and states that $(J,f^{\prime})$ has a proper non-$3$-SIP map. Using the relationship between $G^{\prime}$ and $J$ described in Proposition 8 along with the triangle-inequality, we can obtain a proper non-$3$-SIP map $\ell_{G^{\prime}}$ for $(G^{\prime},f^{\prime})$. By definition, this means that the CCS $\Omega^{3}_{f^{\prime}}(G^{\prime},\ell_{G^{\prime}})$ contains exactly two distinct positive real numbers. Let $\ell_{1}$ and $\ell_{2}$ be squared edge-length maps for the edges of $C^{-1}\cup f^{\prime}$ that agree with $\ell_{G^{\prime}}$ on the edges of $C^{-1}$ and assign different squared lengths from $\Omega^{3}_{f^{\prime}}(G^{\prime},\ell_{G^{\prime}})$ to $f^{\prime}$. Note that the linkages $(C^{-1}\cup f^{\prime},\ell_{1})$ and $(C^{-1}\cup f^{\prime},\ell_{2})$ each have at least one $3$-realization. Therefore, consider the edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for the edges of $(H^{-1}\setminus f)\cup f^{\prime}$ given by Proposition 10. Define $\ell_{G}$ to be the edge-length map for the edges of $G$ that agrees with $\ell_{G^{\prime}}$ on the edges of $G^{\prime}$ and $\ell^{\prime}_{1}$ on the edges of $H^{-1}$. Clearly, $\ell_{G}$ is a proper non-$3$-SIP map for $(G,f)$, which completes the proof by induction. ∎
Proof of the forward direction of Theorem 4.
Let $(G,f)$ be a graph-nonedge pair such that some atom of $G\cup f$ that contains $f$ has an $f$-preserving forbidden minor. Corollary 3 allows us to assume wlog that $G\cup f$ is an atom. If $(G,f)$ has a reducing edge, then let $[G\cup f]$ be the minor obtained by contracting any such edge. By definition, $[G\cup f]$ is $f$-retaining and has an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. Hence, by Lemma 1, if $([G\cup f]\setminus[f],[f])$ does not have the $3$-SIP, then neither does $(G,f)$. Therefore, we can assume wlog that $(G,f)$ has no reducing edge. Combining this with the fact that $G\cup f$ is an atom shows that $(G,f)$ is minimal, and thus the theorem follows from Theorem 7. ∎
5. Open problems and conjectures
Open Problem 1.
Give a polynomial time algorithm to decide the characterization in Theorem 4.
There is likely such an algorithm based on the polynomial time of [17] for finite forbidden minors.
Open Problem 2.
Use the characterization in Theorem 4 to characterize $3$-convex graph-nonege-set pairs, similar to how the characterization of $2$-convex graph-nonege-set pairs in [11] is based on their characterization of graph-nonege pairs with the $2$-SIP.
We believe our characterization in Theorem 4 extends to any dimension as follows.
Conjecture 1.
A graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbidden minor.
Since the number of forbidden minors for $d$-flattenability is conjectured to grow asymptotically with the number of forbidden minors for partial $d$-trees [5], which is at least $75$ for $d=4$, a proof of this conjecture must avoid using properties of the explicit $d$-flattenability forbidden minors. To this end, we posit the following problem.
Open Problem 3.
For any integer $d\geq 0$, give a characterization of the set of $d$-flattenability forbidden minor, other than the set of minors that characterize $d$-flattenable graphs.
Appendix A Proofs of the lemmas in Section 2
A.1. Proof of Lemma 2 [Edge deletions that preserve $d$-Cayley-connectivity]
Lemma 2 is proved using Lemma 8, below. For any integer $n>0$, consider a subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections. For any point $x\in S^{\prime}$, the fiber $S(x)$ of $x$ through $S$ is the preimage of $x$ under the projection map from $S$ to $S^{\prime}$. Recall that in this paper, all sets are closed and compact, all maps are closed, and connected implies path-connected, unless otherwise specified.
Lemma 8 (Connected projections of disconnected sets have disconnected fibers).
For any integer $n>0$, consider a disconnected compact subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections. If $S^{\prime}$ is connected, then there exists a point $x\in S^{\prime}$ whose fiber $S(x)$ is disconnected.
Proof.
Assume to the contrary that $S^{\prime}$ is connected and the fiber of any of its points through $S$ is connected. We can assume wlog that $S$ has exactly two connected components $S_{1}$ and $S_{2}$. Consider any two points $p\in S_{1}$ and $q\in S_{2}$ and let $p^{\prime}$ and $q^{\prime}$ be their projections in $S^{\prime}$. By assumption, there exists a path $P$ in $S^{\prime}$ between $p^{\prime}$ and $q^{\prime}$. Since $S_{1}$ and $S_{2}$ are compact, the fiber $S(x)$ of any point $x$ along $P$ is a subset of either $S_{1}$ or $S_{2}$. Hence, define $C(x)$ to be $1$ if $S(x)\subseteq S_{1}$, and $2$ otherwise. Also, let $Y$ be the set of all points $y$ along $P$ such that $C(y)=C(p^{\prime})$ and let $Z$ be the set of all points $z$ along $P$ such that $C(z)=C(q^{\prime})$. Note that $C(p^{\prime})\neq C(q^{\prime})$, and so $Y\cap Z=\emptyset$. However, this implies that $P$ is disconnected, which is a contradiction. Thus, the lemma is proved. ∎
Remark. Lemma 8 is true if we replace $S$ with any subset of a topological vector space.
Proof of Lemma 2.
Assume to the contrary that $(G,F)$ and $(G\cup F,F^{\prime})$ are $d$-Cayley-connected but $(G,F\cup F^{\prime})$ is not. By definition, there exists a squared edge-length map $\ell$ such that the CCS $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ is disconnected. The CCS $\Omega^{d}_{F}(G,\ell)$ is connected, by assumption. Note that $\Omega^{d}_{F}(G,\ell)$ is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$. Hence, letting $S=\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ and $S^{\prime}=\Omega^{d}_{F}(G,\ell)$, Lemma 8 shows that there exists a point $x\in S^{\prime}$ such that $S(x)$ is disconnected. However, this implies that there exists a squared edge-length map $\ell^{\prime}$ such that the CCS $\Omega^{d}_{F^{\prime}}(G\cup F,\ell^{\prime})$ is disconnected, which contradicts the fact that $(G\cup F,F^{\prime})$ is $d$-Cayley-connected. Therefore, $(G,F\cup F^{\prime})$ is $d$-Cayley-connected.
Finally, we show that if $(G,F\cup F^{\prime})$ is $d$-Cayley-connected, then so is $(G,F^{\prime})$. By assumption, the CCS $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ is connected for any squared edge-length map $\ell$. Since $\Omega^{d}_{F^{\prime}}(G,\ell)$ is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$, it is connected. This completes the proof. ∎
A.2. Proof of Lemma 3 [Clique-gluing]
Lemma 3 is proved using Lemma 9, below. For any two induced subgraphs $H_{1}$ and $H_{2}$ of a graph $G$, $H_{1}\cap H_{2}$ is the subgraph of $G$ induced by $V(H_{1})\cap V(H_{2})$.
Lemma 9 (Intersection of CCSs).
Consider a graph-nonedge pair $(G,f)$, a cs $C$ of $G\cup f$, and the set $\{G_{i}\}$ of all $C$-components of $G\cup f$ that contain $f$. Also, consider any linkage $(G,\ell)$ and let $\ell_{i}$ be the restriction of $\ell$ to the edges of $G_{i}\setminus f$. For any integer $d\geq 0$, the following equation is true:
$\Omega^{d}_{f}(G,\ell)=\bigcap_{i}\Omega^{d}_{f}(G_{i}\setminus f,\ell_{i}).$
Proof.
Let $I=\Omega^{d}_{f}(G,\ell)$ and $I_{i}=\Omega^{d}_{f}(G_{i}\setminus f,\ell_{i})$. First, we show that $I\subseteq\bigcap_{i}I_{i}$. For any point $x\in I$, there exists a $d$-realization of $(G,\ell)$ in which $f$ attains the squared length $x$. From this realization, we get a $d$-realization of each linkage $(G_{i}\setminus f,\ell_{i})$ in which $f$ attains the squared length $x$. Hence, $x$ is contained in $\bigcap_{i=1}^{n}I_{i}$.
Next, we show that $\bigcap_{i}I_{i}\subseteq I$. Let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$ and let $\ell^{\prime}_{i}$ be the restriction of $\ell$ to the edges of $H_{i}\setminus f$. Also, let $y$ be any point in $\bigcap_{i}I_{i}$ and let $p_{i}$ be any $d$-realization of $(H_{i}\setminus f,\ell^{\prime}_{i})\}$ such that if $H_{i}$ contains $f$, then $f$ attains the squared length $y$ in $p_{i}$. For any two $C$-components $H_{i}$ and $H_{j}$, observe that the restrictions of $p_{i}$ and $p_{j}$ to the vertex set of $H_{i}\cap H_{j}$ are equivalent up to translations, rotations, and overall reflection. Hence, we can construct a $d$-realization of $(G,\ell)$ in which $f$ attains the squared length $y$ by gluing all the realizations in $\{p_{i}\}$ together at their common intersections. Thus, $y$ is contained in $I$. ∎
An induced subgraph $H$ of a graph $G$ is preserved in a minor $[G]$ if each pair of its vertices is preserved in $[G]$. $H$ is weakly retained in $[G]$ if it is preserved and each of its nonedges is retained.
Proof of Lemma 3.
For one direction, assume that each pair $(G_{i}\setminus f,f)$ has the $d$-SIP. Consider any linkage $(G,\ell)$ and let $\ell_{i}$ be the restriction of $\ell$ to the edges of $G_{i}\setminus f$. Also, let $I=\Omega^{d}_{f}(G,\ell)$ and $I_{i}=\Omega^{d}_{f}(G_{i}\setminus f,\ell_{i})$. Since each set $I_{i}$ is a single interval, by assumption, Lemma 9 shows that $I$ is a single interval, and so $(G,f)$ has the $d$-SIP.
For the other direction, let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$, $H^{\prime}_{i}=H_{i}\setminus f$, and $C^{\prime}=C\setminus f$. Assume wlog that $H_{1}$ contains $f$ and $(H_{1}\setminus,f)$ does not have the $d$-SIP. We will show that there exists an induced minor $[G]$ such that $([G],[f])$ does not have the $d$-SIP, and then apply the contrapositive of Lemma 1 to complete the proof. Let $[G]_{A}$ be the minor obtained by contracting all edges in $H^{\prime}_{i}\setminus C^{\prime}$ for each $H^{\prime}_{i}$ with $i\neq 1$. If $C^{\prime}$ does not contain $f$, then consider the minor $[G]_{B}$ obtained from $[G]_{A}$ by contracting any edge incident to each vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ with $i\neq 1$. Note that $[f]_{B}$ is a nonedge of $[G]_{B}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{B}$, and $[G]_{B}=[H^{\prime}_{1}]_{B}$. Hence, since $(H^{\prime}_{1},f)$ does not have the $d$-SIP, neither does $([G]_{B},[f]_{B})$, as desired.
Finally, if $C^{\prime}$ contains $f$, then consider the minor $[G]_{B}$ obtained from $[G]_{A}$ as follows. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ adjacent to some vertex that is not an endpoint of $[f]_{A}$, contract the edge between two vertices. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ that has degree $1$, contract the edge incident to it. Note that $[f]_{B}$ is a nonedge in $[G]_{B}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{B}$, and $[G]_{B}$ can be obtained from $[H^{\prime}_{1}]_{B}$ by adding vertices of degree $2$ that neighbor the endpoints of $[f]_{B}$. Using these facts along with the triangle-inequality and the assumption that $(H_{1},f)$ does not have the $d$-SIP, it is easy to see that $([G]_{B},[f]_{B})$ does not have the $d$-SIP, as desired. This completes the proof. ∎
Appendix B Proofs of the converse direction propositions in Section 3
B.1. Proof of Proposition 1 [Graphs with winged graph minors]
The proof of Proposition 1 requires Lemmas 10 and 11, below. The proof of Lemma 10 requires the following operation that transforms one minor $[G]_{A}$ of a graph $G$ into another minor $[G]_{B}$. Refer to Figure 16. Consider an edge $[uv]_{A}$ of $[G]_{A}$, the subgraph $H$ of $G$ induced by the set $[u]_{A}^{-1}$, a vertex $x\in[u]_{A}^{-1}$ with a neighbor in the set $[v]_{A}^{-1}$, and a connected component $J$ of $H\setminus x$. The minor $[G]_{B}$ obtained from $[G]_{A}$ by replacing $[u]_{A}^{-1}$ with $V(J)$ and $[v]_{A}^{-1}$ with $[v]_{A}^{-1}\cup(V(H)\setminus V(J))$ is said to be obtained via the vertex exchange from $[u]_{A}^{-1}$ to $[v]_{A}^{-1}$ that exchanges $x$ and fixes $J$.
Figure 16. A vertex exchange that exchanges $x$ and fixes $J$ is shown in red. A component exchange that exchanges $K$ is shown in purple. See the discussion below and Appendix B.2.
Figure 17. Graphs in Lemma 10.
Lemma 10 (Transforming a minor in Figure 17 into one in Figures 8 or 18).
Let $G$ be a $2$-connected graph and $e$ be one of its edges. If some minor $[G]_{A}$ is the graph in Figure 17(a) (resp. Figure 17(b)) such that $[e]_{A}$ is $w_{1}w_{2}$, then some minor $[G]_{B}$ is one of the graphs in Figures 8(a) (resp. Figure 8(b)) or 18(a)-18(d) (resp. 18(e)-18(i)) such that $[e]_{B}$ is $w_{1}w_{2}$. Furthermore, letting $\{w_{i}\}$ be the vertex sets of $[G]_{A}$ and $[G_{B}]$, we have $[w_{i}]_{A}^{-1}=[w_{i}]_{B}^{-1}$ for any $i\notin\{1,2\}$.
Proof.
Let $e=uv$ and $[G]$ be a minor that is the graph in Figure 17(a) (resp. Figure 17(b)) and such that $[u]$ is $w_{1}$, $[v]$ is $w_{2}$, and $[u]^{-1}$ has minimum size. Let $H$ be the subgraph of $G$ induced by $[u]^{-1}$ and $x$ be any vertex in $[u]^{-1}\setminus\{u\}$ with a neighbor in $[v]^{-1}$, which exists since $G$ is $2$-connected. Also, let $J$ be the connected component of $H\setminus x$ that contains $u$. First, we show how to choose $x$ such that $J$ contains some vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. If $J$ contains such a vertex, then our current choice for $x$ works. Otherwise, since there exists a path in $G\setminus x$ between $u$ and some vertex in $G\setminus([u]^{-1}\cup[v]^{-1})$, some connected component $K$ of $H\setminus x$ contains some vertex $y$ with a neighbor in $[v]^{-1}$ and some vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. Note that the connected component of $H\setminus y$ that contains $u$ contains both $J$ and $x$. Hence, we can set $x$ to be $y$ and repeat this argument until $J$ has the desired property.
Next, consider the minor $[G]_{A}$ obtained via the vertex exchange from $[u]^{-1}$ to $[v]^{-1}$ that exchanges $x$ and fixes $J$. If $x$ does not have a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$, then our above choice for $x$ ensures that $[G]_{A}$ is the graph in Figure 17(a) (resp. Figure 17(b)) such that $[u]_{A}$ is $w_{1}$ and $[v]_{A}$ is $w_{2}$. However, note that $[u]_{A}^{-1}$ is a proper subset of $[u]^{-1}$, which contradicts the minimality of $[u]^{-1}$. Therefore, $x$ has some neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$, and thus $[G]_{A}$ clearly has the desired properties. ∎
Figure 18. Graphs in Lemma 11. Solid line-segments are edges. Dashed line-segments are either edges or nonedges. Line-segments between $w_{2}$ and the blue circle represent an edge or nonedge between $w_{2}$ and some vertex in the circle.
Lemma 11 (Graphs in Figure 18 have $w_{1}w_{2}$-preserving forbidden minors).
For any edge $e$ of a graph $G$, if some minor $[G]$ is one of the graphs in Figure 18 such that $[e]$ is $w_{1}w_{2}$, then $G$ has an $e$-preserving forbidden minor.
Proof.
For each graph in Figure 18, a solid line-segment represents an edge and a dashed line-segment represents a nonedge. If a line-segment is between $w_{2}$ and the blue circle, then the endpoints of this edge or nonedge is $w_{2}$ and some vertex in the circle. Observe that each of these graphs has a $w_{1}w_{2}$-preserving forbidden minor. ∎
Proof of Proposition 1.
Assume that $G$ is $2$-connected and has a forbidden minor but no $e$-preserving forbidden minor. Then, there exists a minor $[G]$ in which $[e]$ is an edge and from which a forbidden minor can be obtained by contracting $[e]$. Hence, $[G]$ is one of the graphs in Figures 8, 18, or 17 such that $[e]$ is $w_{1}w_{2}$. Our assumptions along with Lemmas 10 and 11 show that $[G]$ is one of the graphs in Figure 8. This completes the proof. ∎
B.2. Proof of Proposition 2 [Winged graph minors with singelton wing tips]
The proof of Proposition 2 requires Lemma 12, below. The proof of this lemma requires the following operation, similar to a vertex exchange from Appendix B.1, that transforms one minor $[G]_{A}$ of a graph $G$ into another minor $[G]_{B}$. Refer to Figure 16. Consider an edge $[uv]_{A}$ of $[G]_{A}$, the subgraph $H$ of $G$ induced by the set $[u]_{A}{-1}$, a connected subgraph $J$ of $H$, and a connected component $K$ of $H\setminus J$ that contains a vertex with a neighbor in the set $[v]_{A}^{-1}$. The minor $[G]_{B}$ obtained by replacing $[u]_{A}^{-1}$ with $[u]_{A}^{-1}\setminus V(K)$ and $[v]_{A}^{-1}$ with $[v]_{A}^{-1}\cup V(K)$ is said to be obtained via the component exchange from $[u]_{A}^{-1}$ to $[v]_{A}^{-1}$ that exchanges $K$. For any $f$-winged graph minor $[G\cup f]$ of an winged graph pair $(G,f)$, with $f=uv$, if there exist vertices $s\in[u]^{-1}$ and $t\in[v]^{-1}$ such that either these are the only vertices in $[u]^{-1}\cup[v]^{-1}$ or $\{s,t\}$ is an $xy$-separator of $G\cup f$ for any two vertices $x$ in $[u]^{-1}\cup[v]^{-1}$ and $y$ not in $[u]^{-1}\cup[v]^{-1}$, then we call $s$ and $t$ wing tips of $[G\cup f]$.
Lemma 12 (Wing tips of winged graph minors).
For any winged graph pair $(G,f)$, with $f=uv$, and any $f$-winged graph minor $[G\cup f]$, there exists an $f$-winged graph minor $[G\cup f]_{A}$ with wing tips such that $[u]_{A}^{-1}\cup[v]_{A}^{-1}$ is a subset of $[u]^{-1}\cup[v]^{-1}$ and both $[f]_{A}$ and $[f]$ contain the same subset of edges of $G$.
Proof.
Let $[G\cup f]_{A}$ be any $f$-winged graph minor whose vertex set is $\{[a_{i}]_{A}\}$ such that (i) $[u]_{A}^{-1}\cup[v]_{A}^{-1}$ is a subset of $[u]^{-1}\cup[v]^{-1}$ and both $[f]_{A}$ and $[f]$ contain the same subset of edges of $G$ and (ii) $[u]_{A}^{-1}\cup[v]_{A}^{-1}$ has minimum size over all minors satisfying (i). Assume that there exist at least two vertices in $[u]_{A}^{-1}$ with neighbors in $[a_{3}]_{A}^{-1}\cup[a_{4}]_{A}^{-1}$. Then, we can choose $x$ to be one such vertex other than $u$. Let $H$ be the subgraph of $G\cup f$ induced by $[u]_{A}^{-1}$ and let $J$ be the connected component of $H\setminus x$ that contains $u$. Assume wlog that $x$ has a neighbor in $[a_{3}]_{A}^{-1}$. If $J$ contains some vertex with a neighbor in $[a_{4}]_{A}^{-1}$, then let $[G\cup f]_{B}$ be the minor obtained via the vertex exchange from $[u]_{A}^{-1}$ to $[a_{3}]_{A}^{-1}$ that exchanges $x$ and fixes $J$. Else, if $J$ contains some vertex with a neighbor in $[a_{3}]_{A}^{-1}$ and $x$ has a neighbor in $[a_{4}]_{A}^{-1}$, then let $[G\cup f]_{B}$ be the minor obtained via the vertex exchange from $[u]_{A}^{-1}$ to $[a_{4}]_{A}^{-1}$ that exchanges $x$ and fixes $J$. Lastly, if no vertex in $[a_{4}]_{A}^{-1}$ neighbors either $x$ or any vertex in $J$, then some connected component $K$ of $H\setminus x$ other than $J$ contains some vertex with a neighbor in $[a_{4}]_{A}^{-1}$. In this case, let $[G\cup f]_{B}$ be the minor obtained via the component exchange from $[u]_{A}^{-1}$ to $[a_{4}]_{A}^{-1}$ that exchanges $K$.
Next, in each of the above cases, the neighborhoods of vertices in $[G\cup f]_{A}$ along with Lemma 11 can be used to show that $[G\cup f]_{B}$ is an $f$-winged graph minor such that $[u]_{B}^{-1}\cup[v]_{B}^{-1}$ is a proper subset of $[u]_{A}^{-1}\cup[v]_{A}^{-1}$ and both $[f]_{B}$ and $[f]_{A}$ contain the same subset of edges of $G$. However, this contradicts Property (ii). Therefore, $[u]_{A}^{-1}$ contains exactly one vertex with a neighbor in $[a_{3}]_{A}^{-1}\cup[a_{4}]_{A}^{-1}$. Symmetrically, $[v]_{A}^{-1}$ contains exactly one vertex with a neighbor in $[a_{5}]_{A}^{-1}\cup[a_{6}]_{A}^{-1}$. Thus, $[G\cup f]_{A}$ satisfies Property (i) and has wing tips, as desired. ∎
Proof of Proposition 2.
Let $f=uv$ and consider any $f$-winged graph minor $[G\cup f]$ with wing tips $s$ and $t$, which exists by Lemma 12, and let $[u]^{-1}\cup[v]^{-1}$ have minimum size over all $f$-winged graph minors with wing tips. If $[u]^{-1}$ and $[v]^{-1}$ each contain exactly one vertex, then the lemma is proved. Otherwise, if $st$ is an edge of $G\cup f$, then $\{s,t\}$ is a cms of $G\cup f$. Since $G\cup f$ has no $f$-preserving forbidden minor, by assumption, clearly neither does any $\{s,t\}$-component $H$ of $G\cup f$ that contains $f$. Hence, since $|V(H)|<|V(G)|$, the converse direction of Theorem 4 along with Lemma 3 show that $(G,f)$ has the $3$-SIP. Therefore, assume that $st$ is a nonedge of $G\cup f$. We will show that either Lemma 2 can be used to prove that $(G,f)$ has the $3$-SIP or there exists an $f$-winged graph minor $[G\cup f]_{A}$ with wing tips such that $[u]_{A}^{-1}\cup[v]_{A}^{-1}$ is a proper subset of $[u]^{-1}\cup[v]^{-1}$. However, this contradicts the minimality of $[u]^{-1}\cup[v]^{-1}$, and so $(G,f)$ must have the $3$-SIP.
Since $f$ is an edge of $G\cup f$, $f$ is not $st$. Consequently, if $(G,st)$ and $(G\cup st,f)$ both have the $3$-SIP, then Lemma 2 shows that $(G,f)$ has the $3$-SIP. We will show that $(G\cup st,f)$ has the $3$-SIP. Observe that $\{s,t\}$ is a cms of $G\cup\{st,f\}$, and exactly one $\{s,t\}$-component $H$ of $G\cup\{st,f\}$ contains $f$. If $(H\setminus f,f)$ does not have the $3$-SIP, then the fact that $|V(H)|<|V(G)|$ along with the contrapositive of the converse direction of Theorem 4 shows that $H$ has an $f$-preserving forbidden minor. However, this clearly implies that $G\cup f$ has an $f$-preserving forbidden minor, which is not the case. Therefore, $(G\cup st,f)$ has the $3$-SIP.
Next, if $(G,st)$ has the $3$-SIP, then $(G,f)$ has the $3$-SIP and the proof is complete, as discussed above. Otherwise, note that $\{s,t\}$ is a clique separator of $G\cup st$, and so Lemma 3 shows that there exists a $\{s,t\}$-component $H$ of $G\cup st$ that contains $st$ and such that $(H\setminus st,st)$ does not have the $3$-SIP. The fact that $|V(H\setminus st)|<|V(G)|$ along with the contrapositive of the converse direction of Theorem 4 shows that $H$ has an $st$-preserving forbidden minor. Note that $H$ must contain $f$, or else $G\cup f$ clearly has an $f$-preserving forbidden minor, which is not the case. For the same reason, $H\cup f$ has no $f$-preserving forbidden minor. Also, since $G\cup f$ is $2$-connected, so is $H\cup f$. Therefore, Proposition 1 shows that $(H,f)$ is a winged graph pair with an $f$-winged graph minor $[H\cup f]_{B}$ whose vertex set is $\{[b_{i}]_{B}\}$.
Finally, since $H\cup f$ has an $st$-preserving forbidden minor, we can choose $[H\cup f]_{B}$ such that, wlog, $s$ is not contained in $[u]_{B}^{-1}\cup[v]_{B}^{-1}$. Assume that $s\in[b_{3}]_{B}^{-1}$. Using Lemma 12, we can additionally assume that $[H\cup f]_{B}$ has wing tips $s^{\prime}$ and $t^{\prime}$. Let $X$ be the set of all vertices in $G$ that are not contained in $H$ and consider the collection $\{[a_{i}]_{A}^{-1}\}$ of sets of vertices such that $[a_{3}]_{A}^{-1}=[b_{3}]_{B}^{-1}\cup X$ and $[a_{i}]_{A}^{-1}=[b_{i}]_{B}^{-1}$ for any $i\neq 3$. Note that no vertex in $X$ neighbors any vertex in $V(H)\setminus\{s,t\}$. Furthermore, since $st$ is an edge of $H\cup f$ and $s\in[b_{3}]_{B}^{-1}$, it is easy to see that there exists an $f$-winged graph minor $[G\cup f]_{A}$ with wing tips $s^{\prime}$ and $t^{\prime}$ and whose vertex set $\{[a_{i}]_{C}^{-1}\}$. Thus, since $[u]_{A}^{-1}\cup[v]_{A}^{-1}=[u]_{B}^{-1}\cup[v]_{B}^{-1}$ and $[u]_{B}^{-1}\cup[v]_{B}^{-1}\subsetneq V(H)\subseteq[u]^{-1}\cup[v]^{-1}$, we arrive at the desired contradiction. ∎
B.3. Proof of Proposition 3 [Pins of a winged graph minor]
The proof of Proposition 3 requires Lemmas 13, 14, and 15, below. The proofs of these lemmas are similar and involve inspecting many cases, so we omit the proofs of the latter two lemmas, but the proofs are available upon request.
Lemma 13 (Case: $|X|$ or $|Y|$ is $1$ and $|A|$ or $|B|$ is at least $2$).
Let $(G,f)$ be a winged graph pair, $[G\cup f]$ be a $f$-winged graph minor whose vertex set is $\{[c_{i}]\}$, $X\subseteq[c_{3}]^{-1}$ and $Y\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{1}]^{-1}$, and
  1. (1)
    if $[G\cup f]$ is the graph in Figure 8(a), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{6}]^{-1}$; otherwise,
  2. (2)
    if $[G\cup f]$ is the graph in Figure 8(b), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{7}]^{-1}$ and $[c_{6}]^{-1}\cup[c_{7}]^{-1}$, respectively.
If exactly one set in $\{X,Y\}$ has size $1$ and some set in $\{A,B\}$ has size at least $2$, then there exists an $f$-winged graph minor $[G\cup f]_{A}$ whose vertex set is $\{[d_{i}]_{A}\}$ such that $[d_{3}]_{A}^{-1}\cup[d_{4}]_{A}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$ and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same set of edges of $G$.
Proof.
We prove the lemma for the case where Statement (1) is true. The proof for the case where Statement (2) is true is similar, so we omit the details, but we highlight the differences at the end of the proof. Wlog, assume that $|Y|=1$ and $|X|>1$. We will show how to perform a sequence of vertex and component exchanges that transform $[G\cup f]$ into the $f$-winged graph minor $[G\cup f]_{A}$ with the desired properties. Let $H$ be the subgraph of $G\cup f$ induced by $[c_{3}]^{-1}$ and consider any vertex $a\in A$ with a neighbor in $[c_{5}]^{-1}$. Since $|X|>1$, some connected component $H_{1}$ of $H\setminus a$ contains a vertex in $X$. There are four cases.
Case 1: $H_{1}$ contains some vertices with neighbors in $[c_{4}]^{-1}$ and $[c_{6}]^{-1}$.
Let $[G\cup f]_{A}$ be obtained via the vertex exchange from $[c_{3}]^{-1}$ to $[c_{5}]^{-1}$ that exchanges $a$ and fixes $H_{1}$. Clearly, $[G\cup f]_{A}$ is one of the graphs in Figures 18(a) - 18(d) or 8(a) such that $[f]_{A}$ is $w_{1}w_{2}$, $[d_{3}]_{A}^{-1}\cup[d_{4}]_{A}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$, and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same subset of edges of $G$. Using Lemma 11, we see that either $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction, or $[G\cup f]_{A}$ has the desired properties.
Note 1.
In the remainder of the proof, we examine many cases for which similar arguments apply. As to not overburden the proof, we will state how to obtain $[G\cup f]_{A}$ and make observations about neighborhoods of vertices, and we leave it to the reader to use these observations along with Lemma 11 to verify that either $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction, or $[G\cup f]_{A}$ has the desired properties.
Case 2: $H_{1}$ does not contain a vertex with a neighbor in either $[c_{4}]^{-1}$ or $[c_{6}]^{-1}$.
Note that $H\setminus H_{1}$ contains some vertices with neighbors in $[c_{4}]^{-1}$ and $[c_{6}]^{-1}$, and let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $H_{1}$.
Case 3: $H_{1}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$ but no vertex with a neighbor in $[c_{4}]^{-1}$.
Let $a^{\prime}$ be any vertex in $H_{1}$ with a neighbor in $[c_{6}]^{-1}$ and note that $H\setminus H_{1}$ is a subgraph of some connected component $J_{1}$ of $H\setminus a^{\prime}$ and contains some vertex with a neighbor in $[c_{4}]^{-1}$. If some connected component $J_{2}$ of $H\setminus a^{\prime}$ other than $J_{1}$ contains some vertex in $X$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $J_{2}$. Otherwise, $J_{1}$ contains some vertex in $X$ since $|X|\geq 2$. Hence, let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[c_{3}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $a^{\prime}$ and fixes $J_{1}$.
Case 4: $H_{1}$ contains some vertex with a neighbor in $[c_{4}]^{-1}$ but no vertex with a neighbor in $[c_{6}]^{-1}$.
This case contains many subcases. The reader should keep in mind the edges and nonedges of $[G\cup f]$ as well as the assumptions leading to each subcase.
Note that $H\setminus H_{1}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$. If $H\setminus H_{1}$ contains some vertex with a neighbor in $[c_{4}]^{-1}$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $H_{1}$. Otherwise, if $H_{1}$ contains some vertex with a neighbor in $[c_{5}]^{-1}$, then let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[c_{3}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $a$ and fixes $H_{1}$. Hence, assume that $A\subseteq V(H-H_{1})$. If some connected component $H_{2}$ of $H\setminus a$ other than $H_{1}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $H_{2}$. Otherwise, $a$ has some neighbor in $[c_{6}]^{-1}$. If $H_{2}$ contains some vertex with a neighbor in $[c_{5}]^{-1}$, then $[G\cup f]_{A}$ is the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{5}]^{-1}$ that exchanges $H_{2}$.
Therefore, assume that $A=\{a\}$, and so $|B|\geq 2$ by assumption. Hence, there exists a vertex $b\in B\setminus Y$ since $|Y|=1$. Wlog, assume that $b$ neighbors some vertex in $[c_{5}]^{-1}$. Let $K$ be the subgraph of $G\cup f$ induced by $[c_{4}]^{-1}$ and let $K_{1}$ be the connected component of $K\setminus b$ that contains the vertex $y\in Y$. By arguments similar to those above, we can conclude that either (a) $K_{1}$ contains some vertex with a neighbor in $[c_{3}]^{-1}$ but no vertex with a neighbor in $[c_{6}]^{-1}$ or (b) $K_{1}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$ but no vertex with a neighbor in $[c_{3}]^{-1}$.
Assume that Statement (a) is true, and so $K\setminus K_{1}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$. If $K_{1}$ contains some vertex with a neighbor in $[c_{5}]^{-1}$, then let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $b$ and fixes $K_{1}$. Otherwise, we have that $B\subseteq V(K\setminus K_{1})$. Hence, some connected component $K_{2}$ of $K\setminus b$ other than $K_{1}$ contains some vertex in $B$. If $K_{2}$ contains some vertex with a neighbor in $[c_{6}]^{-1}$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $K_{2}$. Otherwise, $b$ must have some neighbor in $[c_{6}]^{-1}$. In this case, let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{5}]^{-1}$ that exchanges $K_{2}$.
Next, assume that Statement (b) is true, and so $K\setminus K_{1}$ contains some vertex with a neighbor $w\in[c_{3}]^{-1}$. By a previous assumption, we have $w\in V(H_{1})$. Hence, $H\setminus H_{1}$ is a subgraph of some connected component $J_{1}$ of $H\setminus w$. If some connected component $J_{2}$ of $H\setminus w$ other than $J_{1}$ contains some vertex in $X$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $J_{2}$. Otherwise, $J_{1}$ contains some vertex in $X$ since $|X|\geq 2$. If $w\in X$, then consider the minor $[G\cup f]_{B}$ obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $K_{1}$, and let it have have the vertex set $\{[b_{i}]_{B}\}$. Let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[b_{3}]_{B}^{-1}$ to $[b_{4}]_{B}^{-1}$ that exchanges $w$ and fixes $J_{1}$. Otherwise, if some connected component $H_{2}$ of $H\setminus a$ other than $H_{1}$ contains some vertex in $X$, then let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $H_{2}$. Therefore, it must be that $J_{1}\cap H_{1}$ contains some vertex in $X$. If $a\in X$, then consider the minor $[G\cup f]_{B}$ obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $K_{1}$, and let it have the vertex set $\{[b_{i}]_{B}\}$. Let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[b_{3}]_{B}^{-1}$ to $[b_{4}]_{B}^{-1}$ that exchanges $H_{1}$.
Lastly, by the above arguments, it suffices to assume that $X\subseteq V(J_{1}\cap H_{1})$. Consider any vertex $x_{1}\in X$ and let $R_{1}$ be the connected component of $H\setminus x_{1}$ that contains $w$. If $R_{1}$ contains $a$, then let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[c_{3}]^{-1}$ to $[c_{1}]^{-1}$ that exchanges $x_{1}$ and fixes $R_{1}$. Otherwise, let $R_{2}$ be the connected component of $H\setminus x_{1}$ that contains $a$. Since $|X|\geq 2$, there exists a vertex $x_{2}\in X$ in some connected component of $H\setminus x_{1}$. By the above arguments, this connected component must be either $R_{1}$ or $R_{2}$. If $x_{2}$ is contained in $R_{1}$, then consider the minor $[G\cup f]_{B}$ obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $K_{1}$, and let it have the vertex set $\{[b_{i}]_{B}\}$. Let $[G\cup f]_{A}$ be the minor obtained via the component exchange from $[b_{3}]_{B}^{-1}$ to $[b_{4}]_{B}^{-1}$ that exchanges $R_{1}$. Otherwise, let $[G\cup f]_{B}$ be obtained via the component exchange from $[c_{4}]^{-1}$ to $[c_{6}]^{-1}$ that exchanges $K_{1}$. Let $[G\cup f]_{A}$ be the minor obtained via the vertex exchange from $[b_{3}]_{B}^{-1}$ to $[b_{4}]_{B}^{-1}$ that exchanges $x_{1}$ and fixes $R_{2}$.
The cases above are exhaustive, and so the lemma is proved for the case where Statement (1) is true.
Finally, consider the case where Statement (2) is true. As stated at the beginning of this proof, the argument is similar, so we omit the details. However, we note the differences in the argument by examining the case where $H_{1}$ some contains vertices with neighbors in $[c_{4}]^{-1}$ and $[c_{7}]^{-1}$. Let $[G\cup f]_{A}$ be obtained via the vertex exchange from $[c_{3}]^{-1}$ to $[c_{5}]^{-1}$ that exchanges $a$ and fixes $H_{1}$. Clearly, $[G\cup f]_{A}$ is some graph $G^{\prime}$ that can be obtained from one of the graphs in Figures 18(e) - 18(i) or 8(b) by adding edges between its vertices, $[f]_{A}$ is $w_{1}w_{2}$, $[d_{3}]_{A}^{-1}\cup[d_{4}]_{A}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$, and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same subset of edges of $G$. If $G^{\prime}$ is the graph Figure 8(b), then $[G\cup f]_{A}$ has the desired properties. Otherwise, it is easy to see that $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction.
This completes the proof. ∎
Lemma 14 (Case: $|A|$ or $|B|$ is $1$ and $|X|$ and $|Y|$ are at least $2$).
Let $(G,f)$ be a winged graph pair, $[G\cup f]$ be any $f$-winged graph minor whose vertex set is $\{[c_{i}]\}$, $X\subseteq[c_{3}]^{-1}$ and $Y\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{1}]^{-1}$, and
  1. (1)
    if $[G\cup f]$ is the graph in Figure 8(a), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in either $[c_{5}]^{-1}$ or $[c_{6}]^{-1}$; otherwise,
  2. (2)
    if $[G\cup f]$ is the graph in Figure 8(b), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{7}]^{-1}$ and $[c_{6}]^{-1}\cup[c_{7}]^{-1}$, respectively.
If exactly one set in $\{A,B\}$ has size $1$ and both $X$ and $Y$ have size at least $2$, then there exists an $f$-winged graph minor $[G\cup f]_{A}$ whose vertex set is $\{[d_{i}]_{A}\}$ such that $[d_{3}]_{A}^{-1}\cup[d_{4}]_{A}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$ and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same set of edges of $G$.
Lemma 15 ($|X|$, $|Y|$, $|A|$, and $|B|$ are at least $2$).
Let $(G,f)$ be a winged graph pair, $[G\cup f]$ be any $f$-winged graph minor whose vertex set is $\{[c_{i}]\}$, $X\subseteq[c_{3}]^{-1}$ and $Y\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{1}]^{-1}$, and
  1. (1)
    if $[G\cup f]$ is the graph in Figure 8(a), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in either $[c_{5}]^{-1}$ or $[c_{6}]^{-1}$; otherwise,
  2. (2)
    if $[G\cup f]$ and the graph in Figure 8(b), then let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{7}]^{-1}$ and $[c_{6}]^{-1}\cup[c_{7}]^{-1}$, respectively.
If $X$, $Y$, $A$, and $B$ each has size at least $2$, then there exists an $f$-winged graph minor $[G\cup f]_{A}$ whose vertex set is $\{[d_{i}]_{A}\}$ such that $[d_{3}]_{A}^{-1}\cup[d_{4}]_{A}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$ and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same set of edges of $G$.
Proof of Proposition 3.
We proceed by induction on the size $n$ of $[a_{3}]^{-1}\cup[a_{4}]^{-1}$. The proposition is immediate when $n=2$. Assume the proposition is true when $n\leq k$ for any $k\geq 2$, and we will prove it when $n=k+1$. Let $X\subseteq[a_{3}]^{-1}$ and $Y\subseteq[a_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[a_{1}]^{-1}$. If $[G\cup f]$ is the graph in Figure 8(a), then let $A\subseteq[a_{3}]^{-1}$ and $B\subseteq[a_{4}]^{-1}$ be the subsets of all vertices with neighbors in either $[a_{5}]^{-1}$ or $[a_{6}]^{-1}$. Otherwise, if $[G\cup f]$ is the graph in Figure 8(b), then let $A\subseteq[a_{3}]^{-1}$ and $B\subseteq[a_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[a_{5}]^{-1}\cup[a_{7}]^{-1}$ and $[a_{6}]^{-1}\cup[a_{7}]^{-1}$, respectively. If either $|X|=|Y|=1$ or $|A|=|B|=1$ is true, then the proposition is immediate. Otherwise, by applying either Lemma 13, 14, or Lemma 15 we can obtain an $f$-winged graph minor $[G\cup f]_{A}$ whose vertex set is $\{[b_{i}]_{A}\}$ such that $[b_{3}]_{A}^{-1}\cup[b_{4}]_{A}^{-1}$ is a proper subset of $[a_{3}]^{-1}\cup[a_{4}]^{-1}$ and both $[f]_{A}^{-1}$ and $[f]^{-1}$ contain the same set of edges of $G$. Therefore, the inductive hypothesis states that there exists a minimal $uv$-separator $\{x^{\prime},y^{\prime}\}$ of $G\setminus[f]_{A}^{-1}$ such that $x^{\prime}\in[b_{3}]_{A}^{-1}$ and $y^{\prime}\in[b_{4}]_{A}^{-1}$.
Finally, observe that $x^{\prime},y^{\prime}\in[a_{3}]^{-1}\cup[a_{4}]^{-1}$ and $\{x^{\prime},y^{\prime}\}$ is a $uv$-separator of $G\setminus[f]^{-1}$. Hence, these observations along with the paths in $G\setminus[f]^{-1}$ implied by the existence of $[G\cup f]$ show that, wlog, $x^{\prime}\in[a_{3}]^{-1}$ and $y^{\prime}\in[a_{4}]^{-1}$. Thus, setting $x=x^{\prime}$ and $y=y^{\prime}$ completes the proof. ∎
B.4. Proof of Proposition 4 [Chained winged-graph minors]
The proof of Proposition 4 requires Lemmas 16 and 17, below.
Lemma 16 (No pin-preserving forbidden minor).
For any winged-graph pair $(G,f)$ and any pin $xy$ of an $f$-winged graph minor of $G\cup f$, $G\cup\{f,xy\}$ has no $f$-preserving forbidden minor.
Proof.
The lemma is immediate if $xy$ is an edge of $G\cup f$. Otherwise, assume to the contrary that $f$ is preserved in some forbidden minor $[G\cup\{f,xy\}]$. Let $[G\cup f]_{A}$ be an $f$-winged graph minor such that $xy$ is one of its pins. By Lemma 12, we can assume wlog that $[G\cup f]_{A}$ has wing tips $s$ and $t$. Since $[G\cup f]_{A}$ has at least two pins, by Proposition 3, $C=\{x,y,t\}$ is a separator of $G\cup\{f,xy\}$. Furthermore, since every forbidden minor is $4$-connected and has at least five vertices, there is a unique $C$-component $H$ of $G\cup\{f,xy\}$ that contains some vertex $w$ such that $[w]$ is not contained in $[C]$. If $H$ contains $f$, then the paths in $G\cup f$ given by the existence of $[G\cup f]_{A}$ imply the existence of some minor $[G\cup f]_{B}$ in which $H$ is preserved and $[C]_{B}$ is a clique. Combining these facts shows that $[G\cup f]_{B}$ has an $[f]_{B}$-preserving forbidden minor, and thus $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction.
Next, if $H$ does not contain $f$, then the existence of $[G\cup\{f,xy\}]$ implies the existence of a minor $[G\cup\{f,xy\}]_{D}$ that preserves $H$, has an $[f]_{D}$-preserving forbidden minor, and has the property that $[w]_{D}$ is contained in $[C]_{D}$ for every vertex $w\in V(G)\setminus V(H)$. Hence, both endpoints of $[f]_{D}$ are contained $[C]_{D}$. If $[f]_{D}=[xy]_{D}$, then $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$. Otherwise, assume wlog that $[f]_{D}=[xt]_{D}$. Then, we must have $[s]_{D}=[x]_{D}$. Since $\{x,y,t\}$ is a separator of $G\cup f$, $[y]_{A}^{-1}\setminus V(H)$ must be a subset of $[x]_{D}^{-1}\cup[y]_{D}^{-1}$. Therefore, it is easy to see that $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$. Thus, $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$ in either case, and so $G\cup f$ clearly has an $f$-preserving forbidden minor, which is a contradiction. Thus, $G\cup\{f,xy\}$ has no $f$-preserving forbidden minor. ∎
Lemma 17 (Graph-pin pairs have the $3$-SIP).
Consider any winged graph pair $(G,f)$ and any pin $xy$ of an $f$-winged graph minor $[G\cup f]$. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then $(G\setminus[f]^{-1},xy)$ has the $3$-SIP.
Proof.
The lemma is immediate if $xy$ is an edge of $G\setminus[f]^{-1}$. Otherwise, note that $\{x,y\}$ is a clique separator of $(G\setminus[f]^{-1})\cup xy$. Hence, by Lemma 3, it suffices to show that show that $(H\setminus xy,xy)$ has the $3$-SIP for each $\{x,y\}$-component $H$ of $(G\setminus[f]^{-1})\cup xy$ that contains $xy$. If $(H\setminus xy,xy)$ does not have the $3$-SIP, then, since $|V(H)|<|V(G)|$, the contrapositive of the converse direction of Theorem 4 shows that $H$ has an $xy$-preserving forbidden minor. However, this fact along with the paths in $G\setminus[f]^{-1}$ given by the existence of $[G\cup f]$ imply that $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus xy,xy)$ has the $3$-SIP, and the proof is complete. ∎
Proof of Proposition 4.
Let $(G^{\prime},f)$ be any $f$-winged graph pair with an $f$-winged graph minor $[G^{\prime}\cup f]$ such that (i) $G$ is a spanning subgraph of $G^{\prime}$ and $(G,f)$ has the $3$-SIP if $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley-connected and (ii) the number $k$ of pins of $[G^{\prime}\cup f]$ that are edges of $G^{\prime}\cup f$ is maximum over all winged graph pairs satisfying (i). Assume that some pin $xy$ of $[G^{\prime}\cup f]$ is a nonedge of $G^{\prime}\cup f$. Using Lemma 16, we see that $(G^{\prime}\cup xy,f)$ is a winged graph pair. Also, the minor $[G^{\prime}\cup\{f,xy\}]_{A}$ such that $[u]_{A}=[v]_{A}$ if and only if $[u]=[v]$, for any two vertices $u,v\in V(G^{\prime})$, is an $f$-winged graph minor. Note that $G$ is a spanning subgraph of $G^{\prime}\cup xy$ and the number of pins of $[G^{\prime}\cup\{f,xy\}]_{A}$ that are edges of $G^{\prime}\cup\{f,xy\}$ is strictly greater than $k$. Also, since $G$ is a spanning subgraph of $G^{\prime}$, Lemma 17 shows that $(G^{\prime}\setminus[f]^{-1},xy)$ has the $3$-SIP. Therefore, Lemma 2 shows that $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley connected if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected. Consequently, $(G,f)$ has the $3$-SIP if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected. Combining the above facts contradicts (ii). Thus, every pin of $[G^{\prime}\cup f]$ is an edge of $G^{\prime}\cup f$, and the proof is complete. ∎
B.5. Proof of Proposition 5[$(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected]
The proof of Proposition 5 requires Lemmas 19 and 22, below. Lemma 19 is proved using Lemma 18, also below.
Lemma 18 (No forbidden minor preserves endpoints of pins or $f$).
Consider any winged graph pair $(G,f)$ with a chained $f$-winged graph minor $[G\cup f]$. For any two distinct vertices $x$ and $w$ of $G$ such that $x$ is an endpoint of some pin of $[G\cup f]$ and $w$ is an endpoint of either some pin of $[G\cup f]$ or $f$, $(G\setminus[f]^{-1})\cup xw$ has no $xw$-preserving forbidden minor.
Proof.
Assume to the contrary that $xw$ is preserved in some forbidden minor of $G\setminus[f]^{-1}$. Let $xy$ be a pin of $[G\cup f]$. Since $xy$ is an edge of $G$ and no forbidden minor has a separator of size two, some $\{x,y\}$-component of $G\setminus[f]^{-1}$ has an $xw$-preserving forbidden minor. However, combining this with the existence of $[G\cup f]$ implies that $G\cup f$ has some $xw$-preserving forbidden minor that is also $f$-preserving, which is a contradiction. Thus, the lemma is proved. ∎
Consider a winged graph pair $(G,f)$ with $f=uv$ and a chained $f$-winged graph minor $[G\cup f]$ whose ordered set of pins is $\{x_{i}y_{i}\}$. Also, let $H_{u}(x_{i}y_{i})$ and $H_{v}(x_{i}y_{i})$ be the $\{x,y\}$-components of $G\setminus[f]^{-1}$ that contain $u$ and $v$, respectively. The graphs $H_{u}(x_{1}y_{1})$ and $H_{v}(x_{n}y_{n})$ and the $\{x_{i+1},y_{i+1}\}$-component of $H_{v}(x_{i}y_{i})$ that contains $x_{i}y_{i}$ are the links of $[G\cup f]$.
Lemma 19 (Links are $3$-connected).
Consider any winged graph pair $(G,f)$, where $G\cup f$ is $2$-connected, and any chained $f$-winged graph minor $[G\cup f]$ with singleton wing tips. If the converse direction of Theorem 4 is true for all graph-nonedge pairs whose graphs have strictly fewer vertices than $G$, then either $(G,f)$ has the $3$-SIP or each link of $[G\cup f]$ is $3$-connected.
Proof.
Assume that some link $L$ of $[G\cup f]$ is not $3$-connected. We will show that $(G,f)$ has the $3$-SIP. First, we show that $L$ is $2$-connected. If this is not the case, then $L$ has some separator $\{w\}$. Since each pin of $[G\cup f]$ is an edge of $G$ and $[G\cup f]$ has singleton wing tips, $\{w\}$ is clearly a separator of $G$. Also, since $G\cup f$ is $2$-connected, there are exactly two $\{w\}$-components of $G$: the unique ones containing the endpoints $u$ and $v$ of $f$, respectively. Hence, we have $w\notin\{u,v\}$. Letting the vertex set of $[G\cup f]$ be $\{[a_{i}]\}$, the above facts above show that $w\notin[a_{1}]^{-1}\cup[a_{2}]^{-1}$. However, the paths in $G$ given by the existence of $[G\cup f]$ show that $u$ and $v$ are contained in the same $\{w\}$-component of $G$, which is a contradiction. Therefore, $L$ is $2$-connected.
Next, since $L$ is $2$-connected but not $3$-connected, it has some minimal separator of size two. We prove that any such separator $\{x,y\}$ has the following four properties. Observe that the definition of a link of $[G\cup f]$ prevents $xy$ from being a pin of $[G\cup f]$. Let $wz$ be any pin of $[G\cup f]$ contained in $L$ and let $w^{\prime}\notin\{w,z\}$ be a vertex in $L$ that is an endpoint of either $f$ or a pin of $[G\cup f]$.
Property (i): $\{x,y\}$ is a separator of $G$. This follows easily from the fact that each pin of $[G\cup f]$ is an edge of $G$.
Property (ii): $w$, $z$, and $w^{\prime}$ are all contained in the same $\{x,y\}$-component of $G$. Since each pin of $[G\cup f]$ is an edge of $G$, $w$ and $z$ are contained in the same $\{x,y\}$-component $H$ of $G$. Hence, if $w^{\prime}$ is not contained in $H$, then $\{x,y\}$ must be a $uv$-separator of $G$, since $\{w,z\}$ is a $uv$-separator of $G$. Along with the fact that $[G\cup f]$ has singleton wing tips and the paths in $G$ given by the existence of $[G\cup f]$, we get that $xy$ is a pin of $[G\cup f]$, which contradicts the above observation. Therefore, $w^{\prime}$ is contained in $H$.
Property (iii): if $xy$ is a nonedge of $G$, then $(G\cup xy,f)$ has the $3$-SIP. Using Property (ii) and the paths in $G$ given by the existence of $[G\cup f]$, we see that $u$ and $v$ are contained in the same $\{x,y\}$-component of $G$. Hence, $\{x,y\}$ is a separator of $G\cup f$. Consequently, by Lemma 3, it suffices to show that $(H\setminus f,f)$ has the $3$-SIP for each $\{x,y\}$-component $H$ of $G\cup\{xy,f\}$ that contains $f$. Since $G\cup f$ has no $f$-preserving forbidden minor and $\{x,y\}$ is one of its minimal separators, $H$ has no $f$-preserving forbidden minor. Therefore, since $|V(H)|<|V(G)|$, the converse direction of Theorem 4 shows that $(H\setminus f,f)$ has the $3$-SIP.
Property (iv): if $xy$ is an edge of $G$, then $(G,f)$ has the $3$-SIP. The proof of this property is similar to the proof of Property (iii), so we omit the details.
Finally, we use Properties (i) - (iv) to show that $(G,f)$ has the $3$-SIP. If $xy$ is an edge of $G$, then this follows from Property (iv). Hence, assume that every separator of $L$ that has size two is a nonedge of $G$. Let $\{x,y\}$ now be a separator of $L$ such that the $\{x,y\}$-component $H_{wz}$ of $L$ containing $w$, $z$, and $w^{\prime}$, which exists by Property (ii), contains the maximum number of vertices. We will show that $(G,xy)$ has the $3$-SIP and then apply Property (iii) and Lemma 2 to complete the proof. By Property (i) and Lemma 3, it suffices to show that $(H\setminus xy,xy)$ has the $3$-SIP for each $\{x,y\}$-component $H$ of $L\cup xy$ that contains $xy$. To do this, we show that $H$ has no $xy$-preserving forbidden minor and then apply the converse direction of Theorem 4.
Assume to the contrary that some $\{x,y\}$-component $H$ of $L\cup xy$ has an $xy$-preserving forbidden minor. Using Lemma 18, we see that no $\{x,y\}$-component of $L\cup xy$ has a $wz$-preserving forbidden minor. Also, $H_{wz}$ is $2$-connected since $L$ is $2$-connected. These facts easily imply that $H$ is $H_{wz}$. Proposition 1 shows that $(H_{wz}\setminus wz,wz)$ is a winged graph pair. Using Lemma 10, we see that there exists a $wz$-winged graph minor $[H_{wz}]_{A}$ such that $x$ is not contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. Additionally, Lemma 12 allows us to assume that $[H_{wz}]_{A}$ has wing tips $x^{\prime}$ and $y^{\prime}$. Using Lemma 18, we see that $w$, $z$, and $w^{\prime}$ must be contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. Consequently, $\{x^{\prime},y^{\prime}\}$ is a separator of $H_{wz}$, and hence of $L$. Wlog, assume that $[x^{\prime}x]_{A}$ is an edge. Since $xy$ is an edge of $H_{wz}$, $y$ is either contained in $[x^{\prime}]_{A}^{-1}$ or not contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. In the former case, it must be that $y=x^{\prime}$. Therefore, in either case, all paths from any vertex in $\{w,z,w^{\prime}\}$ to any vertex in $\{x,y\}$ contain some vertex in $\{x^{\prime},y^{\prime}\}$. However, this implies that the the $\{x^{\prime},y^{\prime}\}$-component of $L$ that contains $w$, $z$, and $w^{\prime}$ has strictly more vertices than $H_{wz}$, which contradicts our maximality assumption on $\{x,y\}$. Thus, $H_{wz}$ has no $xy$-preserving forbidden minor, and so the proof is complete. ∎
Next, Lemma 22 is proved using Lemma 21, below, and Theorems 5 and 6, stated in Section 2 Lemma 21 is proved using Lemma 20, also below.
Lemma 20 ($3$-flattenable and not minimally $3$-connected implies partial $3$-tree).
If a graph $G$ is $3$-flattenable and $3$-connected and contains a $3$-connected proper spanning subgraph, then $G$ is a partial $3$-tree.
Proof.
Assume that $G$ is $3$-connected and $3$-flattenable and contains a $3$-connected proper spanning subgraph. In [5] it was shown that any $3$-flattenable graph is a subgraph of a graph that can be constructed from $K_{4}$’s, $V_{8}$’s, and $C_{5}\times C_{2}$’s using $1$-sums, $2$-sums, and $3$-sums. Furthermore, since neither $V_{8}$ nor $C_{5}\times C_{2}$ contains a $K_{3}$ subgraph, these graphs can only be involved in $1$-sums and $2$-sums. Hence, $G$ is a subgraph of some graph $H$ that can be constructed as described. Furthermore, since $G$ is $3$-connected, $H$ is either a $3$-tree, $V_{8}$, or $C_{5}\times C_{2}$. In the latter two cases, every vertex of $H$ has degree three. However, this implies that no proper spanning subgraph of $G$ is $3$-connected, which contradicts our assumption. Thus, $H$ is a $3$-tree, and so the lemma is proved. ∎
Lemma 21 (Edge additions to links that preserve partial $3$-tree).
Consider any winged graph pair $(G,f)$ with a chained $f$-winged graph minor $[G\cup f]$, and any of its links $L$ that is $3$-connected. For any two distinct vertices $x$ and $w$ of $L$ such that $x$ is an endpoint of some pin of $[G\cup f]$ and $w$ is an endpoint of either some pin of $[G\cup f]$ or $f$, $L\cup xw$ is a partial $3$-tree.
Proof.
Note that $L\cup xw$ is $3$-connected and $L$ is one of its $3$-connected spanning proper subgraphs. If we can show that $L\cup xw$ is $3$-flattenable, then the lemma follows from Lemma 20. Consequently, assume that $L\cup xw$ is not $3$-flattenable, and hence has a forbidden minor. By Lemma 18, $xw$ is contracted in every such minor. Hence, Proposition 1 shows that $(L,xw)$ is a winged graph pair. Therefore, either Proposition 3, if some $f$-winged graph minor of $L\cup xw$ has singleton wing tips, or Lemma 12, otherwise, shows that $L$ has a separator of size two. However, this contradicts the fact that $L$ is $3$-connected. Thus, $L\cup xw$ has no forbidden minor, and so the proof is complete. ∎
Lemma 22 ($(G,ux_{2})$ and $(G,uy_{2})$ both have the $3$-covering map property).
Let $(G,f)$ be a winged graph pair, with $f=uv$, that has a chained $f$-winged graph minor with singleton wing tips and whose ordered set of pins is $\{x_{i},y_{i}\}$. Also, consider the links $L_{0}$, containing $u$ and $x_{1}y_{1}$, and $L_{1}$, containing $x_{1}y_{1}$ and $x_{2}y_{2}$, of this minor. If $L_{0}$ and $L_{1}$ are $3$-connected, then $(G,ux_{2})$ and $(G,uy_{2})$ have the $3$-covering map property.
Proof.
Wlog, we will prove that $(G,ux_{2})$ has the $3$-covering map property. Since $\{x_{2},y_{2}\}$ is a clique separator of $G$, it suffices to show that $(L=L_{0}\cup L_{1},ux_{2})$ has this property. Consider any squared edge-length map $\ell$ and the set $\{X_{i}\}$ of connected components of the CS $\mathcal{C}^{3}(L,\ell)$. Given a set of nonedges $F$ of $L$, let $\phi_{F}$ be the Cayley map from $\mathcal{C}^{3}(L,\ell)$ to the CCS $\Omega^{3}_{F}(L,\ell)$. We must show that $\phi_{ux_{2}}(X_{i})=\phi_{ux_{2}}(X_{j})$ for any $i$ and $j$. To do this, we show that $\phi_{F_{0}\cup F_{1}}(X_{i})=\phi_{F_{0}\cup F_{1}}(X_{j})$, where $F_{0}$ is the set of nonedges in $\{ux_{1},uy_{1}\}$ and $F_{1}$ is the set of nonedges in $\{x_{1}x_{2},y_{1}x_{2}\}$. Then, for each point $p\in\phi_{F_{0}\cup F_{1}}(X_{i})$, the tetrahedral-inequality applied to the squared lengths $p$ and $\ell(x_{1}y_{1})$ solely determines the interval $\phi_{ux_{2}}(\phi^{-1}_{F_{0}\cup F_{1}}(p))$, because $\{x_{1},y_{1}\}$ is a clique separator of $L$. Combining the above facts proves the lemma.
We complete the proof by showing that $\phi_{F_{0}\cup F_{1}}(X_{i})=\phi_{F_{0}\cup F_{1}}(X_{j})$. Let $\ell_{0}$ and $\ell_{1}$ be the restrictions of $\ell$ to $E(L_{0})$ and $E(L_{1})$, respectively. Also, let $\{Y_{i}\}$ be the collection containing each connected component $A$ of $\mathcal{C}^{3}(L_{0},\ell_{0})$ and each set $A^{r}$ obtained from $A$ be performing an overall reflection on each of its realizations. Define the collection $\{Z_{i}\}$ using $\mathcal{C}^{3}(L_{1},\ell_{1})$ analogously. Since $\{x_{1},y_{1}\}$ is a clique separator of $L$, there exist components $Y_{j}$ and $Z_{k}$ such that each realization in $X_{i}$ can be obtained from some pair of realizations $(p\in Y_{j},q\in Z_{k})$ via rigid-body motions. Furthermore, since $Y_{j}$ and $Z_{k}$ are connected, it is easy to see that every pair of realizations $(p\in Y_{j},q\in Z_{k})$ can be combined via rigid-body motions to obtain a realization in $X_{i}$. Therefore, it suffices to show that $(L_{0},F_{0})$ and $(L_{1},F_{1})$ have the $3$-covering map property. Since $L_{0}$ is $3$-connected, Lemma 21 shows that $L_{0}\cup ux_{1}$ and $L_{0}\cup uy_{1}$ are partial $3$-trees. Thus, by Theorems 5 and 6, $(L_{0},F_{0})$ has the desired property. An identical argument shows that $(L_{1},F_{1})$ has the $3$-covering map property. This completes the proof. ∎
Proof of Proposition 5.
Let $\{x_{i}y_{i}\}$ be the ordered set of pins of the chained $f$-winged graph minor $[G\cup f]$. Consider the links $L_{0}$, containing $u$ and $x_{1}y_{1}$, and $L_{1}$, containing $x_{1}y_{1}$ and $x_{2}y_{2}$, of $[G\cup f]$, and let $L=L_{0}\cup L_{1}$. By Lemma 19, we can assume wlog that $L_{0}$ and $L_{1}$ are both $3$-connected. In this case, we will show that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. Since $[G\cup f]$ has singleton wing tips and its pins are edges of $G$, it suffices to show that $(L,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. By definition, we must show that the CCS $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is connected for any squared edge-length map $\ell$. Consider the CS $\mathcal{C}^{3}(L,\ell)$ and its set of connected components $\{X_{i}\}$. Let $Y_{i}$ be the image of $X_{i}$ under the Cayley map $\phi_{ux_{2},uy_{2}}$. Observe that $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is the union of the sets $Y_{i}$. Furthermore, each $Y_{i}$ is connected since $X_{i}$ is connected and $\phi_{ux_{2},uy_{2}}$ is continuous. Therefore, we complete the proof by showing that $Y_{i}\cap Y_{j}\neq\emptyset$ for any $i$ and $j$.
Let $Y_{i}(ux_{2})$ and $Y_{j}(ux_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$ onto $ux_{2}$, respectively. Also, let $Y_{i}(uy_{2})$ and $Y_{j}(uy_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$ onto $uy_{2}$, respectively. Since $Y_{i}$ is connected, it contains a continuous curve $\tau_{i}$ whose coordinate projections are $Y_{i}(ux_{2})$ and $Y_{i}(uy_{2})$. Similarly, $Y_{j}$ contains a continuous curve $\tau_{j}$ whose coordinate projections are $Y_{j}(ux_{2})$ and $Y_{j}(uy_{2})$. Additionally, Lemma 22 can be used to show that $Y_{i}(ux_{2})=Y_{j}(ux_{2})$ and $Y_{i}(uy_{2})=Y_{j}(uy_{2})$. Therefore, $\tau_{i}$ and $\tau_{j}$ are contained in $Y_{i}(ux_{2})\times Y_{i}(uy_{2})$. Thus, the Intermediate Value theorem shows that these curves share some point, and so $Y_{i}\cap Y_{j}\neq\emptyset$. This completes the proof. ∎
Appendix C Proofs of the forward direction propositions in Section 4
C.1. Proof of Proposition 6 [No Type (2) edge implies non-$3$-SIP map]
The proof of Proposition 6 requires Lemmas 23 and 25, below.
Lemma 23 (No Type (2) edge properties).
Let $(G,f)$ be a minimal graph-nonedge pair, where $f=uv$, $[G\cup f]$ be an $f$-preserving forbidden minor, and consider the subgraph $H=[G\cup f]\setminus\{[u],[v]\}$. If $G$ has no Type (2) edge, then the following statements are true:
  1. (i)
    any edge of $G$ that does not share an endpoint with $f$ is preserved in $[G\cup f]$,
  2. (ii)
    $|[w]^{-1}|=1$ for any vertex $[w]\in V(H)$,
  3. (iii)
    any vertex $x\in([u]^{-1}\cup[v]^{-1})\setminus\{u,v\}$ neighbors $u$, $v$, and at least two vertices in the subgraph $H^{-1}$ of $G\cup f$, but no vertex in $[x]^{-1}\setminus\{u,v\}$, and
  4. (iv)
    the sets $[u]^{-1}$ and $[v]^{-1}$ each have cardinality at most $2$.
Proof.
Assume that $G$ has no Type (2) edge. For Statement (i), consider any edge $e$ of $G$ that does not share an endpoint with $f$. Since $f$ is clearly retained in the minor of $G\cup f$ obtained by contracting $e$, $e$ is not of Type (3). Hence, $e$ must be of Type (1). Since $[G\cup f]$ is $f$-preserving, this implies it is also $e$-preserving.
For Statement (ii), consider any vertex $[w]\in V(H)$. If $|[w]^{-1}|>1$, then there exists an edge of $G$ that is contracted in $[G\cup f]$ and that does not share an endpoint with $f$. However, this contradicts Statement (i), and so we have $|[w]^{-1}|=1$.
For Statement (iii), if $([u]^{-1}\cup[v]^{-1})\setminus\{u,v\}$ is empty, then we are done. Otherwise, let $x$ be any vertex in this set and assume wlog that $x\in[u]^{-1}$. Since $x\notin\{u,v\}$, there exists an edge $xy$ of $G$ that is contracted in $[G\cup f]$. Statement (i) tells us that $u$ is the only neighbor of $x$ in $[x]^{-1}\setminus\{u,v\}$. Furthermore, $xu$ is clearly not of Type (1), and so it must be of Type (3). Hence, $xv$ is an edge of $G\cup f$. It remains to show that $x$ neighbors at least two vertices in $H^{-1}$.
Assume that $x$ does not neighbor any vertex in $H^{-1}$. Since $G\cup f$ is an atom, by assumption, $x$ neighbors some vertex $y$ other than $u$ or $v$. By assumption and the discussion above, we have $y\in[v]^{-1}\setminus\{v\}$. Let $J$ be the subgraph of $G\cup f$ induced by $[u]^{-1}$. By the discussion above, $J\setminus x$ is connected and contains $u$. Hence, the minor obtained via the vertex exchange (see Section B.1) from $[u]^{-1}$ to $[v]^{-1}$ that exchanges $x$ and fixes $J\setminus x$ is an $f$-preserving forbidden minor in which $xy$ is contracted. However, since neither $x$ nor $y$ is an endpoint of $f$, this contradicts Statement (i), and so $x$ must neighbor some vertex $w$ in $H^{-1}$. If $w$ is the only vertex in $H^{-1}$ that $x$ neighbors, then $xw$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor either is or has an $f$-preserving forbidden minor. Therefore, we arrive at a similar contradiction, and so $x$ neighbors at least two vertices in $H^{-1}$.
Finally, for Statement (iv), assume wlog that $|[u]^{-1}|>2$. Consider any two distinct vertices $x,y\in([u]^{-1}\setminus\{u\})$ and observe that $[u]$ has exactly three neighbors in $H$, say $[w_{1}]$, $[w_{2}]$, and $[w_{3}]$. By Statements (ii) and (iii), we can assume wlog that $x$ neighbors both $w_{1}$ and $w_{2}$. Furthermore, $x$ and $y$ share at least one neighbor in $H^{-1}$, say $w_{1}$. Wlog, we can assume that some vertex in $[u]^{-1}\setminus x$ neighbors $w_{3}$. Therefore, we arrive at a similar contradiction by considering the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w_{2}]^{-1}$ that exchanges $x$ and fixes $J\setminus x$. Thus, the lemma is proved. ∎
Lemma 25 is proved using Lemma 24, below.
Lemma 24 (Transferring non-$3$-SIP through $K_{4}$).
Let $G$ be obtained from $K_{4}$ be deleting one of its edges $f$, let $H$ be one of its $K_{3}$ subgraphs, and let $f^{\prime}$ be the edge of $H$ that does not share an endpoint with $f$. Also, let $\ell_{1}$ and $\ell_{2}$ be squared edge-length maps for $H$ that assign positive real numbers to each edge and differ only in the numbers they assign to $f^{\prime}$. If the linkages $(H,\ell_{1})$ and $(H,\ell_{2})$ both have some $3$-realization, then there exist squared edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for $G$ such that
  1. (i)
    they agree with each other on the edges not in $H$ and assign these edges positive real numbers,
  2. (ii)
    $\ell^{\prime}_{i}$ agrees with $\ell_{i}$ on the edges of $H$, and
  3. (iii)
    the CCSs $\Omega^{3}_{f}(G,\ell^{\prime}_{1})$ and $\Omega^{3}_{f}(G,\ell^{\prime}_{2})$ each contain exactly one positive value, and these values are distinct.
Proof.
Let the vertex set of $G$ be $\{v_{i}\}$, where $H$ does not contain $v_{1}$, $f=v_{1}v_{2}$, and $f^{\prime}=v_{3}v_{4}$. Also, let $a=\ell_{1}(v_{2}v_{3})$, $b=\ell_{1}(v_{2}v_{4})$, $c=\ell_{1}(v_{3}v_{4})$, and $d=\ell_{2}(v_{3}v_{4})$. Wlog, assume that $c<d$. Set $\ell_{1}^{\prime}(v_{1}v_{3})$ and $\ell_{1}^{\prime}(v_{1}v_{4})$ to be positive numbers such that $c=|\ell_{1}^{\prime}(v_{1}v_{3})-\ell_{1}^{\prime}(v_{1}v_{4})|$, $d=\ell_{1}^{\prime}(v_{1}v_{3})+\ell_{1}^{\prime}(v_{1}v_{4})$, and $\ell_{1}^{\prime}(v_{1}v_{3})$ is not equal to $a$ or $b$. It is easy to see that Conditions (i)-(iii) are satisfied, and so the lemma is proved. ∎
Lemma 25 (Every pair in Figure 12 has a proper non-$3$-SIP map).
Every graph-nonedge pair in Figure 12 has a proper non-$3$-SIP map.
Proof.
For the graph-nonedge pair $(G,f)$ in Figure 12(a), let $\ell$ be the squared edge-length map that assigns $1$ to each edge of $G$. It is easy to see that $\Omega^{3}_{f}(G,\ell)$ contains exactly two points: 00 and a non-zero value. A proper non-$3$-SIP map $\ell^{\prime}$ for the graph-nonedge pair $(G^{\prime},f^{\prime})$ in Figure 12(b) is constructed in the proof of Theorem 8 in [5]. Lemma 24, $\ell$, and $\Omega^{3}_{f}(G,\ell)$ can be used to construct a non-$3$-SIP squared edge-length map for the graph-nonedge pair in Figure 12(c). Similarly, Lemma 24, $\ell^{\prime}$, and $\Omega^{3}_{f^{\prime}}(G^{\prime},\ell^{\prime})$ can be used to construct a non-$3$-SIP squared edge-length map for the graph-nonedge pair in Figure 12(d). ∎
Proof of Proposition 6.
By Lemma 25, every graph-nonedge pair in Figure 12 has a proper non-$3$-SIP map. It remains to show that these are the only minimal graph-nonedge pairs with no Type (2) edge. Consider any graph-nonedge pair $(G,f)$, where $f=uv$. If $(G,f)$ is a pair in Figure 12, then it is easy to verify that $(G,f)$ is minimal and has no Type (2) edge. Next, we show that if $(G,f)$ is minimal and has no Type (2) edge, then it is some pair in Figure 12. Since $(G,f)$ is minimal, $G\cup f$ is an atom and has an $f$-preserving forbidden minor $[G\cup f]$. Furthermore, Statements (i)-(iv) of Lemma 23 are true and show that $5\leq|V(G)|\leq 8$. Let $H=[G\cup f]\setminus\{[u],[v]\}$ and consider the subgraph $H^{-1}$ of $G\cup f$. If $|[u]^{-1}|=|[v]^{-1}|=1$, then Statement (ii) shows that $(G,f)$ is the pair in either Figure 12(a) or 12(b). Hence, assume wlog that $|[u]^{-1}|>1$. By Statement (iv), we have $[u]^{-1}=\{u,x\}$. There are three cases.
Case 1: $u$ has no neighbor in $H^{-1}$.
If $|[v]^{-1}|=1$, then Statement (iii) shows that $\{x,v\}$ is a clique separator of $G\cup f$, contradicting the fact that $G\cup f$ is an atom. Otherwise, Statement (iv) shows that $[v]^{-1}=\{v,y\}$, and so all paths in $G\cup f$ from $u$ to any vertex in $H^{-1}$ contains some vertex in the set $\{v,x,y\}$. Furthermore, Statement (iii) shows that $xy$ is a nonedge and $vx$ and $vy$ are edges of $G\cup f$. Hence, $vy$ is clearly of Type (2), contradicting our assumption.
Case 2: $u$ has more than one neighbor in $H^{-1}$.
Let $w$ be the vertex in $H^{-1}$ that neighbors $x$ but not $u$ if such a vertex exists, or any vertex in $H^{-1}$ that neighbors $x$ otherwise. Also, let $J$ be the subgraph of $G\cup f$ induced by $[u]^{-1}$. By Statement (iii), $J\setminus x$ is connected and contains $u$. Hence, $xw$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor either is or has an $f$-preserving forbidden minor. However, this contradicts Statement (i).
Case 3: $u$ has exactly one neighbor in $H^{-1}$.
Let $w$ be the vertex in $H^{-1}$ that neighbors $u$. Note that $x$ neighbors each vertex in $H^{-1}\setminus w$. If $vw$ is not an edge of $G\cup f$, then $f$ is retained in the minor obtained by contracting $uw$, and this minor clearly has an $f$-preserving forbidden minor. Hence, $uw$ is not of Type (1) or (3). However, since $uw$ is not of Type (2) by assumption, it must be reducing. This contradicts the fact that $(G,f)$ is minimal, and so $vw$ must be an edge of $G\cup f$.
Next, consider the case where $|[v]^{-1}|=1$. If $xw$ is a nonedge of $G\cup f$, then $(G,f)$ is the pair in either Figure 12(c) or 12(d). Otherwise, $\{x,v,w\}$ is a clique separator of $G\cup f$, contradicting the fact that $G\cup f$ is an atom. Lastly, consider the case where $|[v]^{-1}|>1$. The discussion above shows that $w$ is the only neighbor of $v$ in $H^{-1}$. Let $w^{\prime}$ be any vertex in $H^{-1}\setminus w$ that neighbors $x$. Also, let $J$ be defined as in Case 2. Then, $xw^{\prime}$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w^{\prime}]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor has an $f$-preserving forbidden minor. However, this contradicts Statement (i).
The cases above are exhaustive, and so the lemma is proved. ∎
C.2. Proof of Proposition 7 [Graphs with edge-separating minors have an expanded edge-separating cms]
The proof of Proposition 7 requires Lemma 28, below. This lemma is proved using Lemma 27, below, which in turn is proved using Lemma 26, also below.
Lemma 26 (Edge-preserving forbidden minor implies edge-preserving induced forbidden minor).
If an edge of a graph is preserved in some forbidden minor, then some such minor is induced.
Proof.
If an edge $e$ of a graph $G$ is preserved in some $K_{5}$ minor, then it is clearly preserved in some induced $K_{5}$ minor. Otherwise, if $e$ is preserved in some $K_{2,2,2}$ minor, then consider the minor $[G]$ obtained by performing all contractions between $G$ and this $K_{2,2,2}$ minor. If $[G]$ is $K_{2,2,2}$, then we are done. Otherwise, $[G]$ has $K_{2,2,2}$ as a proper spanning subgraph, in which case $[G]$ has an $[e]$-preserving $K_{5}$ minor. This completes the proof. ∎
Lemma 27 (Edge-preserving forbidden minors extend minors of atoms).
If an edge of a graph $G$ is preserved in some forbidden minor, then some such minor is an extension of a minor of some atom of $G$.
Proof.
Observe that $K_{5}$ and $K_{2,2,2}$ are both atoms. Hence, the lemma follows by applying Lemma 26 followed by Lemma 58 in Appendix E. ∎
Lemma 28 (Graphs with edge-separating forbidden minors have an edge-separating pair).
Given a graph $G$ and one of its edges $e$, if $G$ has an $e$-separating forbidden minor, then that minor has an $e$-separating pair.
Proof.
By definition, any $e$-separating forbidden minor $[G]$ has an $[e]$-preserving forbidden minor but none of its atoms that contain $[e]$ have such a minor. Hence, Lemma 27 shows that some atom $M$ of $[G]$ does not contain some endpoint $[u]$ of $[e]$ and has an $[e]$-preserving forbidden minor. The above facts imply the existence of some clique minimal $[ux]$-separator $C$ of $[G]$, for any vertex $[x]$ of $M\setminus C$, such that $M\setminus C$ is non-empty. Thus, $(C,M)$ is an $e$-separating pair of $[G]$, and the proof is complete. ∎
Proof of Proposition 7.
If $[G]$ is an $e$-separating minor, then Lemma 28 shows that it has an $e$-separating pair. Thus, $G$ clearly has a top-level expanded $e$-separating pair, and hence a top-level expanded $e$-separating cms. ∎
C.3. Proof outline for Proposition 8 [Minimal pair inductive structure]
The proof of Proposition 8 is very technical, and so we split it into the following four sections.
C.3.1. Proof of Lemma 5[Top-Level Structure]
Here we prove Lemma 5 from Section 4 using Lemmas 30 and 32, below. Lemma 30 is proved using Lemma 29, also below.
Lemma 29 (Vertex-disjoint paths after contraction in $(H^{-1}\setminus f)\setminus C^{-1}$).
Let $C^{-1}$ be any expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. Also, consider the minor $[G\cup f]$ obtained by contracting any edge of $(H^{-1}\setminus f)\setminus C^{-1}$ and any $f$-preserving forbidden minor $[I^{-1}]_{A}$. There exists a pair $xy$ of distinct vertices in $C^{-1}$ that is preserved in $[I^{-1}]_{A}$ and some paths $P_{ux}$ and $P_{vy}$ in $G\cup f$ between $u$ and $x$ and between $v$ and $y$, respectively, that each contain exactly one vertex in $I^{-1}$ and such that $[P_{ux}]$ and $[P_{vy}]$ are vertex-disjoint.
Proof.
Since $f$ is preserved in $[I^{-1}]_{A}$ and $(G\cup f)\setminus C^{-1}$ is disconnected, there exists a pair of distinct vertices $x^{\prime}y^{\prime}$ in $C^{-1}$ and an extension $[G\cup f]_{B}$ of $[I^{-1}]_{A}$ such that $[f]_{B}=[x^{\prime}y^{\prime}]_{B}$. This implies that $x^{\prime}y^{\prime}$ is preserved in $[I^{-1}]_{A}$ and some vertex-disjoint paths $P_{ux^{\prime}}$ and $P_{vy^{\prime}}$ in $G\cup f$ between $u$ and $x^{\prime}$ and between $v$ and $y^{\prime}$, respectively, each contain exactly one vertex in $I^{-1}$. If $[P_{ux^{\prime}}]$ and $[P_{vy^{\prime}}]$ are vertex-disjoint, then setting $xy=x^{\prime}y^{\prime}$ proves the lemma.
Otherwise, letting $ab$ be the edge of $G\cup f$ that is contracted in $[G\cup f]$, it must be wlog that $P_{ux^{\prime}}$ contains $a$ and $P_{vy^{\prime}}$ contains $b$. Since $G\cup f$ is an atom, there exists a path $P_{vz}$ in $G\cup f$ between $v$ and some vertex $z$ in $C^{-1}$ that does not contain either $a$ or $b$. Wlog, we can choose $P_{vz}$ such that it contains exactly one vertex in $C^{-1}$ and does not contain $u$. There are two cases, each with several subcases. In each case, it will be clear that our choices for $P_{ux}$ and $P_{vy}$ each contain exactly one vertex in $I^{-1}$. We will show that these paths are vertex-disjoint and at least one of them does not contain $a$ or $b$. These facts imply that $[P_{ux}]$ and $[P_{vy}]$ are vertex-disjoint, which completes the proof.
Case 1: $[z]_{A}=[x^{\prime}]_{A}$.
There are several subcases. Given a path $R$ containing two vertices $s$ and $t$, $R(st)$ denotes the path in $R$ between $s$ and $t$. Additionally, we treat a path as a sequence of vertices and write the union of two paths $R_{1}$ and $R_{2}$ as $R_{1}R_{2}$.
Subcase 1: $P_{vz}$ does not contain any vertex in either $P_{ux^{\prime}}(ua)$ or $P_{vy^{\prime}}(by^{\prime})$.
Setting $x=y^{\prime}$, $y=z$, $P_{ux}=P_{ux^{\prime}}(ua)P_{vy^{\prime}}(by^{\prime})$, and $P_{vy}=P_{vz}$ proves the lemma.
Subcase 2: $P_{vz}$ contains some vertex in $P_{ux^{\prime}}(ua)$, and the last of its vertices $w$ contained in either $P_{ux^{\prime}}(ua)$ or $P_{vy^{\prime}}(vb)$ is contained in $P_{ux^{\prime}}(ua)$.
Since $P_{vz}$ does not contain either $a$ or $b$, by assumption, we have $w\neq a$. If $P_{vz}(wz)$ does not contain any vertex in $P_{vy^{\prime}}(by^{\prime})$, then set $x=z$, $y=y^{\prime}$, $P_{ux}=P_{ux^{\prime}}(uw)P_{vz}(wz)$, and $P_{vy}=P_{vy^{\prime}}$. By the properties of $w$ and $P_{vz}$, $P_{ux}$ does not contain either $a$ or $b$. Also, our assumptions imply that neither $P_{ux^{\prime}}(uw)$ nor $P_{vz}(wz)$ contains any vertex in $P_{vy^{\prime}}$, and hence $P_{ux}$ and $P_{vy}$ are vertex-disjoint.
Next, assume that $P_{vz}(wz)$ contains some vertex in $P_{vy^{\prime}}(by^{\prime})$, and let $s$ be the first of its vertices in either $P_{ux^{\prime}}(ax^{\prime})$ or $P_{vy^{\prime}}(by^{\prime})$. If $s$ is contained in $P_{ux^{\prime}}(ax^{\prime})$, then set $x=x^{\prime}$, $y=y^{\prime}$, $P_{ux}=P_{ux^{\prime}}(uw)P_{vz}(ws)P_{ux^{\prime}}(sx^{\prime})$, and $P_{vy}=P_{vy^{\prime}}$. Otherwise, set $x=y^{\prime}$, $y=x^{\prime}$, $P_{ux}=P_{ux^{\prime}}(uw)P_{vz}(ws)P_{vy^{\prime}}(sy^{\prime})$, and $P_{vy}=P_{vy^{\prime}}(vb)P_{ux^{\prime}}(ax^{\prime})$. In either case, a similar argument applies.
Subcase 3: $P_{vz}$ contains some vertex in $P_{ux^{\prime}}(ua)$, and the last of its vertices $w$ contained in either contained in either $P_{ux^{\prime}}(ua)$ or $P_{vy^{\prime}}(vb)$ is contained in $P_{vy^{\prime}}(vb)$.
This case is similar to Subcase 2, so we omit the details.
Subcase 4: $P_{vz}$ contains some vertex in $P_{vy^{\prime}}(by^{\prime})$ but no vertex in $P_{ux^{\prime}}(ua)$, and the first of its vertices $s$ contained in either $P_{ux^{\prime}}(ax^{\prime})$ or $P_{vy^{\prime}}(by^{\prime})$ is contained in $P_{ux^{\prime}}(ax^{\prime})$.
In this case, set $x=y^{\prime}$, $y=x^{\prime}$, $P_{ux}=P_{ux^{\prime}}(ua)P_{vy^{\prime}}(by^{\prime})$, and $P_{vy}=P_{vz}(vs)P_{ux^{\prime}}(sx^{\prime})$. An argument similar to those in Subcase 2 applies.
Subcase 5: $P_{vz}$ contains some vertex in $P_{vy^{\prime}}(by^{\prime})$ but no vertex in $P_{ux^{\prime}}(ua)$, and the first of its vertices $s$ contained in either $P_{ux^{\prime}}(ax^{\prime})$ or $P_{vy^{\prime}}(by^{\prime})$ is contained in $P_{vy^{\prime}}(by^{\prime})$.
In this case, set $x=x^{\prime}$, $y=y^{\prime}$, $P_{ux}=P_{ux^{\prime}}$, and $P_{vy}=P_{vz}(vs)P_{uy^{\prime}}(sy^{\prime})$. An argument similar to those in Subcase 2 applies.
Case 2: $[z]_{A}\neq[x^{\prime}]_{A}$.
This case is similar to Case 1, so we omit the details.
The cases above are exhaustive, and so the lemma is proved. ∎
Lemma 30 (No Type (1) edge in $(H^{-1}\setminus f)\setminus C^{-1}$).
Let $C^{-1}$ be any expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, with $f=uv$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. For any minor $[G\cup f]$ obtained by contracting any edge of $(H^{-1}\setminus f)\setminus C^{-1}$, $[I^{-1}]$, and hence $[G\cup f]$, has an $[f]$-preserving forbidden minor.
Proof.
By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{A}$. Hence, since $I^{-1}$ is preserved in $[G\cup f]$, it suffices to show that the edge contracted in $[G\cup f]$ is contracted in some extension of $[I^{-1}]_{A}$. This follows easily from Lemma 29. ∎
Next, Lemma 32 is proved using Lemma 31, below, which is in turn proved using Lemma 4 from Section 4. We prove Lemma 4 below. An induced subgraph $H$ of a graph $G$ is retained in a minor $[G]$ if each pair of vertices in $H$ is retained in $[G]$.
Proof of Lemma 4.
For Statement (1), since $(G\cup f)\setminus C^{-1}$ is disconnected and $G\cup f$ is an atom, $C^{-1}$ is not a clique. Hence, Figure 13 shows all cases for $C^{-1}$ if $|V(C^{-1})|\leq 5$. We complete the proof by showing that $|V(C^{-1})|\leq 5$. Assume to the contrary that this is not the case. Consider any $f$-separating minor $[G\cup f]_{A}$ such that $(C,M)$ is one of its $f$-separating pairs. Then, $C$ is a clique on at least five vertices. Consequently, since any atom $J$ of $[G\cup f]_{A}$ that contains $[f]_{A}$ has no $[f]_{A}$-preserving forbidden minor, by definition, $J$ cannot contain $C$. This implies that there exists a clique minimal $[ux]_{A}$-separator $E$ of $[G\cup f]_{A}$, for some endpoint $[u]_{A}$ of $[f]_{A}$ and any vertex $[x]_{A}$ in $C\setminus E$, such that $C\setminus E$ is non-empty. Clearly, $(E^{-1},M^{-1})$ is an expanded $f$-separating pair of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this shows that $C^{-1}$ is not in the top-level of $(G,f)$, which is a contradiction. Therefore, we have $|V(C^{-1})|\leq 5$, and so Statement (1) is proved.
Next, for Statement (2), observe that $C^{-1}$ is retained in $[H^{-1}\cup I^{-1}]$, and $[H^{-1}]$ and $[I^{-1}]$ are vertices of $[H^{-1}\cup I^{-1}]$. Hence, Statement (1) shows that $[C^{-1}]$ is one the graphs in Figure 13. Consequently, $[H^{-1}\cup I^{-1}]$ is some graph $G^{\prime}$ that can be obtained from one of the graphs in Figure 14 by deleting some edges, each of which has one endpoint in $\{a,b\}$ and endpoint in $\{w_{i}\}$. Furthermore, $[C^{-1}]$ is the subgraph of $G^{\prime}$ induced by the set $\{w_{i}\}$. If $G^{\prime}$ can be obtained without deleting any edges, then we are done. Otherwise, assume wlog that the above-mentioned minor $[G\cup f]_{A}$ is obtained by contracting $w_{1}w_{2}$. It is easy to see that either $G\cup f$ is not an atom or, for any endpoint $[u]_{A}$ of $[f]_{A}$ and any vertex $[x]_{A}$ of $M\setminus C$, $C$ is not a minimal $[ux]_{A}$-separator of $[G\cup f]_{A}$. In either case, we get a contradiction, and thus Statement (2) is proved. ∎
Lemma 31 (Weakly retained subgraphs that are contained in atoms).
Let $C^{-1}$ be an expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and $J^{-1}$ be either its expanded $f$-component or expanded minor component. Also, let $G^{\prime}=(G\cup f)\setminus(J^{-1}\setminus C^{-1})$ and $[G\cup f]$ be the minor obtained by contracting any edge of $G\cup f$. If $G^{\prime}$ is weakly retained in $[G\cup f]$, then $[G^{\prime}]\setminus E$ is connected for any cms $E$ of $[G\cup f]$ contained in $[J^{-1}]$.
Proof.
Assume that $G^{\prime}$ is weakly retained in $[G\cup f]$. Then, the endpoints of the edge contracted in $[G\cup f]$ must be contained in $J^{-1}$, and at least one of these endpoints is not contained in $C^{-1}$. Hence, we have $[J^{-1}]^{-1}=J^{-1}$. Since $E$ is contained in $[J^{-1}]$, this implies that $E^{-1}$ is contained in $J^{-1}$. Therefore, $[G^{\prime}]\setminus E$ is connected if and only if $G^{\prime}\setminus A$ is connected, where $A=E^{-1}\cap C^{-1}$. We complete the proof by showing that $G^{\prime}\setminus A$ is connected.
If $A$ is empty, then we are done. Otherwise, by definition, $G^{\prime}$ contains either the expanded $f$-component or expanded minor component of $C^{-1}$, say $K^{-1}$. Hence, using Lemma 4 Statement (2) and the observation that $A$ is a clique, it is easy to see that $K^{-1}\setminus A$ is connected. Furthermore, since $C^{-1}$ is not a clique, $C^{-1}\setminus A$ is non-empty. Therefore, since $G\cup f$ is an atom, for any vertex $x$ in $G^{\prime}\setminus K^{-1}$, there exists a path in $G\cup f$ between $x$ and some vertex in $C^{-1}\setminus A$ that does not contain any vertex of $A$. Note that we can choose this path such that it is contained in $G^{\prime}$. Combining this with the fact that $K^{-1}\setminus A$ is connected shows that $G^{\prime}\setminus A$ is connected. ∎
Lemma 32 (No Type (2) edge in $(H^{-1}\setminus f)\setminus C^{-1}$).
Let $C^{-1}$ be an expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let $H^{-1}$ be its expanded $f$-component. If $C^{-1}$ is in the top level of $(G,f)$, then no edge of $(H^{-1}\setminus f)\setminus C^{-1}$ is of Type (2).
Proof.
Assume to the contrary that $C^{-1}$ is in the top level of $(G,f)$ but some edge $ab$ of $(H^{-1}\setminus f)\setminus C^{-1}$ is of Type (2). Let $[G\cup f]$ be the $f$-separating minor obtained by contracting $ab$ and let $G^{\prime}=(G\cup f)\setminus(H^{-1}\setminus C^{-1})$. We will show that (i) $[G^{\prime}]$ is contained in some atom $N$ of $[G\cup f]$, (ii) there exists an $f$-separating pair $(E,N)$ of $[G\cup f]$, and (iii) $[C^{-1}]\setminus E$ is non-empty. Since $[C^{-1}]\setminus E$ is contained in both $[G^{\prime}]$ and $[H^{-1}]$, (i) - (iii) show that the $f$-component $J$ of $(E,N)$ is a proper subgraph of $[H^{-1}]$. Therefore, $J^{-1}$ is a proper subgraph of $H^{-1}$, which contradicts the fact that $C^{-1}$ is in the top-level of $(G,f)$. Thus, once we show (i) - (iii) are true, the proof will be complete.
To prove (i), consider any cms $E$ of $[G\cup f]$. Observe that $E$ must contain $[a]$, or else $E^{-1}$ is clearly a cms of $G\cup f$, which is a contradiction. Hence, since $E$ is a clique and $[C^{-1}]$ is a separator of $[G\cup f]$ that does not contain $[a]$, $E$ must be contained in $[H^{-1}]$. Consequently, since $G^{\prime}$ is retained in $[G\cup f]$, Lemma 31 shows that $[G^{\prime}]\setminus E$ is connected. Since $E$ was arbitrary, some atom $N$ of $G\cup f$ contains $G^{\prime}$, as desired.
Next, to prove (ii), note that $G^{\prime}$ contains the expanded minor component $I^{-1}$ of $C^{-1}$, and $[I^{-1}]$ has an $[f]$-preserving forbidden minor, by Lemma 30. This implies that $N$ has an $[f]$-preserving forbidden minor. Hence, since $[G\cup f]$ is $f$-separating, $N$ does not contain some endpoint $[u]$ of $[f]$, and so there exists a clique minimal $[ux]$-separator $E$ of $[G\cup f]$ for any vertex $[x]$ of $N\setminus E$. Therefore, $(E,N)$ is an $f$-separating pair of $[G\cup f]$.
Finally, to prove (iii), observe that $C^{-1}$ is retained in $[G\cup f]$. Hence, by Lemma 4 Statement (1), $[C^{-1}]$ is not a clique. Thus, since $E$ is a clique, we get that $[C^{-1}]\setminus E$ is non-empty, which completes the proof. ∎
Finally, we prove Lemma 5.
Proof of Lemma 5.
Since $w$ is a vertex of $H^{-1}\setminus(C^{-1}\cup\{u,v\})$, it must be adjacent to some vertex $z$ in $H^{-1}\setminus C^{-1}$. Note that Lemmas 30 and 32 imply that $wz$ is of Type (3). Hence, $z$ must be either $u$ or $v$, and both $wu$ and $wv$ are edges. Next, since $G\cup f$ is an atom, $w$ must be adjacent to some vertex other than $u$ or $v$. The same two lemmas imply that this vertex is contained in $C^{-1}$. Thus, the lemma is proved. ∎
C.3.2. Proof of Lemma 6 [Pruned possible graphs for $C^{-1}$]
Here we prove Lemma 6 using Lemma 34, below, which is in turn proved using Lemma 33, also below.
Lemma 33 (Each atom of top-level $f$-component contains $f$).
Let $(G,f)$ be a minimal graph-nonedge pair and let $C$ be a top-level $f$-separating cms of an $f$-separating minor $[G\cup f]$ whose $f$-component is $H$. Then, every atom of $H$ contains $[f]$.
Proof.
If $H$ is an atom, then we are done. Otherwise, it suffices to show that every cms of $H$ contains $[f]$. Assume to the contrary that some cms $E$ of $H$ does not contain some endpoint $[u]$ of $[f]$. We will show that this implies $C$ is not in the top-level of $[G\cup f]$, which is a contradiction. Let $H_{1}$ be the $E$-component of $H$ that contains $[u]$, $K$ be any other $E$-component of $H$, and $[x]$ be any vertex of $K\setminus E$. Since $[u]$ and $[x]$ are contained in distinct $E$-components of $H$, they are distinct vertices and $[ux]$ is a nonedge of $[G\cup f]$. Hence, Lemma 5 tells us that $[x]$ must be contained in $C$.
Next, since $C$ is a clique, $K$ is the unique $E$-component of $H$ that contains $C$. Hence, $E$ is a cms of $[G\cup f]$, $H_{1}$ is an $E$-component of $[G\cup f]$, and the $E$-component $J$ of $[G\cup f]$ that contains $K$ also contains $C$. This implies that $J$ contains $G^{\prime}=[G\cup f]\setminus(H\setminus C)$, which contains some atom $M$ such that $(C,M)$ is an $f$-separating pair of $[G\cup f]$. Therefore, some subgraph $E^{\prime}$ is such that $(E^{\prime},M)$ is an $f$-separating pair of $[G\cup f]$.
Finally, note that $E^{\prime}$ does not contain $[x]$. Hence, a similar argument shows that $G^{\prime}$ and $M$ are both contained in a unique $E^{\prime}$-component $J^{\prime}$ of $[G\cup f]$. Since $E^{\prime}$ is a minimal $[uy]$-separator for any vertex $[y]$ of $M\setminus E^{\prime}$, $J^{\prime}$ does not contain $[u]$. Thus, letting $H_{2}$ be the $E^{\prime}$-component of $[G\cup f]$ that contains $[u]$, we get that $H_{2}^{-1}$ is a proper subgraph of the expanded $f$-component of $C^{-1}$. This shows that $C$ is not in the top-level of $[G\cup f]$, as desired, and so the proof is complete. ∎
Figure 19. See Lemma 34. All possible expanded $f$-components of an expanded $f$-separating cms that has exactly five vertices and is in the top-level of a minimal graph-nonedge pair $(G,f)$. Dashed line-segments may be edges or nonedges.
Lemma 34 (Cases for $H^{-1}$ when $C^{-1}$ has $5$ vertices).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$. If $|V(C^{-1})|=5$, then the expanded $f$-component of $C^{-1}$ is one of the graphs in Figure 19 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the set $\{w_{i}\}$.
Proof.
Consider any $f$-separating minor $[G\cup f]$ obtained by contracting a single edge of $C^{-1}$ and such that $C$ is one of its $f$-separating cms’s. If the $f$-component $H$ of $C$ is the graph in Figure 8(a) such that $[f]$ is $w_{1}w_{2}$ and $C$ is the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$, then we can use Lemma 4 Statement (2) to show that the expanded $f$-component of $C^{-1}$ has the desired properties. Hence, we complete the proof by showing that $H$ has these properties. We prove this by first showing that (i) $H$ has no $[f]$-preserving forbidden minor. Second, we use (i) to show that (ii) some minor $[H]_{A}$ is the graph in Figure 8(a) such that $[f]_{A}$ is $w_{1}w_{2}$ and $|[y]_{A}^{-1}|=1$ for each vertex $y\in V(C)$. Third and finally, letting $f=uv$, we use (i) and (ii) to show that (iii) $|[u]_{A}^{-1}|=|[v]_{A}^{-1}|=1$. This completes the proof.
For Statement (i), since $H$ is a $C$-component of $[G\cup f]$, Lemma 60 in Appendix E states that any atom $J$ of $H$ is also an atom of $[G\cup f]$. Additionally, by Lemma 33, $J$ contains $[f]$. Since $[G\cup f]$ is $f$-separating, these facts imply that $J$ has no $[f]$-preserving forbidden minor. Therefore, Statement (i) follows from the contrapositive of Lemma 27 in Appendix C.2.
For Statement (ii), note that $C$ does not contain any endpoint of $[f]$, or else we can use Lemma 4 Statement (2) to show that $H$ has an $[f]$-preserving $K_{5}$ minor, which contradicts Statement (i). Using the same lemma, we see that $H$ has a $K_{5}$ minor $[H]_{B}$ such that $|[y]_{B}^{-1}|=1$, for each vertex $y\in V(C)$. By Statement (i), $f$ is contracted in $[H]_{B}$. Hence, there exists a minor $[H]_{A}$ from which $[H]_{B}$ can be obtained by contracting $[f]_{A}$. Clearly, $[H]_{A}$ is one of the graphs in Figures 8(a), 17(a), or 18(a)-18(d) such that $[f]_{A}$ is $w_{1}w_{2}$ and $[C]_{A}$ is the subgraph induced by $\{w_{3},w_{4},w_{5},w_{6}\}$. Since $G\cup f$ is an atom, $H$ is $2$-connected. Therefore, Lemma 10 allows us to assume that $[H]_{A}$ is one of the graphs in Figures 8(a) or 18(a)-18(d). Statement (ii) now follows from Statement (i) and Lemma 11.
Finally, for Statement (iii), wlog assume to the contrary that $|[u]_{A}^{-1}|>1$. If either $[u]$ has a neighbor in $C$ or $|[u]_{A}^{-1}|>2$, then we can use Lemma 5 to contradict Statement (i). Otherwise, let $[x]$ be the unique vertex in $[u]_{A}^{-1}\setminus\{[u]\}$. If $|[v]_{A}^{-1}|=1$, then $\{[x],[v]\}\}$ is a clique separator of $H$. However, since neither $[x]$ nor $[v]$ is contained in $C$, we have $1=|[x]^{-1}|=|[v]^{-1}|$. This implies that $G\cup f$ is not as atom, which is a contradiction. Hence, assume that $|[v]_{A}^{-1}|>1$. Similar arguments show that $[v]$ has no neighbor in $C$, there exists a unique vertex $[x^{\prime}]$ in $[v]_{A}^{-1}\setminus\{[v]\}$, and $1=|[x]^{-1}|=|[x^{\prime}]^{-1}|$. Hence, $\{x,x^{\prime}\}$ is a size two separator of $G\cup f$, and it is easy to see that in any $f$-preserving forbidden minor of $G\cup f$ some edge $e$ in the set $\{ux,ux^{\prime},vx,vx^{\prime}\}$ is contracted. Furthermore, the minor obtained by contracting $e$ has an $f$-separating pair whose $f$-component $J$ is such that $J^{-1}$ is a proper subgraph of $H^{-1}$. This contradicts the fact that $C$ is in the top-level of $[G\cup f]$, and so Statement (iii) is proved. This completes the proof. ∎
Proof of Lemma 6.
By Lemma 4 Statement (1), $C^{-1}$ is one of the graphs in Figure 13. We will show that $C^{-1}$ is not any of the graphs in Figures 13(e)-(i). Assume to the contrary that $C^{-1}$ is one of these graphs. We will show that $(G,f)$ has a reducing edge, and hence is not minimal, which is a contradiction. Let $H^{-1}$ and $I^{-1}$ be the expanded $f$-component and expanded minor component of $C^{-1}$, respectively. By Lemma 34, $H^{-1}$ is one of the graphs in Figure 19 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by the set $\{w_{i}\}$. We consider each case for $H^{-1}$.
If $H^{-1}$ is the graph in either Figure 19(a) or 19(b), then let $[G\cup f]$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]$ and $[H^{-1}]$ is an atom that contains $[f]$ such that $[H^{-1}]\cup[w_{1}w_{4}]$ has an $f$-preserving forbidden minor. Hence, some atom $J$ of $[G\cup f]$ contains $[H^{-1}]$. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Since $G^{\prime}$ is preserved in $[G\cup f]$, Lemma 4 Statement (2) tells us that some path in $[G^{\prime}]$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[C^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. Note that the only vertices in this path that are also in $[H^{-1}]$ are $[w_{1}]$ and $[w_{4}]$. Therefore, Lemma 66 shows that some path in $J$ between $[w_{1}]$ and $[w_{4}]$ does not contain any vertex in $[H^{-1}]\setminus\{[w_{1}],[w_{4}]\}$. The above facts imply that $J$ has an $[f]$-preserving forbidden minor, and so $uw_{1}$ is reducing, as desired.
If $H^{-1}$ is the graph in Figure 19(c), then let $[G\cup f]$ be the minor obtained by contracting $uw_{1}$. Observe that $[H^{-1}]$ is an atom with an $f$-preserving forbidden minor. Hence, since some atom of $[G\cup f]$ contains $[H^{-1}]$, $uw_{1}$ is reducing, as desired. A similar argument shows that $vw_{4}$ is reducing if $H^{-1}$ is the graph in Figure 19(d), and $vw_{3}$ is reducing if $H^{-1}$ is the graph in Figure 19(e).
Next, assume that $H^{-1}$ is the graph in either Figure 19(f), 19(g), 19(h), or 19(i), or $H^{-1}$ is the graph in 19(l) and $uw_{2}$ is an edge. Since $I^{-1}$ has an $f$-preserving forbidden minor, $G\cup f$ has an $f$-preserving forbidden minor in which some edge $e$ with $v$ as an endpoint is contracted. We will show that $e$ is reducing. The previous observation shows that $e$ is not of Type (1). Let $[G\cup f]$ be the minor obtained by contracting $e$. Note that $f$ is retained in $[G\cup f]$, and so $e$ is not of Type (3). We complete the proof by showing that $e$ is not of Type (2). Since $G^{\prime}$ is the union of $C^{-1}$ and some non-empty subset of connected components of $(G\cup f)\setminus C^{-1}$ and since $G\cup f$ is an atom, Lemma 61 in Appendix E states that, for any cms $E$ of $G^{\prime}$, each connected component of $G^{\prime}\setminus E$ contains some vertex of $C^{-1}\setminus E$. Along with the observation that $C^{-1}$ is an atom, this shows that $G^{\prime}$ is an atom. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, we get that $[G^{\prime}]$ is an atom. Additionally, note that $[H^{-1}]$ is an atom. Since $[C^{-1}]$ is not a clique, $[G\cup f]=[G^{\prime}]\cup[H^{-1}]$, and $[C^{-1}]=[G^{\prime}]\cap[H^{-1}]$, Lemma 64 shows that $[G\cup f]$ is an atom, and so $e$ is not of Type (2). Therefore, $e$ is reducing, as desired.
Finally, assume that $H^{-1}$ is the graph in Figure 19(j). If either $uw_{1}$ or $vw_{3}$ is not of Type (1), then an argument similar to those above shows that at least one of these edges is reducing. Hence, we complete the proof by showing that at least one of these edges is not of Type (1). By definition, $I^{-1}$ has an $f$-preserving forbidden minor $[I^{-1}]_{A}$. If either $w_{1}w_{4}$ or $w_{3}w_{5}$ is preserved in $[I^{-1}]_{A}$, then there exists an extension $[G\cup f]$ of $[I^{-1}]_{A}$ in which either $uw_{1}$ or $vw_{3}$ is contracted, and we are done. Otherwise, note that $w_{2}w_{3}$ is preserved in $[I^{-1}]_{A}$ since $f$ is preserved in $[I^{-1}]_{A}$. Therefore, there exists an extension $[G\cup f]$ of $[I^{-1}]_{A}$ in which both $uw_{1}$ and $vw_{3}$ are contracted, and hence neither of these edges is of Type (1). Similar arguments apply if either $H^{-1}$ is the graph in Figure 19(k) or $H^{-1}$ is the graph in Figure 19(l) and $uw_{2}$ is a nonedge.
The cases above are exhaustive, and so the proposition is proved. ∎
C.3.3. Proof of Lemma 7 [Unpruned possible graphs for top-level $H^{-1}$]
Here we prove Lemma 7 from Section 4. Consider a minimal graph-nonedge pair $(G,f)$ with an expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. This lemma enumerates all possible graphs for $H^{-1}$. Its proof is highly technical and has many cases, which are split amongst Lemmas 39, 40, and 41, below. The proofs of these lemmas require Lemmas 36 and 38, below, which are proved using Lemmas 35 and 37, also below.
Lemma 35 (Conditions such that $I^{-1}\setminus C^{-1}$ contains exactly one vertex).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component and expanded minor component be $H^{-1}$ and $I^{-1}$, respectively. Also, consider the minor $[G\cup f]$ obtained by contracting all edges in $I^{-1}\setminus C^{-1}$. If $F$ is a non-empty set of nonedges of $C^{-1}$ such that (i) some vertex of $C^{-1}$ is an endpoint of each nonedge in $F$, (ii) the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of each nonedge in $[F]$, and (iii) $H^{-1}\cup F$ has an $f$-preserving forbidden minor, then $I^{-1}\setminus C^{-1}$ contains exactly one vertex.
Proof.
Assume to the contrary that $F$ is a non-empty set of nonedges of $C^{-1}$ that satisfy Statements (i)-(iii) but $I^{-1}\setminus C^{-1}$ contains more than one vertex. Then, there exists some edge $e$ in $I^{-1}\setminus C^{-1}$. We will show that $e$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since neither endpoint of $e$ is contained in $H^{-1}$, $f$ is retained in be the minor $[G\cup f]_{A}$ obtained by contracting $e$, and so $e$ is not of Type (3). Combining this with Statements (i)-(iii) shows that $[G\cup f]_{A}$ has an $f$-preserving forbidden minor, and so $e$ is not of Type (1). We complete the proof by showing that $e$ is not of Type (2).
Since $G\cup f$ is an atom, any cms of $[G\cup f]_{A}$ must contain the vertex $[e]_{A}$. Hence, since $[C^{-1}]_{A}$ is a separator of $[G\cup f]_{A}$, every cms of $[G\cup f]_{A}$ is contained in $[I^{-1}]_{A}$. Furthermore, observe that $G^{\prime}=(G\cup f)\setminus(I^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]_{A}$. Therefore, Lemma 31 applies and shows that $[G^{\prime}]_{A}$ is contained in some atom $K$ of $[G\cup f]_{A}$. Note that $K$ contains $[H^{-1}]_{A}$, and so $K\cup[F]_{A}$ has an $[f]_{A}$-preserving forbidden minor, by Statement (iii). Also, since any cms of $[G\cup f]_{A}$ contains $[e]_{A}$, $K$ contains $[e]_{A}$. For each endpoint $y$ of each nonedge in $[F]_{A}$, Statement (ii) can be used to obtain some path in $[G\cup f]_{A}$ between $[e]_{A}$ and $y$ that contains exactly one vertex in $[H^{-1}]_{A}$. Consequently, Lemma 66 in Section E allows us to choose this path to be contained in $K$. Combining this with Statements (i)-(iii) shows that $K$ has an $[f]_{A}$-preserving forbidden minor, and thus $e$ is not of Type (2). ∎
Lemma 36 ($H^{-1}$ has no $f$-preserving forbidden minor).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. Then, the following statements are true.
  1. (1)
    If $C^{-1}$ is the graph in either Figure 13(a), 13(c), or 13(d), then, for any non-empty subset $F$ of nonedges of $C^{-1}$, $H^{-1}\cup F$ has no $f$-preserving forbidden minor.
  2. (2)
    If $C^{-1}$ is the graph in Figure 13(b), then there exists a nonedge $f^{\prime}$ of $C^{\prime}$ such that $H^{-1}\cup f^{\prime}$ has no $f$-preserving forbidden minor.
Consequently, $H^{-1}$ has no $f$-preserving forbidden minor.
Proof.
Let $I^{-1}$ be the expanded minor component of $C^{-1}$ and consider the minor $[G\cup f]$ obtained by contracting every edge in $I^{-1}\setminus C^{-1}$. Also, let $[G\cup f]_{A}$ be an $f$-separating minor such that $C$ is one of its $f$-separating cms’s. If Statement (1) is false, then some non-empty set $F$ of nonedges $C^{-1}$ is such that $H^{-1}\cup F$ has an $f$-preserving forbidden minor. Using Lemma 4 Statement (2), we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of each nonedge in $F$. If Statement (2) is false, then, for any nonedge $f^{\prime}$ of $C^{-1}$, $H^{-1}\cup f^{\prime}$ has an $f$-preserving forbidden minor. Using Statement (2) of Lemma 4, we see that the vertex $[I^{-1}\setminus C^{-1}]$ neighbors the endpoints of some nonedge of $C^{-1}$. In both cases, Lemma 35 applies and shows that $I^{-1}\setminus C^{-1}$ contains exactly one vertex. However, since $|V(C^{-1})|\leq 4$, this implies that $[I^{-1}]_{A}$ has no forbidden minor, which is a contradiction. This completes the proof. ∎
Lemma 37 (Properties of cms in minor of $H^{-1}$).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. Also, let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$ be weakly retained in some $f$-separating minor $[G\cup f]$ that is obtained by contracting some edge in $H^{-1}$. If $E$ is a cms of $[G\cup f]$ that is contained in $[H^{-1}]$, then any $E$-component of $[G\cup f]$ that does not contain some endpoint of $[f]$ does not have an $[f]$-preserving forbidden minor.
Proof.
Assume to the contrary that there exists a cms $E$ of $[G\cup f]$ that is contained in $[H^{-1}]$ and such that some $E$-component $J$ of $[G\cup f]$ that does not contain some endpoint of $[f]$ has an $[f]$-preserving forbidden minor. Wlog, we can assume that $J$ contains $E$. Let $M$ be the atom of $[G\cup f]$ that is contained in $J$ and has an $[f]$-preserving forbidden minor, given by Lemma 27. Then, there exists a subgraph $E^{\prime}$ of $E$ such that $(E^{\prime},M)$ is an $f$-separating pair of $[G\cup f]$. We will show that the $f$-component $K$ of $(E^{\prime},M)$ is a proper subgraph of $[H^{-1}]$. If this is true, then $K^{-1}$ is a proper subgraph of $H^{-1}$, since the edge contracted to obtain $[G\cup f]$ is contained in $H^{-1}$. However, this contradicts the fact that $C^{-1}$ is in the top-level of $G\cup f$, and so the lemma is proved.
We complete the proof by showing that the $f$-component $K$ of $(E^{\prime},M)$ is a proper subgraph of $[H^{-1}]$. Since $J$ contains $E$, $J$ is contained in some $E^{\prime}$-component $J^{\prime}$ of $[G\cup f]$. Also, since $G^{\prime}$ is weakly retained in $[G\cup f]$ and $E^{\prime}$ is contained in $[H^{-1}]$, Lemma 31 tells us that $[G^{\prime}]$ is contained in some $E^{\prime}$-component of $[G\cup f]$. If $J^{\prime}$ does not contain $[G^{\prime}]$, then $J^{\prime}$ is contained in $[H^{-1}]$. However, this implies that $H^{-1}$ has an $f$-preserving forbidden minor. Combining this with Lemma 6 contradicts Lemma 36. Therefore, assume that $J^{\prime}$ contains $[G^{\prime}]$. Since $C^{-1}$ is weakly retained in $[G\cup f]$ and is not a clique, by Lemma 4 Statement (1), $[C^{-1}]\setminus E^{\prime}$ is non-empty. Note that $[C^{-1}]\setminus E^{\prime}$ is contained in both $J^{\prime}\setminus E^{\prime}$ and $[H^{-1}]$. Consequently, we get that $K$ is a proper subgraph of $[H^{-1}]$, as desired. ∎
Lemma 38 (Edges that are not Type (2) in $H^{-1}$).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. Also, let $u$ be an endpoint of $f$ not in $C^{-1}$ and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. If there exists an edge $ab$ of $H^{-1}\setminus u$ such that $G^{\prime}$ is weakly retained in the minor obtained by contracting $ab$ and either
  1. (i)
    for any vertex $z$ of $C^{-1}\setminus\{a,b\}$, there exists some path in $H^{-1}$ between $z$ and $u$ that does not contain any vertex of $(C^{-1}\setminus z)\cup\{a,b\}$ or
  2. (ii)
    for any subgraph $X$ of $C^{-1}$ that is a clique and contains $V(C^{-1})\cap\{a,b\}$ and for any connected component $Y$ of $C^{-1}\setminus X$, there exists some path in $H^{-1}$ between $u$ and some vertex in $Y$ that does not contain any vertex of $(C^{-1}\setminus Y)\cup\{a,b\}$,
then $ab$ is not of Type (2).
Proof.
Assume to the contrary that some edge $ab$ as described in the Lemma statement is of Type (2). Then, the minor $[G\cup f]$ obtained by contracting $ab$ is $f$-separating, and hence has an $f$-separating pair $(E,M)$, by Lemma 28. We first make some observations about $E$ and the $E$-component $I$ of $[G\cup f]$ that contains $M$. By definition, $I$ does not contain some endpoint of $f$ and has an $[f]$-preserving forbidden minor. Hence, Lemma 37 applies and tells us that $E$ is not contained in $[H^{-1}]$. Also, since $u$ is not an endpoint of $ab$, $[u]$ is contained in $[H^{-1}]\setminus[C^{-1}]$. Hence, $[G\cup f]\setminus[C^{-1}]$ is disconnected. Since $E$ is a clique that is not contained in $[H^{-1}]$, this implies that $E$ is contained in $[G^{\prime}]$. Consequently, neither $E$ nor $I$ contains $[u]$.
Next, by the definition of $(E,M)$, for some endpoint $[v]$ of $f$ is and each vertex $[x]$ of $M\setminus E$, $E$ is a minimal $[vx]$-separator of $[G\cup f]$. Since $[f]$ is an edge of $[G\cup f]$ and $I$ does not contain $[u]$, this true when we set $v=u$. Furthermore, since $I\setminus E$ is connected, we get that $E$ is a minimal $[uy]$-separator of $[G\cup f]$ for each vertex $[y]$ of $I$. We complete the proof by contradicting this statement.
First, letting $X=E^{-1}\cap C^{-1}$, we show that $C^{-1}\setminus X$ is non-empty. Since $G\cup f$ is an atom, $E^{-1}$ must contain $ab$. Letting $A=E^{-1}\cap G^{\prime}$, the above facts show that $E^{-1}=A\cup a$. Since $G^{\prime}$ is weakly retained in $[G\cup f]$, $A$ is a clique. Consequently, $X$ is a clique. Also, since $C^{-1}$ is contained in $G^{\prime}$ and is not a clique, by Lemma 6, $C^{-1}\setminus X$ is non-empty.
Second, we show that $I^{-1}$ is contained in $H^{-1}\cup A$. Consider any vertex $w$ of $G\cup f$ and let $J^{-1}$ be the expanded $f$-component of $E^{-1}$. Note that $X$ contains $V(C^{-1})\cap\{a,b\}$. If $w$ is contained in $C^{-1}\setminus X$, then Statements (i) and (ii) and the properties of $E^{-1}$ show that $x$ is contained in $J^{-1}$. Else, if $x$ is contained in $G^{\prime}\setminus(C^{-1}\cup A)$, then, since $G\cup f$ is an atom, $A$ is a clique, and $C^{-1}\setminus X$ is non-empty, there exists some path in $G^{\prime}$ between $x$ and some vertex $z$ in $C^{-1}\setminus X$ that does not contain any vertex in $A$. Since we have shown that $z$ is contained in $J^{-1}$, we get that $x$ is contained in $J^{-1}$. Therefore, the only vertices potentially not in $J^{-1}$ are those in $H^{-1}\cup A$.
Finally, let $[z]$ be any vertex of $I\setminus E$. Since $I^{-1}$ is contained in $H^{-1}\cup A$ and each vertex of $C^{-1}\setminus X$ is contained in $J^{-1}$, $[z]$ is contained in $[H^{-1}]\setminus[C^{-1}]$. Furthermore, since $[C^{-1}]$ is a separator of $[G\cup f]$ and every path between $[z]$ and $[u]$ contains some vertex in $[E]$, by definition, every path between $[z]$ and $[u]$ contains some vertex in $E\cap[C^{-1}]$. Thus, since $E\cap[C^{-1}]$ is both a clique and a proper subgraph of $E$, $E$ is not a minimal clique $[uz]$-separator, which is a contradiction. This completes the proof. ∎
Lemma 39 (Cases for $H^{-1}$ when $C^{-1}$ is graph in Figure 13(a) and contains endpoint of $f$).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 13(a) and contains exactly one endpoint of $f$, then $H^{-1}$ is one of the graphs in Figures 15(a), 15(e), or 15(n) such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
Proof.
Let $f=uv$, where $u$ is not contained in $C^{-1}$. If $u$ is the only vertex of $H^{-1}\setminus C^{-1}$, then the fact that $G\cup f$ is an atom implies that $H^{-1}$ is one of the graphs in Figures 15(a) or 15(n) and has the desired properties. Otherwise, we will consider the two cases for $v$ in $C^{-1}$.
Case 1: $v$ is either $w_{1}$ or $w_{3}$.
We will show that some $H^{-1}$ is the graph in Figure 15(e) has the desired properties. Wlog, let $v=w_{1}$. By Lemma 5, each vertex of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors $u$, $w_{1}$, and at least one vertex of $C^{-1}\setminus w_{1}$. First we show that no such vertex neighbors $w_{2}$. If this is not the case, then let $x$ be any such vertex that neighbors $w_{2}$. We will show that $xw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{2}$, $xw_{2}$ is not of Type (3). To see that $xw_{2}$ is not of Type (1) or (2), consider the $f$-separating minor $[G\cup f]_{A}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating cms’s. Since $G\cup f$ is an atom, there exists some path in $G\cup f$ from $u$ to $w_{3}$ that does not contain any vertex in $\{x,w_{1},w_{2}\}$. Also, the definition of $C$ shows that some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor. By combining these facts, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{2}$ is not of Type (1). Lastly, observe that $G^{\prime}=(G\cup f)\setminus(H^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]$. Furthermore, Statement (i) of Lemma 38 is true if we set $ab=xw_{2}$. Consequently, this lemma shows that $xw_{2}$ is not of Type (2), and so the above facts show that $xw_{2}$ is reducing.
Next, we show that $u$ neighbors $w_{2}$. Assume this is not the case. The above argument shows that the neighborhood of any vertex $x$ in $H^{-1}\setminus(C^{-1}\setminus u)$ is $\{u,w_{1},w_{3}\}$. Since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, $xw_{3}$ is not of Type (3). Also, since some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, $[G\cup f]$ clearly has an $[f]$-preserving forbidden minor, and so $xw_{3}$ is not of Type (1). Hence, $xw_{3}$ must be of Type (2). Let $E$ be the subgraph of $[G\cup f]$ induced by $\{[w_{1}],[w_{3}]\}$ and let $I^{-1}$ be the expanded minor component of $C^{-1}$. Observe that $I^{-1}$ is preserved in $[G\cup f]$ and $E$ is a $[uy]$-separator for any vertex $[y]$ of $[I^{-1}]\setminus E$. Since $I^{-1}$ has an $f$-preserving forbidden minor, Lemma 27 can be used to show that there exists an atom $M$ of $[G\cup f]$ that has an $[f]$-preserving forbidden minor and such that $E$ is a $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Furthermore, since $G\cup f$ is an atom and $E$ is a clique separator of $[G\cup f]$ that contains exactly two vertices, it is easy to see that $E$ is a minimal $[uy]$-separator for any vertex $[y]$ of $M\setminus E$. Therefore, $(E,M)$ is an $f$-separating pair of $[G\cup f]$. However, note that the $[f]$-component of $E$ is $[H^{-1}]\setminus[w_{2}]$, and so the expanded $f$-component of $E^{-1}$ is $H^{-1}\setminus w_{2}$. This shows that $C^{-1}$ is not in the top-level of $G\cup f$, which is a contradiction.
Finally, we show that either $u$ neighbors $w_{3}$ or $H^{-1}\setminus(C^{-1}\setminus u)$ contains at least two vertices, which completes the proof. If this is not the case, then $H^{-1}\setminus(C^{-1}\setminus u)$ contains exactly one vertex $x$ and all paths in $H^{-1}$ between $u$ and $w_{3}$ contain some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$, then $C^{-1}$ is clearly not in the top-level of $G\cup f$, which is a contradiction. Otherwise, we arrive at a similar contradiction by considering the minor $[G\cup f]$ obtained by contracting $xw_{1}$ and the subgraph of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{2}]\}$.
Case 2: $v$ is $w_{2}$.
We will show this case is not possible. Consider the case where some vertex $x$ of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors each vertex of $C^{-1}$. Since $G\cup f$ is an atom, there exists a path in $H^{-1}$ between $u$ and $w_{1}$ that does not contain any vertex in $\{x,w_{2},w_{3}\}$. Similarly, there exists a path in $H^{-1}$ between $u$ and $w_{3}$ that does not contain any vertex in $\{x,w_{1},w_{2}\}$. Hence, $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving $K_{5}$ minor, which contradicts Lemma 36.
Next, consider the case where no vertex of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors each vertex of $C^{-1}$. A similar argument as in Case 1 shows that no vertex of $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors either $w_{1}$ or $w_{3}$. However, this implies that some vertex in $H^{-1}\setminus(C^{-1}\setminus u)$ neighbors only $u$ and $v$, which contradicts Lemma 5.
The cases above are exhaustive, and so the lemma is proved. ∎
Lemma 40 ($C^{-1}$ contains endpoint of $f$ if it is graph in Figure 13(a)).
If any top-level expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair $(G,f)$ is the graph in Figure 13(a), then $C^{-1}$ contains some endpoint of $f$.
Proof.
Assume to the contrary that $C^{-1}$ does not contain any endpoint of $f$. Let $f=uv$ and let $H^{-1}$ be the expanded $f$-component of $C^{-1}$. There are five cases.
Case 1: $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ is empty and either $u$ or $v$ neighbors $w_{2}$.
Wlog, assume that $v$ neighbors $w_{2}$. Since $G\cup f$ is an atom, $u$ neighbors some vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$, and either $u$ or $v$ neighbors $w_{3}$. Consider the case where $v$ neighbors $w_{3}$. If $v$ neighbors $w_{1}$, then $u$ must neighbor $w_{3}$. If, additionally, $u$ neighbors $w_{2}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36. Otherwise, an argument as in Case 1 of Lemma 39 shows that $vw_{2}$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. Therefore, assume $v$ does not neighbor $w_{1}$. If, additionally, $u$ does not neighbor $w_{3}$, then a similar argument shows that $uw_{1}$ is reducing, which is a contradiction. Otherwise, if $u$ neighbors $w_{2}$, then $\{u,w_{2},w_{3}\}$ is a cs of $G\cup f$, which is a contradiction. Else, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
Next, consider the case where $u$ neighbors $w_{3}$ and $v$ does not. If $u$ neighbors $w_{2}$, then the $\{u,w_{1},w_{2}\}$ is a cs of $G\cup f$, which is a contradiction. Otherwise, a similar argument shows that $vw_{2}$ is reducing, which is a contradiction.
Case 2: $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ is empty and no vertex in $\{u,v\}$ neighbors $w_{2}$.
Since $G\cup f$ is an atom, if either $u$ or $v$ neighbors both $w_{1}$ and $w_{3}$, then $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$. Consider the $f$-separating minor $[G\cup f]_{A}$ obtained by contracting some edge in $\{w_{1}w_{2},w_{2}w_{3}\}$ and that has $C$ as one of its $f$-separating cms’s. The definition of $C$ shows that some $C$-component of $[G\cup f]_{A}$ that does not contain $[f]_{A}$ has an $[f]_{A}$-preserving forbidden minor. This implies that some edge $e\in\{vw_{1},vw_{3}\}$ is contracted in $[G\cup f]_{A}$, and so $e$ is not of Type (1). Note that the unique atom of the minor $[G\cup f]$, obtained by contracting $e$, that contains $[u]$ is $K_{3}$, and hence $e$ is clearly not of Type (3). Consequently, $e$ is of Type (2). Let $E$ be the cms of $[G\cup f]$ induced by the set $\{[w_{1}],[w_{3}]\}$. Using an argument as in Case 1 of Lemma 39, we see that $E^{-1}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this contradicts the fact that $C^{-1}$ is in the top-level of $G\cup f$.
Next, assume that neither $u$ nor $v$ neighbors both $w_{1}$ and $w_{3}$. Since $G\cup f$ is an atom, we can assume wlog that $u$ neighbors $w_{1}$ and $v$ neighbors $w_{3}$. An argument as in Case 1 of Lemma 39, except using Statement (ii) of Lemma 38 instead of Statement (i), shows that $uw_{1}$ is reducing.
Case 3: $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains some vertex that neighbors $w_{2}$.
Let $x$ be any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ that neighbors $w_{2}$. Also, assume wlog that $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$. If every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$, then the subgraph induced by these three vertices is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, if some path in $G\cup f$ between either $u$ or $v$ and $w_{1}$ does not contain any vertex in $\{x,w_{2},w_{3}\}$, then an argument as in Case 1 of Lemma 39 shows that $xw_{2}$ is reducing, which is a contradiction. Else, note that $f$ is retained in the minor obtained by contracting $xw_{2}$, and so $xw_{2}$ is not of Type (3). Also, an argument as in Case 1 of Lemma 39 shows that $xw_{2}$ is not of Type (1). Hence, $xw_{2}$ of of Type (2). Combining our assumption on the paths in $G\cup f$ with an argument as in Case 1 of Lemma 39, we see that the subgraph of $G\cup f$ induced by $\{x,w_{2},w_{3}\}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
Case 4: $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ does not contain any vertex that neighbors $w_{2}$ and contains some vertex that neighbors exactly one vertex in $\{w_{1},w_{3}\}$.
By our assumption and Lemma 5, any vertex in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ neighbors some vertex in $\{w_{1},w_{3}\}$. Consider the case where some vertex $x$ in this subgraph neighbors exactly one vertex in $\{w_{1},w_{3}\}$, wlog say $w_{1}$. Also, assume that every path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ contains some vertex in $\{x,w_{1},w_{2}\}$. If $w_{1}w_{2}$ is contracted in $[G\cup f]_{A}$, then the subgraph induced by these three vertices is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction. Otherwise, since $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{1}$, $xw_{1}$ is not of Type (3). Also, since $G\cup f$ is an atom it has some path between either $u$ or $v$ and $w_{2}$ that does not contain any vertex in $\{w_{1},w_{2}\}$. Hence, $[G\cup f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $xw_{1}$ is not of Type (1). Consequently, $xw_{1}$ is of Type (2). We arrive at the same contradiction as above by using an argument as in Case 1 of Lemma 39.
Next, assume some path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ does not contain any vertex in $\{x,w_{1},w_{2}\}$. We will show that $xw_{1}$ is reducing with respect to $(G,f)$, which is a contradiction. As noted above, $xw_{1}$ is not of Type (3). A similar argument as above along with the existence of the above path shows that $xw_{1}$ is not of Type (1). Lastly, observe that $G^{\prime}=(G\cup f)\setminus(H^{-1}\setminus C^{-1})$ is weakly retained in $[G\cup f]$ and Statement (ii) of Lemma 38 is true if we set $ab=xw_{1}$. Therefore, this lemma shows that $xw_{1}$ is not of Type (2), and so $xw_{1}$ is reducing.
Case 5: $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ is non-empty and each of its vertices neighbors both $w_{1}$ and $w_{3}$ but not $w_{2}$.
If either $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains at least three vertices or $u$ and $v$ each neighbor both $w_{1}$ and $w_{3}$, then $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36. Also, if exactly one vertex in $\{u,v\}$, wlog say $v$, neighbors either $w_{2}$ or exactly one vertex in $\{w_{1},w_{3}\}$, then an argument as in Case 1 shows that some edge incident on $v$ is reducing, which is a contradiction. Hence, assume none of these situations occur. Additionally, assume that $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains exactly two vertices, say $x$ and $y$. If all paths between either $u$ or $v$ and any vertex in $C^{-1}$ contains some vertex in $\{x,y\}$, then consider the minor $[G\cup f]$ obtained by contracting $vx$. The unique atom of $[G\cup f]$ that contains $[f]$ is $K_{3}$, and so $vx$ is not of Type (3). Also, since $[G\cup f]_{A}$ has an $[f]_{A}$-preserving forbidden minor, it is easy to see that $[G\cup f]$ has an $[f]$-preserving forbidden minor, and so $vx$ is not of Type (3). Consequently, $vx$ is of Type (2). Using as argument as in Case 1 of Lemma 39, we see that the subgraph of $G\cup f$ induced by $\{v,x,y\}$ is an expanded $f$-separating cms of $G\cup f$ whose expanded $f$-component is a proper subgraph of $H^{-1}$, which is a contradiction.
Next, assume that some path between either $u$ or $v$ and some vertex in $C^{-1}$ does not contain any vertex in $\{x,y\}$. Since $x$ and $y$ are the only vertices in $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$, every such path contains exactly one edge. For all cases of the neighborhoods of $u$ and $v$ that satisfy the above assumptions, we see that $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36.
Finally, assume that $H^{-1}\setminus(C^{-1}\setminus\{u,v\})$ contains exactly one vertex, say $x$. An argument as in Case 4 shows that some path in $G\cup f$ between either $u$ or $v$ and $w_{3}$ does not contain any vertex in $\{x,w_{1},w_{2}\}$. Similarly, some path in $G\cup f$ between either $u$ or $v$ and $w_{1}$ does not contain any vertex in $\{x,w_{2},w_{3}\}$. These facts along with our assumption show that $w_{1}$ and $w_{3}$ each neighbor some vertex in $\{u,v\}$. If either $u$ or $v$ neighbors $w_{2}$, then the above facts imply the existence of edges such that $H^{-1}\cup w_{1}w_{3}$ has an $f$-preserving forbidden minor, which contradicts Lemma 36. Otherwise, the above facts imply the existence of edges such that $G\cup f$ is not an atom, which is a contradiction.
The above cases are exhaustive, and so the lemma is proved. ∎
Lemma 41 (Remaining cases for $H^{-1}$).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$, and let its expanded $f$-component be $H^{-1}$. If $C^{-1}$ is the graph in Figure 13(b), 13(c), or 13(d), then $H^{-1}$ is one of the graphs in Figure 15 not mentioned in Lemma 39 such that $f$ is the red edge and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
The proof of Lemma 41 is omitted here since it is very similar to the proofs of Lemmas 39 and 40, but it is available upon request.
Proof of Lemma 7.
The lemma follows immediately from Lemmas 6, 39, 40, and 41. ∎
C.3.4. Proof of Proposition 8 [Minimal pair inductive structure]
Consider a minimal graph-nonedge pair $(G,f)$ with an expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$. First, we prove Lemma 44, below, which eliminates some of the graphs in Figure 15 as possibilities for $H^{-1}$. The proof of this lemma requires Lemmas 42 and 43, also below.
Figure 20. Graphs in the proof of Lemma 42.
Lemma 42 ($I^{-1}$ has $f^{\prime}$-preserving forbidden minor for each nonedge of $C^{-1}$).
Let $C^{-1}$ be a top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$ whose expanded minor component is $I^{-1}$. $I^{-1}$ has an $f^{\prime}$-preserving forbidden minor for each nonedge $f^{\prime}$ of $C^{-1}$.
Proof.
By the definition of $C^{-1}$, $C$ is an $f$-separating cms of some minor $[G\cup f]$, and the minor component $I$ of $C$ has an $[f]$-preserving forbidden minor $[I]_{A}$. Note that each nonedge $f^{\prime}$ of $C^{-1}$ is preserved in $[G\cup f]$. Hence, it suffices to show that $I$ has an $[f^{\prime}]$-preserving forbidden minor, which we demonstrate now. By Lemma 6, $C^{-1}$ is one of the graphs in Figures 13(a)-13(d). This implies that $C$ is either $K_{2}$ or $K_{3}$. If $C$ is $K_{2}$, then $[I]_{A}$ is $[f^{\prime}]$-preserving, since it is $[f]$-preserving. Otherwise, let $[I]_{B}$ be the minor obtained by contracting all edges that are contracted in $[I]_{A}$ except for any edge of $C$. If $[I]_{B}=[I]_{A}$, then we are done. Otherwise, since $[I]_{A}$ is $[f]$-preserving, it is obtained from $[I]_{B}$ by contracting exactly one edge of $[C]_{B}$. Since $[I]_{A}$ is either $K_{5}$ or $K_{2,2,2}$, this implies that some spanning subgraph of $[I]_{B}$ is one of the graphs in Figure 20 such that $[C]_{B}$ is the subgraph induced by $\{x_{i}\}$.
Next, assume wlog that contracting $x_{1}x_{2}$ yields $[I]_{A}$. If the above-mentioned some spanning subgraph of $[I]_{B}$ is the graph in Figure 20(a), then $[I]_{A}$ is $K_{5}$. This implies that $x_{3}$ neighbors each vertex in $\{y_{1},y_{2},y_{3}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and wlog $x_{1}$ has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{B}$-preserving $K_{5}$ minor of $[I]_{B}$, and hence an $[f^{\prime}]$-preserving $K_{5}$ minor of $I$.
Finally, if the above-mentioned some spanning subgraph of $[I]_{B}$ is the graph in Figure 20(b), then $[I]_{A}$ is $K_{2,2,2}$. This implies wlog that $x_{3}$ neighbors each vertex in $\{y_{1},y_{3},y_{4}\}$, the union of the neighborhoods of $x_{1}$ and $x_{2}$ is $\{y_{1},y_{2},y_{3},x_{3}\}$, and $x_{1}$ wlog has at least two neighbors in $\{y_{1},y_{2},y_{3}\}$, say $y_{1}$ and $y_{2}$. For all cases of the neighborhoods of $x_{1}$ and $x_{2}$, we can clearly obtain an $[f^{\prime}]_{B}$-preserving forbidden minor of $[I]_{B}$, and hence an $[f^{\prime}]$-preserving forbidden minor of $I$. This completes the proof. ∎
[…]
Lemma 43 (Structure necessary for induction).
Let $C^{-1}$ be any top-level expanded $f$-separating cms of a minimal graph-nonedge pair $(G,f)$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Some nonedge $f^{\prime}$ of $C^{-1}$ is contained in some atom $J$ of $G^{\prime}\cup f^{\prime}$ that has an $f^{\prime}$-preserving forbidden minor. If $C^{-1}$ is the graph in Figure 13(b), then this is true for any nonedge of $C^{-1}$. Furthermore, no clique subgraph of $C^{-1}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$.
Proof.
Lemma 42 states that the expanded minor component of $C^{-1}$ has an $f^{\prime}$-preserving forbidden minor for each nonedge $f^{\prime}$ of $C^{-1}$. Hence, Lemma 27 shows that some atom $J$ of $G^{\prime}\cup f^{\prime}$ has an $f^{\prime}$-preserving forbidden minor. First, we demonstrate that we can choose $f^{\prime}$ to be contained in $J$. Note that $G\cup f$ is an atom, $(G\cup f)\setminus C^{-1}$ is disconnected, and $G^{\prime}$ is the union of $C^{-1}$ and some non-empty set of connected components of $(G\cup f)\setminus C^{-1}$. Furthermore, by inspecting all possibilities for $C^{-1}$ given by Lemma 6, we see that either $C^{-1}$ is an atom or, for any cms $E$ of $C^{-1}$, there are exactly two $E$-components of $C^{-1}$, both of which are atoms. Therefore, Lemma 63 in Appendix E shows that the atom graph of $G^{\prime}$ is a path.
Next, if $G^{\prime}$ is an atom, then Lemma 59 Statement (1) in Appendix E shows that any choice of $f^{\prime}$ is contained in $J$. Note that if $C^{-1}$ is the graph in Figure 13(b), then $C^{-1}$ is an atom. Hence, Lemma 61 in Appendix E shows that $G^{\prime}$ is an atom. Combining these facts demonstrates that any choice for $f^{\prime}$ is contained in $J$. Otherwise, if $G^{\prime}$ is not an atom, then the atom graph of $G^{\prime}$ has two distinct leaves $y_{1}$ and $y_{2}$ with not necessarily distinct neighbors $x_{1}$ and $x_{2}$, respectively. Using an argument similar to the proof of Lemma 63, we see that $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ contains some vertex $u$ of $C^{-1}$, $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ contains some vertex $v$ of $C^{-1}$, and $G^{\prime}(y_{1})\setminus G^{\prime}(x_{1})$ and $G^{\prime}(y_{2})\setminus G^{\prime}(x_{2})$ do not share any vertices. These facts show that $uv$ is a nonedge of $C^{-1}$. Therefore, Lemma 59 Statement (2) shows that $J$ contains $uv$.
Lastly, we show that no clique subgraph of $C^{-1}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. It suffices to consider any clique subgraph $E$ of $C^{-1}$ that is contained in $J$ and does not contain any endpoint of $f^{\prime}$. Since $J$ has a forbidden minor and $|V(C^{-1})|\leq 4$, $J$ contains some vertex $x$ of $G^{\prime}\setminus C^{-1}$. Note that $E$ does not contain $x$. Consider the case where $C^{-1}$ is the graph in Figure 13(b), and let $f=uv$. If $J$ either contains $C^{-1}$ or does not contain any vertex in $C^{-1}\setminus\{u,v\}$, then the claim is immediate. Otherwise, let $y$ be the unique vertex contained in both $J$ and $C^{-1}\setminus\{u,v\}$. Note that the only case for $E$ is the subgraph induced by $\{y\}$. Since $J$ is an atom $vy$ is an edge, there exists a path in $J$ between $x$ and $u$ that does not contain any vertex in $\{v,y\}$. Similarly, there exists a path in $J$ between $x$ and $v$ that does not contain any vertex in $\{u,y\}$. Since these paths do not contain $f^{\prime}$ or $y$, they connect $u$ and $v$ in $(J\setminus f^{\prime})\setminus y$.
Finally, if $C^{-1}$ is not the graph in Figure 13(b), then Lemma 4 Statement (2) shows that there exist paths in $G^{\prime}$ that connect $x$ to the endpoints of each nonedge of $C^{-1}$ and that each contain exactly one vertex in $C^{-1}$. Furthermore, Lemma 66 shows that we can choose these paths to be contained in $J$. Therefore, clearly no clique subgraph of $C^{-1}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. ∎
Lemma 44 (Pruning cases for $H^{-1}$).
The expanded $f$-component $H^{-1}$ of any expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair is one of the graphs in Figures 15(a) - 15(m), where $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), 15(e), 15(i), 15(k), or 15(l), then at least one dashed line-segment is a nonedge.
Proof.
By Proposition 7, $H^{-1}$ is one of the graphs in Figure 15. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$ and consider the nonedge $f^{\prime}$ of $C^{-1}$ and the atom $J$ of $G^{\prime}\cup f^{\prime}$ given by Lemma 43. For each case of $H^{-1}$ in Figures 15(n) - 15(r) or Figures 15(c), 15(d), 15(i), 15(k), or 15(l) such that every dashed line-segment is an edge, observe that there exists a minor $[G\cup f]$ obtained via a single contraction such that $f$ is retained in $[G\cup f]$, $[f]=[f^{\prime}]$, and some induced subgraph of $[G\cup f]$ is $G^{\prime}\cup f^{\prime}$ such that $[f]$ is $f^{\prime}$. These facts along with the existence of $J$ show that the edge contracted in $[G\cup f]$ is reducing with respect to $(G,f)$, which contradicts the minimality of this pair. ∎
Next, we prove Lemma 47, below, using Lemmas 45 and 46, also below.
Lemma 45 (Properties of graphs with $H^{-1}$ in Figures 15(a) - 15(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 15(a) - 15(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. If $G_{2}\cup e$ is an atom and no clique subgraph of $S$ separates the endpoints of $e$ in $G_{2}$, then $G$ is an atom.
Proof.
The lemma will follow by applying Lemma 65 after we show that Statements (1)-(4) of this lemma are satisfied. Statements (1) and (4) are true by assumption. For each case of $G_{1}$, it is easy to verify that either $G_{1}$ or $G_{1}\cup e$ is an atom, and so Statement (2) is satisfied. Also, for any clique subgraph $A$ of $S$, $G_{1}\setminus A$ is connected, and so Statement (3) is satisfied. ∎
Lemma 46 (Additional properties of graphs with $H^{-1}$ in Figures 15(a) - 15(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 15(a) - 15(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. Also, let $e=w_{1}w_{3}$ if $G_{1}$ is the graph in Figure 15(c). If $G_{2}\cup e$ has an $e$-preserving forbidden minor, then the red edge of $G_{1}$ is preserved in some forbidden minor of $G$.
Proof.
Assume that $G_{2}\cup e$ has an $e$-preserving forbidden minor. We will show that $G$ has an $f$-preserving forbidden minor, where $f$ is the red edge of $G_{1}$. If there exists a minor $[G]$ in which $G_{2}$ is preserved and $[f]=[e]$, then this follows from the fact that $G_{2}\cup e$ has an $e$-preserving forbidden minor. Using this fact, it is easy to see that the lemma is true if either (i) $G_{1}$ is one of the graphs in Figures 15(e), 15(f), 15(k), 15(l), or 15(m), (ii) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 15(d) or 15(i), or (iii) $e=w_{1}w_{4}$ and $G_{1}$ is one of the graphs in Figures 15(g) or 15(h).
Next, consider the case where $G_{1}$ is the graph in Figure 15(a). Note that either $w_{1}w_{2}$ or $w_{2}w_{3}$ is preserved in any $e$-preserving minor of $G_{2}$. Also, observe that there exists a minor $[G]$ in which $G_{2}$ is preserved, $[e]$ is an edge, and $[f]$ is either $[w_{1}w_{2}]$ or $[w_{2}w_{3}]$. Combining these facts shows that $G$ has an $f$-preserving forbidden minor. A similar argument applies if either (i) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 15(b), 15(c), 15(g), 15(h), or 15(j) or (ii) $e=w_{2}w_{4}$ and $G_{1}$ is one of the graphs in Figures 15(d), 15(i), or 15(j).
The above cases are exhaustive, and so the proof is complete. ∎
Lemma 47 (Additional structure necessary for induction).
Let $(G,f)$ be a minimal graph-nonedge pair with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. There exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. Furthermore, these statements are true when (i) $f^{\prime}=w_{1}w_{3}$ and $H^{-1}$ is the graph in Figure 15(c), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$, or (ii) $f^{\prime}$ is any nonedge of $C^{-1}$ and $H^{-1}$ is one of the graphs in Figures 15(d), 15(i), or 15(j), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
Proof.
For any nonedge $f^{\prime}$ of $C^{-1}$, consider the statements (a) some atom of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and (b) the pair $(G^{\prime},f^{\prime})$ has no reducing edge. By Lemma 43, (a) is true for some choice of $f^{\prime}$, and for any choice of $f^{\prime}$ if $C^{-1}$ is the graph in Figure 13(b). Assume that none of these choices satisfy (b). We will show that $(G,f)$ has a reducing edge, which, when combined with the assumption that $(G,f)$ is minimal, proves the lemma.
Since any choice for $f^{\prime}$ that satisfies (a) does not satisfy (b), the pair $(G^{\prime},f^{\prime})$ has some reducing edge $h$. Consider the minor $[G\cup f]$ obtained by contracting $h$. By definition, $f^{\prime}$ is retained in $[G^{\prime}]$ and some induced subgraph $K$ of $[G^{\prime}]$ is such that $K\cup[f^{\prime}]$ is an atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. Let $S=[C^{-1}]$, $e=[f^{\prime}]$, $G_{1}=[H^{-1}]$, and $G_{2}=K$. By Lemma 44, $H^{-1}$ is one of the graphs in Figures 15(a) - 15(m) and $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. There are two cases.
Case 1: $G_{1}$ is preserved in $[G\cup f]$.
We will show that $h$ is reducing with respect to $(G,f)$. Observe that some endpoint of $h$ is contained in $G^{\prime}\setminus C^{-1}$. Hence, $f$ is retained in $[G\cup f]$, and so $h$ is not of Type (3). To see that $h$ is not of Type (1) or (2), we show $G_{1}\cup G_{2}$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. By construction, $G_{1}\cup G_{2}$ contains $[f]$. If $C^{-1}$ is the graph in Figure 13(b), then Lemma 43 allows us to choose any of its nonedges to be $f^{\prime}$. Hence, Lemma 46 states that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor.
We complete the proof by showing that $G_{1}\cup G_{2}$ is an atom. If $G_{1}$ is not the graph in Figure 15(a), then it is easy to verify that $G_{1}$ is an atom. Since $G_{2}\cup e$ is an atom, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Hence, assume that $G_{1}$ is the graph in Figure 15(a). We will demonstrate that no clique subgraph of $S$ separates the endpoints of $e$ in $G_{2}$ and then apply Lemma 45 to show that $G_{1}\cup G_{2}$ is an atom. The only clique subgraph to check is the one induced by $\{[w_{2}]\}$. Since $G_{2}\cup e$ has an $e$-preserving forbidden minor and $S$ contains exactly three vertices, there exists some vertex $x$ in $G_{2}\setminus S$. Let $e=[uv]$. If $\{[w_{2}]\}$ separates $[u]$ and $[v]$ in $G_{2}$, then either every path in $G_{2}$ between $x$ and $[u]$ contains some vertex in $\{[v],[w_{2}]\}$ or every path in $G_{2}$ between $x$ and $[v]$ contains some vertex in $\{[u],[w_{2}]\}$. This implies that either $\{[u],[w_{2}]\}$ or $\{[v],[w_{2}]\}$ is a clique separator of $G_{2}\cup e$. Therefore, since $G_{2}\cup e$ is an atom, this implies that $\{[w_{2}]\}$ does not separate $[u]$ and $[v]$ in $G_{2}$.
Case 2: $G_{1}$ is not preserved in $[G\cup f]$.
Since $h$ is contained in $G^{\prime}$, it must be contained in $S$. Combining this with the fact that $e$ is retained in $[G\cup f]$ shows that $H^{-1}$ is one of the graphs in Figures 15(b), 15(f), 15(g), 15(h), 15(k), 15(l), or 15(m). There are $9$ subcases.
Subcase 1: $H^{-1}$ is the graph in Figure 15(b).
Note that $h=w_{2}w_{4}$ and $G_{1}$ is the graph in Figure 15(a). We will show that $h$ is reducing with respect to $(G,f)$. Observe that $f$ is retained in $[G\cup f]$, and so $h$ is not of Type (3). To see that $h$ is not of Type (1) or (2), we show $G_{1}\cup G_{2}$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. By construction, $G_{1}\cup G_{2}$ contains $[f]$. Also, Lemma 46 shows that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor. We will demonstrate that no clique subgraph of $S$ separates the endpoints of $e$ in $G_{2}$ and then apply Lemma 45 to show that $G_{1}\cup G_{2}$ is an atom. The only clique subgraph to check is the one induced by $\{[w_{2}]\}$. Since $f^{\prime}$ satisfies (a), there exists an atom $J$ of $G^{\prime}\cup f^{\prime}$ that contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor. Lemma 43 shows that no clique subgraph of $C^{-1}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. If $\{[w_{2}]\}$ separates the endpoints of $e$ in $G_{2}$, then $\{w_{2},w_{4}\}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. Combining these statements implies that $\{[w_{2}]\}$ does not separate the endpoints of $e$ in $G_{2}$.
Subcase 2: $H^{-1}$ is one of the graphs in Figures 15(f), 15(k), 15(l), or 15(m).
If either $H^{-1}$ is not the graph in Figure 15(f) or $h\neq w_{2}w_{4}$, then we show that $h$ is reducing with respect to $(G,f)$. Observe that $f$ is retained in $[G\cup f]$, and so $h$ is not of Type (3). To see that $h$ is not of Type (1) or (2), we show $G_{1}\cup G_{2}$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. By construction, $G_{1}\cup G_{2}$ contains $[f]$. Also, note that $G_{1}$ is an atom. Hence, since $G_{2}\cup e$ is an atom, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Lastly, since $G_{2}\cup e$ has an $e$-preserving forbidden minor, $G_{1}\cup G_{2}$ clearly has an $[f]$-preserving forbidden minor.
Next, we show that if $H^{-1}$ is the graph in Figure 15(f), then $h\neq w_{2}w_{4}$. If this is not the case, then $f^{\prime}=w_{1}w_{3}$ since it is retained in $[G\cup f]$. Observe that $[G\cup f]$ is $f$-retaining and has an $[f]$-preserving forbidden minor, since it has an $e$-preserving forbidden minor. Let $x$ be the vertex in $H^{-1}\setminus C^{-1}$ that is not an endpoint of $f$. Note that the subgraph $E$ of $[G\cup f]$ induced by $\{[x],[w_{1}],[w_{2}]\}$ is a cms of $[G\cup f]$ such that the $E$-component of $[G\cup f]$ containing $[f]$ is a proper subgraph of $G_{1}$ that has no $[f]$-preserving forbidden minor. These facts imply that $E^{-1}$ is an expanded $f$-separating cms whose expanded $f$-component is a proper subgraph of $H^{-1}$. However, this contradicts the fact that $C^{-1}$ is in the top-level of $(G,f)$. Therefore, $h$ is not $w_{2}w_{4}$.
Subcase 3: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, and $h$ is $w_{2}w_{3}$ or $w_{2}w_{4}$.
Arguments similar to those in Subcases 1 and 2 show that $h$ is reducing with respect to $(G,f)$.
Subcase 4: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{3}$.
Let $u$ be the endpoint of $f$ not contained in $C^{-1}$. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{A}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{A}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
Recall the induced subgraph $K$ of $[G^{\prime}]$ such that $K\cup[f^{\prime}]$ is an atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. Note that the induced subgraph $K^{-1}\cup f^{\prime}$ of $G^{\prime}\cup f^{\prime}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. If $K^{-1}\cup f^{\prime}$ is an atom, then it is contained in some atom $X$ of $G^{\prime}\cup f^{\prime}$ since it is an induced subgraph. Observe that each cms of $X\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Hence, using Lemma 27, we see that some atom of $X\cup w_{1}w_{4}$, and therefore some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor, as desired.
Otherwise, if $K^{-1}\cup f^{\prime}$ is not an atom, then it contains exactly one cms $E$, since $K\cup[f^{\prime}]$ is an atom. Furthermore, $E$ contains exactly one vertex in $\{w_{3},w_{4}\}$ and there are exactly two $E$-components of $K^{-1}\cup f^{\prime}$, both of which are atoms and one of which is $E\cup\{w_{3},w_{4}\}$. These facts imply that the atom graph of $K^{-1}\cup f^{\prime}$ is a path of length two. Let $x$ be the vertex in $\{w_{3},w_{4}\}$ that is not contained in $E$. The neighborhood of $x$ in $K^{-1}\cup f^{\prime}$ is $V(E)$, which is a clique. Since $K^{-1}\cup f^{\prime}$ contains the vertices $w_{1}$, $w_{3}$, and $w_{4}$, the edges $w_{1}w_{3}$ and $w_{3}w_{4}$, and the nonedge $w_{1}w_{4}$, this implies that $x=w_{4}$ and $E$ does not contain $w_{1}$. Therefore, we can use Lemma 59 to see that $K^{-1}\cup\{f^{\prime},w_{1}w_{4}\}$ is an atom. Consequently, it is contained in some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ that contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor, as desired.
Subcase 5: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{4}$.
This case is similar to Subcase 4. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{A}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{A}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
If $K^{-1}\cup f^{\prime}$ is an atom, then it is contained in some atom $X$ of $G^{\prime}\cup f^{\prime}$, since it is an induced subgraph of $G^{\prime}\cup f^{\prime}$. Note that $X$ has a forbidden minor in which $w_{3}w_{4}$ is contracted and both $w_{1}w_{3}$ and $w_{1}w_{4}$ are preserved. Hence, if $X\cup w_{1}w_{3}$ is an atom, then it is contained in some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ that has the desired properties. Otherwise, each cms of $X\cup w_{1}w_{3}$ contains $w_{1}w_{3}$. Hence, Lemma 27 shows that some atom $Y$ of $X\cup w_{1}w_{3}$ contains $w_{1}w_{3}$ and has a forbidden minor in which $w_{3}w_{4}$ is contracted and both $w_{1}w_{3}$ and $w_{1}w_{4}$ are preserved. If $Y$ contains $w_{4}$, then the atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ that contains $Y$ has the desired properties. Otherwise, $Y$ is an atom of $(X\cup w_{1}w_{3})\setminus w_{1}w_{4}$. Therefore, $w_{1}w_{3}$ satisfies (a), and so we can instead choose $f^{\prime}=w_{1}w_{3}$ and apply Subcase 4.
Next, if $K^{-1}\cup f^{\prime}$ is not an atom, then an argument similar to the one in Subcase 4 shows that it contains exactly one cms $E$ that contains $w_{4}$ but not $w_{3}$ and such that there are exactly two $E$-components of $K^{-1}\cup f^{\prime}$, both of which are atoms and one of which is $E\cup w_{3}$. These facts imply that the atom graph of $K^{-1}\cup f^{\prime}$ is a path of length two. Since the neighborhood of $w_{3}$ in $K^{-1}\cup f^{\prime}$ is $E$, which is a clique, and $w_{1}w_{3}$ is a nonedge, $E$ does not contain $w_{1}$. Therefore, we can use Lemma 59 to see that $K^{-1}\cup\{f^{\prime},w_{1}w_{3}\}$ is an atom, and so it is contained in some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ that has the desired properties.
Subcase 6: $H^{-1}$ is the graph in Figure 15(h), $uw_{1}$ is a nonedge, and $h=w_{2}w_{3}$.
An similar argument to that in Subcase 2 applies.
Subcase 7: $H^{-1}$ is the graph in Figure 15(h), $uw_{1}$ is a nonedge, and $h=w_{2}w_{4}$.
We will show that $h$ is reducing with respect to $(G,f)$. Observe that $f$ is retained in $[G\cup f]$, and so $h$ is not of Type (3). To see that $h$ is not of Type (1) or (2), we show $G_{1}\cup G_{2}$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor. By construction, $G_{1}\cup G_{2}$ contains $[f]$. Also, since $G_{2}\cup e$ has an $e$-preserving forbidden minor, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor. We will show that $G_{1}\cup G_{2}$ is an atom by applying Lemma 65, after we show that Statements (1) - (4) in this lemma are true. Statement (1) follows from our assumption that $G_{2}\cup e$ is an atom. Statement (2) follows from the observation that $G_{1}\cup e$ is an atom. For Statement (3) and (4), consider any clique subgraph $A$ of $S$. It is easy to see that $G_{1}\setminus A$ is connected, and so Statement (3) is true. Lastly, to show that Statement (4) is true, we must argue that $\{[w_{2}]\}$ does not separate the endpoints of $e$ in $G_{2}$. Since $f^{\prime}$ satisfies (a), there exists an atom $J$ of $G^{\prime}\cup f^{\prime}$ that contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor. Lemma 43 shows that no clique subgraph of $C^{-1}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. If $\{[w_{2}]\}$ separates the endpoints of $e$ in $G_{2}$, then $\{w_{2},w_{4}\}$ separates the endpoints of $f^{\prime}$ in $J\setminus f^{\prime}$. Combining these statements implies that $\{[w_{2}]\}$ does not separate the endpoints of $e$ in $G_{2}$. Therefore, applying Lemma 3 completes the proof.
Subcase 8: $H^{-1}$ is the graph in Figure 15(h), $uw_{1}$ is a nonedge, $h=w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{4}$.
Let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ other that an endpoint of $f$. We will show that $xw_{1}$ is reducing with respect to $(G,f)$. Observe that the minor $[G\cup f]_{A}$ obtained by contracting $xw_{1}$ is $f$-retaining, and so $xw_{1}$ is not of Type (3). We show that $xw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor.
Since $f^{\prime}$ satisfies (a), some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor. We will show that either $J$ contains $w_{3}$ or $w_{2}w_{3}$ is reducing with respect to $(G^{\prime},f^{\prime})$. In the latter case, Subcase 6 applies. If $J$ does not contain $w_{3}$, then note that the minor $[G\cup f^{\prime}]_{B}$ obtained by contracting $w_{2}w_{3}$ is $f^{\prime}$-retaining, so it $w_{2}w_{3}$ is not of Type (3). Also, some cms $E$ contained in $J$ is such that all paths in $G^{\prime}\cup f^{\prime}$ between $w_{3}$ and any vertex of $J\setminus E$ contain some vertex of $E$. Hence, $J$ is weakly retained in $[G^{\prime}\cup f^{\prime}]_{B}$. Therefore, some atom of $[G^{\prime}\cup f^{\prime}]_{B}$ contains $[J]_{B}$, and consequently contains $[f^{\prime}]_{B}$ and has an $[f^{\prime}]_{B}$-preserving forbidden minor. This shows that $w_{2}w_{3}$ is not of Type (1) or (2), and so it is reducing with respect to $(G^{\prime},f^{\prime})$.
Next, since $J$ contains $w_{3}$, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]_{A}$ such that $f^{\prime}$ is $[f^{\prime}]_{A}$. Let $G_{1}=[C^{-1}]_{A}\cup[u]_{A}$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{A}$. Then, $G_{1}\cup G_{2}$ contains $[f]_{A}$. Also, since $G_{2}$ has an $[f^{\prime}]_{A}$-preserving forbidden minor, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{A}$-preserving forbidden minor. Furthermore, since $G_{1}$ and $G_{2}$ are atoms and $e$ is a nonedge, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Furthermore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{A}$ since it is an induced subgraph of $[G\cup f]_{A}$, and this atom has the desired properties.
Subcase 9: $H^{-1}$ is the graph in Figure 15(h), $uw_{1}$ is a nonedge, $h=w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{3}$.
First, we show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}$, $w_{3}$, and $w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. Recall the induced subgraph $K$ of $[G^{\prime}]$ such that $K\cup[f^{\prime}]$ is an atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. Note that the induced subgraph $K^{-1}\cup f^{\prime}$ of $G^{\prime}\cup f^{\prime}$ contains $w_{1}$, $w_{3}$, and $w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. If $K^{-1}\cup f^{\prime}$ is an atom, then it is contained in some atom $X$ of $G^{\prime}\cup f^{\prime}$ since it is an induced subgraph. Observe that each cms of $X\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Hence, Lemma 27 shows that some atom $Y$ of $X\cup w_{1}w_{4}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. If $Y$ contains $w_{3}$, then we are done. Otherwise, some atom of $G^{\prime}\cup w_{1}w_{4}$ contains $Y$. Therefore, $w_{1}w_{4}$ satisfies (a), and so we can instead choose $f^{\prime}=w_{1}w_{4}$ and apply Subcase 8.
Next, if $K^{-1}\cup f^{\prime}$ is not an atom, then an argument similar to that in Subcase 4 shows that $K^{-1}\cup f^{\prime}$ contains exactly one cms $E$ that contains $w_{3}$ but not $w_{4}$ and such that there are exactly two $E$-components of $K^{-1}\cup f^{\prime}$, both of which are atoms and one of which is $E\cup w_{4}$. These facts imply that the atom graph of $K^{-1}\cup f^{\prime}$ is a path of length two. Since the neighborhood of $w_{4}$ in $K^{-1}\cup f^{\prime}$ is $E$, which is a clique, and $w_{1}w_{4}$ is a nonedge, $E$ does not contain $w_{1}$. Therefore, we can use Lemma 59 to see that $K^{-1}\cup\{f^{\prime},w_{1}w_{4}\}$ is an atom, and so it is contained in some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ that has the desired properties.
Finally, by the above arguments, some induced subgraph $I$ of $[G^{\prime}]_{A}$ is such that $I\cup[w_{1}w_{3}]_{A}$ is an atom that contains $[w_{1}]_{A}$, $[w_{3}]_{A}$, and $[w_{4}]_{A}$ and has a $[w_{1}w_{4}]_{A}$-preserving forbidden minor. Let $G_{1}=[C^{-1}]_{A}\cup[u]_{A}$, $G_{2}=I$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{A}$. Similar to the argument in Subcase 8, we can use Lemma 64 to show that some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor.
The above cases are exhaustive, and so the lemma is proved. ∎
Finally, we prove Lemma 48, below, and then prove Proposition 8.
Lemma 48 (Structure of $G^{\prime}\cup f^{\prime}$).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. If $f^{\prime}=uv$ is a nonedge of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then $G^{\prime}\cup f^{\prime}$ can be obtained from $J$ by adding some vertices whose neighborhoods are $\{u,v\}$.
Proof.
Assume to the contrary that $G^{\prime}\cup f^{\prime}$ cannot be obtained from $J$ by adding vertices whose neighborhood are $\{u,v\}$. Then, there exists a cms $E$ of $G^{\prime}\cup f^{\prime}$ contained in $J$ such that $E$ contains some vertex $x$ that is either the only vertex in $E$ or some vertex other than $u$ or $v$. Also, there exists a vertex $y$ not contained in $J$ and such that $xy$ is an edge of $G^{\prime}\cup f^{\prime}$ and all paths in $G^{\prime}\cup f^{\prime}$ between $y$ and any vertex in $J\setminus E$ contains some vertex in $E$. Hence, $J$ is weakly retained in the minor $[G^{\prime}\cup f^{\prime}]$ obtained by contracting $xy$. Furthermore, $[G^{\prime}\cup f^{\prime}]$ is $f^{\prime}$-retaining and contains some atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. The above facts show that $xy$ is reducing with respect to $(G^{\prime},f^{\prime})$, which is a contradiction. This completes the proof. ∎
Proof of Proposition 8.
The proposition follows easily from Lemmas 47 and 48. ∎
C.4. Proof of Proposition 9 [Pruned possible graphs for top-level $H^{-1}$]
Lemma 49 (Vertices in atom containing $f^{\prime}$ 1).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Also, let $C^{-1}$ be the graph in Figure 13(b). For any nonedge $f^{\prime}$ of $C^{-1}$, if some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then $J$ contains $w_{1}$, $w_{3}$, and $w_{4}$.
Proof.
Wlog, assume $f^{\prime}=w_{1}w_{3}$. Then, $J$ contains $w_{1}$ and $w_{3}$. If $J$ does not contain $w_{4}$, then Lemma 48 states that the neighborhood of $w_{4}$ in $G^{\prime}$ is $\{w_{1},w_{3}\}$. However, Lemma 4 Statement (2) can be used to show that $w_{4}$ neighbors some vertex of $G^{\prime}$ not in $\{w_{1},w_{3}\}$. This is a contradiction, and thus $J$ contains $w_{4}$. ∎
Lemma 50 (Vertices in atom containing $f^{\prime}$ 2).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Also, let $C^{-1}$ be the graph in Figure 13(d). For any nonedge $f^{\prime}$ of $C^{-1}$, if some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then $J$ contains $C^{-1}$.
Proof.
Wlog, assume that $f^{\prime}=w_{1}w_{3}$. By assumption, $J$ contains $w_{1}$ and $w_{3}$. If $J$ does not contain some vertex in $\{w_{2},w_{4}\}$, then Lemma 48 shows that the neighborhood of this vertex in $G^{\prime}$ is $\{w_{1},w_{3}\}$. However, observe that $w_{2}w_{4}$ is an edge of $G^{\prime}$. Thus, $J$ contains $C^{-1}$. ∎
Lemma 51.
Additional pruning of cases for $H^{-1}$ For any minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$, the expanded $f$-component of $C^{-1}$ is one of the graphs in Figures 15(a) - 15(h), where $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), or 15(e), then the dashed line-segment is a nonedge.
Proof.
By Lemma 44, the expanded $f$-component $H^{-1}$ of $C^{-1}$ is one of the graphs in Figures 15(a) - 15(m), where $C^{-1}$ is the subgraph induced by $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), 15(e), 15(i), 15(k), or 15(l), then at least one dashed line-segment is a nonedge. For each case of $H^{-1}$ in Figures 15(i) - 15(m), we will show that $(G,f)$ has a reducing edge, which contradicts the minimality of this pair.
Case 1: $H^{-1}$ is one of the graphs Figures 15(i) or 15(j).
We prove the lemma in the case where $H^{-1}$ is the graph Figure 15(i). The proof in the case where $H^{-1}$ is the graph Figure 15(j) is almost identical, so we omit it. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. By Lemma 47, some atom $J$ of $G^{\prime}\cup w_{1}w_{3}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor and the pair $(G^{\prime},w_{1}w_{3})$ has no reducing edge. Also, Lemma 49 shows that $J$ contains $w_{4}$. If $J$ contains $w_{2}$, then let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ that neighbors $w_{3}$. We will show that $xw_{3}$ is reducing with respect to $(G,f)$. Note that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, and so this edge is not of Type (3). To see that $xw_{3}$ is not of Type (1) or (2), observe that $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]$ such that $w_{1}w_{3}$ is $[w_{1}x]$. Hence, $G_{2}$ is an atom that contains $S=[C^{-1}]$ and has a $[w_{1}x]$-preserving forbidden minor. Let $G_{1}$ be the $\{[u],[w_{1}],[w_{3}]\}$-component of $[H^{-1}]$ that contains $S$. Then, $G_{1}$ is an atom and $S=G_{1}\cap G_{2}$. Since $S$ contains the nonedge $[w_{2}w_{4}]$, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Lastly, $G_{1}\cup G_{2}$ clearly has an $[f]$-preserving forbidden minor. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor.
Next, if $J$ does not contain $w_{2}$, then let $x$ be the vertex of $H^{-1}\setminus C^{-1}$ that neighbors $w_{2}$. We will show that $xw_{2}$ is reducing with respect to $(G,f)$. Note that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{2}$, and so this edge is not of Type (3). To see that $xw_{2}$ is not of Type (1) or (2), we first show that the graph $K=((J\cup w_{2})\setminus w_{1}w_{3})\cup w_{2}w_{4}$ is an atom that has a $w_{1}w_{3}$-preserving forbidden minor. The latter claim follows easily by considering the minor of $K$ obtained by contracting $w_{1}w_{2}$. The former claim will follow by applying Lemma 65. Let $G_{1}$ and $G_{2}$ be the subgraphs of $K$ induced by $\{w_{1},w_{2},w_{3},w_{4}\}$ and $V(J)$, respectively. Observe that $K=G_{1}\cup G_{2}$ and the $S=G_{1}\cap G_{2}$ is the subgraph of $K$ induced by $\{w_{1},w_{3},w_{4}\}$ that contains the nonedge $w_{1}w_{3}$. Furthermore, both $G_{1}\cup w_{1}w_{3}$ and $G_{2}\cup w_{1}w_{3}$ are atoms and $G_{1}\setminus A$ is connected for any clique subgraph of $S$. Lastly, we must show that $w_{4}$ does not separate $w_{1}$ and $w_{3}$ in $G_{2}$. Since $G_{2}\cup w_{1}w_{3}$ has a forbidden minor and $S$ contains exactly three vertices, there exists some vertex in $G_{2}\setminus S$. Additionally, since $G_{2}\cup w_{1}w_{3}$ is an atom, there exist paths that connect this vertex to $w_{1}$ and $w_{3}$ that each contain exactly one vertex in $S$. Therefore, $w_{4}$ does not separate $w_{1}$ and $w_{3}$ in $G_{2}$, and so Lemma 65 shows that $K$ is an atom.
Finally, observe that $K$ is some induced subgraph $G_{1}$ of $[G^{\prime}]$ such that $w_{2}w_{4}$ is $[xw_{4}]$. Let $G_{2}=[H^{-1}]$ and observe that $S=G_{1}\cap G_{2}=[C^{-1}]$ contains the nonedge $[w_{1}w_{3}]$. Since $G_{1}$ and $G_{2}$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Observe that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor. These facts show that the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Therefore, $xw_{2}$ is not of Type (1) or (2), as desired.
Case 2: $H^{-1}$ is one of the graphs in Figures 15(k), 15(l), or 15(m).
By Lemma 47, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. Furthermore, Lemma 50 shows that $J$ contains $C^{-1}$. The remainder of the argument proceeds similarly to Case 1, so we omit it.
The above cases are exhaustive, and so the proof is complete. ∎
Lemma 52 ($H^{-1}$ is not graph in Figure 15(f)).
The graph in Figure 15(f) is not the expanded $f$-component of any top-level expanded $f$-separating cms of a minimal graph-nonedge pair.
Proof.
Assume to the contrary that the graph in Figure 15(f) is the expanded $f$-component $H^{-1}$ of some top-level expanded $f$-separating cms $C^{-1}$ of some minimal graph-nonedge pair $(G,f)$. We will show that $uw_{4}$ is reducing with respect to $(G,f)$, which contradicts the fact that $(G,f)$ is minimal. Observe that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $uw_{4}$, and so $uw_{4}$ is not of Type (3). We show that $uw_{4}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. By Lemma 47, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. If $f^{\prime}=w_{1}w_{4}$, then some atom of $[G\cup f]$ clearly has the desired properties.
In the case where $f^{\prime}=w_{1}w_{3}$, we first argue that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 50, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{A}$. If $w_{1}w_{4}$ is preserved in $[K]_{A}$, then the claim is true. Else, we define a vertex $x$ as follows. If $K$ contains $w_{3}$, then set $x=w_{3}$. Otherwise, some cms $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{3}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Since $[K]_{A}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $E\setminus\{w_{1},w_{4}\}$ contains some vertex $x$ such that $[x]_{A}=[w_{3}]_{A}$.
Next, for either case of $x$, there clearly exists a $w_{1}x$-preserving minor $[K]_{B}$ from which $[K]_{A}$ can be obtained by contracting $[w_{1}w_{4}]_{B}$. Hence, $[K]_{B}$ is one of the graphs in Figures 8, 17, or 18 such that $[w_{1}w_{4}]_{B}$ is the green edge. Since $K$ is an atom, it is $2$-connected, and so Lemma 10 allows us to assume that $[K]_{B}$ is one of the graphs in Figures 8 or 18. Furthermore, since $[K]_{A}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $[w_{3}]_{B}$ is neither $[w_{1}]_{B}$ nor $[w_{4}]_{B}$. Combining this with the fact that $w_{3}$ neighbors both $w_{1}$ and $w_{4}$ shows that $[K]_{B}$ cannot be any graph in Figure 8. Therefore, Lemma 11 shows that $[K]_{B}$ has a $[w_{1}w_{4}]_{B}$-preserving forbidden minor, and hence $K$ has a $w_{1}w_{4}$-preserving forbidden minor.
Finally, we show that there exists some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Note that $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]$ such that $w_{1}w_{4}$ is $[w_{1}w_{4}]$. If $K$ does not contain $w_{3}$, then $N$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor, since $[f]=[w_{1}w_{4}]$. Hence, the atom of $[G\cup f]$ that contains $N$ has the desired properties. Otherwise, if $K$ contains $w_{3}$, then let $G_{1}=[H^{-1}]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ has the desired properties. ∎
Lemma 53 ($H^{-1}$ is not graph in Figures 15(g) or 15(h)).
Neither of the graphs in Figures 15(g) or 15(h) such that $uw_{1}$ is an edge is the expanded $f$-component of any top-level expanded $f$-separating cms of a minimal graph-nonedge pair.
Proof.
Assume to the contrary that one of the graphs in Figures 15(g) or 15(h) such that $uw_{1}$ is an edge is the expanded $f$-component $H^{-1}$ of some top-level expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair $(G,f)$. We will show that either $uw_{1}$ or $w_{2}w_{4}$ is reducing with respect to $(G,f)$, which contradicts the fact that $(G,f)$ is minimal. Observe that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $uw_{1}$, and so $uw_{1}$ is not of Type (3). We attempt to show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 47, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. If $f=w_{1}w_{3}$, then the argument is identical to the proof of Lemma 52, so we omit the details.
Next, consider the case where $f^{\prime}=w_{1}w_{4}$. By Lemma 50, $J$ contains $w_{3}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{3}$ contains $w_{1}w_{3}$. Combining this with the fact that $J\cup w_{1}w_{3}$ has a $w_{1}w_{4}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{3}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{3}$ and has a $w_{1}w_{4}$-preserving forbidden minor $[K]_{A}$. If $K$ contains $w_{4}$, then the claim is true. Otherwise, we show that $w_{2}w_{4}$ is reducing with respect to $(G,f)$. First, we argue that $K$ has a $w_{1}w_{3}$-preserving forbidden minor. If $[K]_{A}$ preserves $w_{1}w_{3}$, then we are done. Else, since $K$ does not contain $w_{4}$, some cms $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{4}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Also, since $[K]_{A}$ preserves $w_{1}w_{4}$ but not $w_{1}w_{3}$, $E\setminus\{w_{1},w_{3}\}$ contains some vertex $x$ such that $[x]_{A}=[w_{4}]_{A}$. Therefore, there clearly exists a $w_{1}x$-preserving forbidden minor $[K]_{B}$ from which $[K]_{A}$ can be obtained by contracting $[w_{1}w_{3}]_{B}$. The remainder of the proof that $K$ has a $w_{1}w_{3}$-preserving forbidden minor is identical to part of the proof of Lemma 52, so we omit the details.
Finally, observe that $f$ is retained in the minor $[G\cup f]_{D}$ obtained by contracting $w_{2}w_{4}$, and so $w_{2}w_{4}$ is not of Type (3). We complete the proof by showing that $w_{2}w_{4}$ is not of Type (1) or (2). Since $K$ does not contain $w_{4}$, observe that $K$ is an atom of $G\cup w_{1}w_{3}$ and is weakly retained in the minor of $G\cup w_{1}w_{3}$ obtained by contracting $w_{2}w_{4}$. Hence, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]_{D}$ such that each vertex $w_{i}$ other than $w_{4}$ is $[w_{i}]_{D}$. Let $G_{1}$ be the subgraph of $[G\cup f]_{D}$ induced by $\{[u]_{D},[w_{1}]_{D},[w_{2}]_{D},[w_{3}]_{D}\}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{D}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{D}$. Also, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{D}$-preserving forbidden minor. We will apply Lemma 65 to show that $G_{1}\cup G_{2}$ is an atom. Note that $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}\cup e$ and $G_{2}\cup e$ are atoms. Furthermore, $G_{1}\setminus A$ is connected for any clique subgraph $A$ of $S$. Lastly, the only clique subgraph of $S$ that could separate $[w_{1}]_{D}$ and $[w_{3}]_{D}$ in $G_{2}$ is $[w_{2}]_{D}$. Since $G_{2}\cup e$ has a forbidden minor and $S$ contains at most three vertices, there exists some vertex in $G_{2}\setminus S$. Additionally, since $G_{2}\cup e$ is an atom, there exist paths that connect this vertex to $[w_{1}]_{D}$ and $[w_{3}]_{D}$ that each contain exactly one vertex in $S$. Therefore, $[w_{2}]_{D}$ does not separate $[w_{1}]_{D}$ and $[w_{3}]_{D}$ in $G_{2}$, and so Lemma 65 shows that $K$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{D}$ that contains $[f]_{D}$ and has an $[f]_{D}$-preserving forbidden minor, and the proof is complete. ∎
Lemma 54 ($H^{-1}$ is not graph in Figure 15(h)).
The graph in Figure 15(h) such that $uw_{1}$ is a nonedge is not the expanded $f$-component of any top-level expanded $f$-separating cms of a minimal graph-nonedge pair.
Proof.
Assume to the contrary that the graph in Figure 15(h) such that $uw_{1}$ is a nonedge is the expanded $f$-component $H^{-1}$ of some top-level expanded $f$-separating cms $C^{-1}$ of some minimal graph-nonedge pair $(G,f)$. Let $x$ be any vertex of $H^{-1}\setminus(C^{-1}\cup u)$. We show that either $xw_{1}$ or $w_{2}w_{3}$ is reducing with respect to $(G,f)$, which contradicts the fact that $(G,f)$ is minimal. Observe that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{1}$, and so $xw_{1}$ is not of Type (3). We attempt to show that $xw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. By Lemma 47, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. There are two cases.
Case 1: $f^{\prime}=w_{1}w_{3}$.
First, we show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor. By Lemma 50, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{A}$. If $K$ contains $w_{3}$, then we are done. Otherwise, we show that $w_{2}w_{3}$ is reducing with respect to $(G,f)$.
Observe that $f$ is retained in the minor $[G\cup f]_{A}$ obtained by contracting $w_{2}w_{3}$, and so $w_{2}w_{3}$ is not of Type (3). To see that $w_{2}w_{3}$ is not of Type (1) or (2), note that an argument similar to to part of the proof of Lemma 53 shows that $K$ has a $w_{1}w_{4}$-preserving forbidden minor. Also, since $K$ does not contain $w_{3}$, $K$ is an atom of $G\cup w_{1}w_{4}$ and is weakly retained in the minor of $G\cup w_{1}w_{4}$ obtained by contracting $w_{2}w_{3}$. Hence, $K\setminus w_{1}w_{4}$ is some induced subgraph $N$ of $[G^{\prime}]_{A}$ such that each vertex $w_{i}$ other than $w_{3}$ is $[w_{i}]_{A}$. Let $G_{1}=[H^{-1}_{A}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{4}]_{A}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{A}$. Furthermore, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{A}$-preserving forbidden minor. Lastly, since $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor, and the proof is complete.
Finally, since $K$ contains $w_{3}$, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G\cup f]$ such that each vertex $w_{i}$ is $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor, as desired.
Case 2: $f^{\prime}=w_{1}w_{4}$.
By Lemma 50, $J$ contains $w_{2}$ and $w_{3}$. Hence, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}\cup w_{1}w_{4}]$ such that each vertex $w_{i}$ is $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Also, since $e$ is a nonedge and both $G_{1}$ and $G_{2}$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in an atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor. Thus, $xw_{1}$ is reducing, as desired.
The above cases are exhaustive, and so the proof is complete. ∎
Proof of Proposition 9.
The proposition follows immediately from Lemmas 51, 52, 53, and 54. ∎
C.5. Proof of Proposition 10 [Extending non-$3$-SIP maps for minimal pairs]
Lemma 55 (Degree $3$ decorations preserve non-$3$-SIP).
Let $G$ be $K_{4}$ and let $H$ be one of its $K_{3}$ subgraphs. Also, let $\ell_{1}$ and $\ell_{2}$ be any squared edge-length maps for $H$ such that some edge is assigned distinct positive real numbers by these maps. If the linkages $(H,\ell_{1})$ and $(H,\ell_{2})$ both have $2$-realizations, then there exist squared edge-length maps $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ for $G$ such that
  1. (i)
    they agree with each other on the edges not in $H$ and assign these edges positive real numbers,
  2. (ii)
    $\ell^{\prime}_{i}$ agrees with $\ell_{i}$ on the edges of $H$,
  3. (iii)
    the linkages $(G,\ell^{\prime}_{1})$ and $(G,\ell^{\prime}_{2})$ both have $2$-realizations.
Proof.
We can clearly choose $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ to satisfy Condition (ii). We show that we can choose these maps to additionally satisfy Conditions (i) and (iii). Let $\{v_{i}\}$ be the vertex set of $G$, let $H$ not contain $v_{4}$, and let $\ell_{1}(v_{1}v_{2})$ and $\ell_{1}(v_{1}v_{2})$ be distinct positive real numbers. Note that there exists a $2$-realization $p$ of $(H,\ell_{1})$ that maps $v_{1}$ to $a_{1}=(0,0)$, $v_{2}$ to $b_{1}=(\ell_{1}(v_{1}v_{2}),0)$, and $v_{3}$ to some point $c_{1}=(x,y)$ with $y\geq 0$. Similarly, there exists a $2$-realization $q$ of $(H,\ell_{2})$ that maps $v_{1}$ to $a_{1}=(0,0)$, $v_{2}$ to $b_{2}=(\ell_{2}(v_{1}v_{2}),0)$, and $v_{3}$ to some point $c_{2}=(w,z)$ with $z\geq 0$. Since $\ell_{1}(v_{1}v_{2})$ and $\ell_{1}(v_{1}v_{2})$ are distinct, there is a unique line $L(b_{1}b_{2})$ that contains $b_{1}$ and $b_{2}$. If $c_{1}=c_{2}$, then consider any point $d$ on the line $\bar{L}(b_{1}b_{2})$ that bisects the line-segment between $b_{1}$ and $b_{2}$. Set $\ell^{\prime}_{i}(v_{1}v_{4})=\|a_{i}-d\|^{2}$, $\ell^{\prime}_{i}(v_{2}v_{4})=\|b_{i}-d\|^{2}$, and $\ell^{\prime}_{i}(v_{3}v_{4})=\|c_{i}-d\|^{2}$. Clearly, $\ell^{\prime}_{1}$ and $\ell^{\prime}_{2}$ satisfy Conditions (i)-(iii), and so the lemma is proved.
Next, if $c_{1}\neq c_{2}$, then there is a unique line $L(c_{1}c_{2})$ that contains $c_{1}$ and $c_{2}$. If $\bar{L}(b_{1}b_{2})$ does not intersect the line $\bar{L}(c_{1}c_{2})$ that bisects the line-segment between $c_{1}$ and $c_{2}$, then a simple algebraic geometry argument shows that we can rotate $b_{1}$ and $c_{1}$ about $(0,0)$ by some angle $\theta$ to obtain the points $b^{\prime}_{1}$ and $c^{\prime}_{1}$, respectively, such that either $c^{\prime}_{1}=c_{2}$ or the line that bisects the line-segment between $b^{\prime}_{1}$ and $b_{2}$ intersects the line that bisects the line-segment between $c^{\prime}_{1}$ and $c_{2}$. In the former case, an argument similar to the one above applies. In the latter case, we can set $d$ to be the intersection of these two lines and then apply a similar argument. This completes the proof. ∎
Proof of Proposition 10.
If $G^{\prime}$ is one of the graphs in Figures 15(a) - 15(c), then it is easy to see that Lemma 24 applies and proves the proposition. Else, if $G^{\prime}$ is the graph in Figure 15(d), then we can first apply Lemma 24 to the subgraph of $G^{\prime}$ obtained by deleting the unlabeled vertices, and then add these vertices back and apply Lemma 55 to prove the proposition. Lastly, if $G^{\prime}$ is the graph in Figure 15(e), then consider the subgraph of $G^{\prime}$ obtained by deleting all but one unlabeled vertex, which we call $x$. We first apply Lemma 24 to this subgraph while treating $xw_{2}$ as $f$, then we apply this lemma again while treating $xw_{2}$ as $f^{\prime}$ and $uw_{1}$ as $f$. Finally, we add the deleted vertices and apply Lemma 55 to prove the proposition. ∎
Appendix D Proof of Theorem 5 [Byproduct: $3$-connected partial $3$-tree star theorem]
The proof of Theorem 5 requires Lemma 20, from Section B.5, and Lemmas 56 and 57, below, which may themselves be independently interesting results about partial $3$-trees.
Lemma 56 (If $G\cup\{wx,wy\}$ has a forbidden minor, then so does $G\cup wx$).
Let $G$ be a $3$-connected graph and $F=\{wx,wy\}$ be some subset of its nonedges. If there exists a forbidden minor $[G\cup F]$ in which $[w]$, $[x]$, and $[y]$ are distinct vertices and $wx$ and $wy$ are retained, then some $f\in F$ is such that $G\cup f$ has a forbidden minor.
Proof.
Let $G^{\prime}=G\cup F$ and assume that the forbidden minor $[G^{\prime}]$ exists. Since $G$ is $3$-connected, it contains at least three internally vertex-disjoint paths $P_{a}$, $P_{b}$, and $P_{c}$ between $w$ and $x$, where $a$, $b$, and $c$ are the first vertices along these paths not contained in $[w]^{-1}$, respectively. Note that $[x]^{-1}\cup[y]^{-1}$ does not contain $a$, $b$, or $c$. Also, since $[w]$ has degree two in $[G^{\prime}]\setminus[F]$, we can choose these paths such that $[a]=[b]$. For any minor $[G^{\prime}]_{\star}$ in which $[w]_{\star}$, $[x]_{\star}$, and $[y]_{\star}$ are distinct vertices and $wx$ and $wy$ are retained, define $a([G^{\prime}]_{\star})$ to be the length of the longest path in $G^{\prime}$ that contains $a$ and consists only of vertices in $V(P_{a})\cap[a]_{\star}^{-1}$. Define $b([G^{\prime}]_{\star})$ similarly. We will show how to obtain a minor $[G^{\prime}]_{A}$ from $[G^{\prime}]$ via a sequence of contractions and component exchanges such that (i) $[G^{\prime}]_{A}$ is a forbidden minor and $[w]_{A}$, $[x]_{A}$, and $[y]_{A}$ are distinct vertices and one of the following statements is true: (ii) either $wx$ or $wy$ is doubled in $[G^{\prime}]_{A}$ or (iii) $a([G^{\prime}]_{A})\geq a([G^{\prime}])$, $b([G^{\prime}]_{A})\geq b([G^{\prime}])$, and at least one of these inequalities is strict. If Statements (i) and (ii) are true, then the lemma is clearly proved. Otherwise, if Statements (i) and (iii) are true, then, since $P_{a}$ and $P_{b}$ are each finite and end with $x$, we can repeat this argument for $[G^{\prime}]_{A}$, and so on, until we obtain a forbidden minor of $G^{\prime}$ for which Statements (i) and (ii) are true.
We complete the proof by showing how to obtain $[G^{\prime}]_{A}$. Let $a^{\prime}$ be the first vertex after $a$ in $P_{a}$ that is not contained in $[a]^{-1}$. Define $b^{\prime}$ similarly. First, we show that we can wlog assume neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. If this is not the case, then assume that $a^{\prime}\in[w]^{-1}$. Let $a^{\prime\prime}$ be the first vertex in $P_{a}$ after $a^{\prime}$ that is not contained in $[w]^{-1}$, and define $S_{a^{\prime}}$ to be the path in $P_{a}$ between $a^{\prime}$ and the vertex that immediately precedes $a^{\prime\prime}$. Let $J$ be the subgraph of $G^{\prime}$ induced by $[w]^{-1}$. Consider the connected component $J_{w}$ of $J\setminus S_{a^{\prime}}$ that contains $w$, and let $J_{a^{\prime}}=J\setminus J_{w}$. Note that $J_{a^{\prime}}$ does not contain any vertex along $P_{b}$ between $w$ and $b$. If $a^{\prime\prime}$ is contained in $[a]^{-1}$ and $J_{w}$ neighbors $[d]^{-1}$, where $[d]$ is the neighbor of $[w]$ other than $[a]$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[w]^{-1}$ to $[a]^{-1}$ that exchanges $J_{a^{\prime}}$ satisfies Statements (i) and (iii). Otherwise, if either $a^{\prime\prime}$ is contained in $[a]^{-1}$ and $J_{w}$ does not neighbor $[d]^{-1}$ or $a^{\prime\prime}$ is contained in $[d]^{-1}$, then the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[w]^{-1}$ to $[d]^{-1}$ that exchanges $J_{a^{\prime}}$ satisfies (i) and $a([G^{\prime}]_{B})=a([G^{\prime}])$ and $b([G^{\prime}]_{B})=b([G^{\prime}])$. By repeating this argument for $[G^{\prime}]_{B}$ and $b^{\prime}$ if necessary, we obtain a minor with the desired property, which we can choose $[G^{\prime}]$ to be to begin with.
Next, let $H$ be the subgraph of $G^{\prime}$ induced by $[a]^{-1}$. Also, let $R_{a}$ be the path in $P_{a}$ between $a$ and the vertex that immediately precedes $a^{\prime}$. Define $R_{b}$ similarly. Consider the connected component $H_{b}$ of $H\setminus R_{a}$ that contains $b$, and let $H_{a}=H\setminus H_{b}$. Note that $H_{a}$ contains $R_{a}$ and $H_{b}$ contains $R_{b}$. Let $V_{a}$ be the subset of all vertices of $G^{\prime}\setminus([a]^{-1}\cup[w]^{-1})$ that the vertices in $H_{a}$ neighbor. Define $V_{b}$ similarly. Note that $a^{\prime}\in V_{a}$ and $b^{\prime}\in V_{b}$ since neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. There are two cases.
Case 1: $[G^{\prime}]$ is $K_{5}$.
If $V_{a}$ is a subset of either $[x]^{-1}$ or $[y]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Else, if $V_{a}$ is a subset of $[d]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (iii). Otherwise, if $V_{a}$ is a subset of $[d]^{-1}\cup[x]^{-1}$ but not a subset $[d]^{-1}$ or $[x]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[y]^{-1}$ that exchanges $H_{b}$ satisfies Statements (i) and (ii). A similar argument works if $V_{a}$ is a subset of $[d]^{-1}\cup[y]^{-1}$ but not a subset $[d]^{-1}$ or $[y]^{-1}$. If $V_{a}$ is a subset of $[x]^{-1}\cup[y]^{-1}$ but not a subset $[x]^{-1}$ or $[y]^{-1}$, then $V_{b}$ contains some vertex in $[d]^{-1}$. Furthermore, using arguments similar to those above, we can assume $V_{b}$ contains some vertex in $[x]^{-1}\cup[y]^{-1}$. Wlog, if this vertex is contained in $[y]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Lastly, the above arguments allow us to assume that $V_{a}$ and $V_{b}$ each contain vertices in $[d]^{-1}$, $[x]^{-1}$, and $[y]^{-1}$, in which case the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii).
Case 2: $[G^{\prime}]$ is $K_{2,2,2}$.
Wlog, assume that the neighborhood of $[a]$ is $\{[d],[e],[w],[x]\}$, where $[e]$ is the vertex in $V([G^{\prime}])\setminus\{[a],[d],[w],[x],[y]\}$. If $V_{a}$ is a subset of $[x]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (ii). Else, if $V_{a}$ is a subset of $[d]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$ satisfies Statements (i) and (iii). Otherwise, if $V_{a}$ is a subset of $[e]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[e]^{-1}$ that exchanges $H_{a}$. Observe that contracting $[bd]_{B}$ yields a $K_{5}$ minor $[G^{\prime}]_{C}$ in which $[w]_{C}$, $[x]_{C}$, and $[y]_{C}$ are distinct vertices and $wx$ and $wy$ are retained. Hence, the existence of $[G^{\prime}]_{A}$ follows from Case 1. Similar arguments apply for all remaining cases of $V_{a}$.
The above cases are exhaustive, and so the lemma is proved. ∎
Lemma 57 (If $G\cup\{wx,wy,wz\}$ has a forbidden minor, then so does $G\cup\{wx,wy\}$).
Let $G$ be a $3$-connected graph and $F=\{wx,wy,wz\}$ be some subset of its nonedges. If there exists a forbidden minor $[G\cup F]$ in which $[w]$, $[x]$, $[y]$, and $[z]$ are distinct vertices and $wx$, $wy$, and $wz$ are retained, then some size two subset $F^{\prime}$ of $F$ is such that $G\cup F^{\prime}$ has a forbidden minor.
Proof.
The proof is similar to the proof of Lemma 56. Let $G^{\prime}=G\cup F$ and assume that $[G^{\prime}]$ exists. Since $G$ is $3$-connected, it contains at least two internally vertex-disjoint paths $P_{a}$ and $P_{b}$ between $w$ and $x$, where $a$ and $b$ are the first vertices along these paths that are not contained in $[w]^{-1}$, respectively. Note that $[x]^{-1}\cup[y]^{-1}\cup[z]^{-1}$ does not contain $a$ or $b$. Also, since $[w]$ has degree one in $[G^{\prime}]\setminus[F]$, we have $[a]=[b]$. We will show how to obtain a minor $[G^{\prime}]_{A}$ from $[G^{\prime}]$ via a sequence of contractions and component exchanges such that $[G^{\prime}]_{A}$ is a forbidden minor in which $[w]_{A}$, $[x]_{A}$, $[y]_{A}$, and $[z]_{A}$ are distinct vertices and either $wx$, $wy$, or $wz$ is doubled. If such a minor exists, then the lemma is clearly proved.
We complete the proof by showing how to obtain $[G^{\prime}]_{A}$. Let $a^{\prime}$ be the first vertex after $a$ in $P_{a}$ that is not contained in $[a]^{-1}$. Define $b^{\prime}$ similarly. First, we show that we can wlog assume neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. If this is not the case, then assume that $a^{\prime}\in[w]^{-1}$. Since $[w]$ has degree one in $[G^{\prime}]\setminus[F]$, the first vertex $a^{\prime\prime}$ after $a^{\prime}$ in $P_{a}$ that is not contained in $[w]^{-1}$ is contained in $[a]^{-1}$. Define $S_{a^{\prime}}$ to be the path in $P_{a}$ between $a^{\prime}$ and the vertex that immediately precedes $a^{\prime\prime}$. Let $J$ be the subgraph of $G^{\prime}$ induced by $[w]^{-1}$. Consider the connected component $J_{w}$ of $J\setminus S_{a^{\prime}}$ that contains $w$, and let $J_{a^{\prime}}=J\setminus J_{w}$. Note that $J_{a^{\prime}}$ does not contain any vertex along $P_{b}$ between $w$ and $b$. The minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[w]^{-1}$ to $[a]^{-1}$ that exchanges $J_{a^{\prime}}$ is a forbidden minor in which $[w]_{B}$, $[x]_{B}$, $[y]_{B}$, and $[z]_{B}$ are distinct vertices and $wx$, $wy$, and $wz$ are retained. By repeating this argument for $[G^{\prime}]_{B}$, and so on, we eventually obtain a minor with the desired property, which we can choose $[G^{\prime}]$ to be to begin with.
Next, let $H$ be the subgraph of $G^{\prime}$ induced by $[a]^{-1}$. Also, let $R_{a}$ be the path in $P_{a}$ between $a$ and the vertex that immediately precedes $a^{\prime}$. Define $R_{b}$ similarly. Consider the connected component $H_{b}$ of $H\setminus R_{a}$ that contains $b$, and let $H_{a}=H\setminus H_{b}$. Note that $H_{a}$ contains $R_{a}$ and $H_{b}$ contains $R_{b}$. Let $V_{a}$ be the subset of all vertices of $G\setminus([a]^{-1}\cup[w]^{-1})$ that the vertices in $H_{a}$ neighbor. Define $V_{b}$ similarly. Note that $a^{\prime}\in V_{a}$ and $b^{\prime}\in V_{b}$ since neither $a^{\prime}$ nor $b^{\prime}$ is contained in $[w]^{-1}$. There are two cases.
Case 1: $[G^{\prime}]$ is $K_{5}$.
If $V_{a}$ is a subset of either $[x]^{-1}$, $[y]^{-1}$, or $[z]^{-1}$, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ has the desired properties. Else, if $V_{a}$ is a subset of the union of any two of these sets, wlog say $[x]^{-1}\cup[y]^{-1}$, but not a subset of any one of these sets, then the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[z]^{-1}$ that exchanges $H_{b}$ has the desired properties. Otherwise, consider the case where $V_{a}$ and $V_{b}$ each contain vertices in $[x]^{-1}$, $[y]^{-1}$, and $[z]^{-1}$, in which case the minor $[G^{\prime}]_{A}$ obtained via the component exchange from $[a]^{-1}$ to $[x]^{-1}$ that exchanges $H_{a}$ has the desired properties.
Case 2: $[G^{\prime}]$ is $K_{2,2,2}$.
Wlog, assume that the neighborhood of $[a]$ is $\{[d],[w],[x],[y]\}$, where $[d]$ is the vertex in $V([G^{\prime}])\setminus\{[a],[w],[x],[y],[z]\}$. If $V_{a}$ is a subset of either $[x]^{-1}$ or $[y]^{-1}$, then the minor obtained via the component exchange from $[a]^{-1}$ to $[V_{a}]^{-1}$ that exchanges $H_{a}$ has the desired properties. Else, if $V_{a}$ is a subset of $[d]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{a}$. Contracting $[bx]_{B}$ yields a minor with the desired properties. Otherwise, if $V_{a}$ is a subset of $[x]^{-1}\cup[y]^{-1}$ but not a subset of $[x]^{-1}$ or $[y]^{-1}$, then consider the minor $[G^{\prime}]_{B}$ obtained via the component exchange from $[a]^{-1}$ to $[d]^{-1}$ that exchanges $H_{b}$. Similar arguments apply for all remaining cases of $V_{a}$.
The above cases are exhaustive, and so the lemma is proved. ∎
Proof of Theorem 5.
Assume that $G\cup wu_{i}$ is a partial $3$-tree for each $wu_{i}$. If $G\cup\{wu_{i}\}$ is $3$-flattenaable, then the lemma follows from Lemma 20. Otherwise, let $F$ be a subset of $\{wu_{i}\}$ of minimum size such that there exists a forbidden minor $[G\cup F]$. Consider the case where some pair $f\in F$ is contracted in every forbidden minor of $G\cup F$, and let $F^{\prime}=F\setminus\{f\}$. Proposition 1 shows that $(G\cup F^{\prime},f)$ is a winged graph pair with an $f$-winged graph minor. Therefore, either Proposition 3, if some such minor has singleton wing tips, or Lemma 12, otherwise, shows that $G\cup F^{\prime}$ has a separator of size two. However, this contradicts the fact that $G$ is $3$-connected.
Next, we can assume that some pair in $F$ is preserved in $[G\cup F]$. Let $X$ be the set of all such pairs and let $F^{\prime}=F\setminus X$. The minimality condition on $F$ implies that each pair in $X$ is retained in $[G\cup F]$. Hence, since $G$ is connected and each vertex of any forbidden minor has degree four, we have $|X|\leq 3$. If $|X|>1$, then applying either Lemma 56 or Lemma 57 followed by Lemma 56 to $G\cup F^{\prime}$ and $X$ shows that the minimality condition on $F$ is violated. Therefore, assume that $X$ contains exactly one pair, say $f=[wz]$. Let $H$ be the subgraph of $G$ induced by $[w]^{-1}$ and let $H_{w}$ be its connected component that contains $w$. Also, let $\{[v_{i}]\}$ be the neighborhood of $[w]$ in $[G\cup F]\setminus[z]$. If each set $[v_{i}]^{-1}$ contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$, then $G\cup f$ clearly has a forbidden minor, which contradicts our assumption that $G\cup f$ is $3$-flattenable. Otherwise, if exactly two sets in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$ and $[v_{2}]^{-1}$, each contain some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$, then some connected component $J$ of $H\setminus H_{w}$ contains both some vertex in $\{u_{i}\}$ and some vertex whose neighborhood in $G$ contains some vertex in $[v_{3}]^{-1}$. The minor $[G\cup F]_{A}$ obtained via the component exchange from $[w]^{-1}$ to $[v_{3}]^{-1}$ that exchanges $J$ either is or has a forbidden minor in which at least two pairs in $\{wu_{i}\}$ are preserved, which is a contradiction.
Lastly, assume that exactly one set in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$, contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$. Let $\{J_{i}\}$ be the set of connected components of $H\setminus H_{w}$. Observe that each $J_{i}$ contains some vertex $u_{j}$ and some vertex whose neighborhood in $G$ contains some vertex in at least one of the sets $[v_{i}]^{-1}$ or $[z]^{-1}$. Furthermore, $[v_{2}]^{-1}$ and $[v_{3}]^{-1}$ each contain some vertex whose neighborhood in $G$ contains some vertex in some $J_{i}$. If the set $\{J_{i}\}$ contains more than one component, then wlog let $J_{1}$ contain some vertex whose neighborhood in $G$ contains some vertex in $[v_{2}]^{-1}$. Also, let $x$ be any vertex with $[x]\neq[w]$ and whose neighborhood in $G$ contains some vertex in $J_{2}$. Let $[x]=[v_{3}]$ if possible. We arrive at the same contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to $[x]^{-1}$ that exchanges $J_{2}$. On the other hand, if $J_{1}$ is the only component in $\{J_{i}\}$, then let $y$ be any vertex in $J_{1}$ whose neighborhood in $G$ contains some vertex in $[v_{2}]^{-1}$. Since $G$ is $3$-connected, there exists some vertex $x$ with $[x]\neq[w]$ and whose neighborhood in $G$ contains a vertex in some connected component $K$ of $J_{1}\setminus y$. Let $[x]=[v_{3}]$ if possible. We arrive at the same contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to $[x]^{-1}$ that exchanges $K$. This completes the proof. ∎
Appendix E Results about atoms
Here we collect various results about atoms.
Lemma 58 (Induced minors of $G$ that are atoms extend induced minors of atoms of $G$).
If an induced minor $[G]$ of a graph $G$ is an atom, then $[G]$ is an extension of some induced minor of some atom of $G$.
Proof.
Assume that $[G]$ is an atom. We proceed by induction on the number $n$ of cms’s of $G$. When $n=0$, the lemma is immediate. Assume the lemma is true when $n\leq k$, for any integer $k\geq 0$, and we will prove it when $n=k+1$. Let $C$ be any cms of $G$. Since $[G]$ is an induced minor and an atom, at most one $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$. Let $H$ be any $C$-component of $G$ that contains $C$, which exists since $C$ is a minimal separator. If no $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$, then each vertex of $G\setminus C$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Otherwise, we can wlog choose $C$ and $H$ such that $H$ additionally contains some vertex that is not contained in $[C]^{-1}$. In this case, each vertex of $G\setminus H$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Therefore, in either case, $[G]$ is an extension of some induced minor $[H]_{A}$ that is an atom.
Next, observe that $H$ has strictly fewer cms’s than $G$. Therefore, by the inductive hypothesis, $[H]_{A}$ is an extension of some induced minor $M$ of some atom of $H$. Finally, the lemma follows from the observation that $[G]$ is an extension of $M$. ∎
Lemma 59 (Conditions under which nonedge $f$ is contained in each atom of $G\cup f$).
For any nonedge $f$ of a graph $G$ whose atom graph is a path, each cms, and consequently each atom, of $G\cup f$ contains $f$ if either of the following statements are true:
  1. (1)
    $G$ is an atom.
  2. (2)
    The atom graph of $G$ contains two distinct leaves $y_{1}$ and $y_{2}$, whose not necessarily distinct neighbors are $x_{1}$ and $x_{2}$, respectively, and the endpoints of $f$ are contained in the graphs $G(y_{1})\setminus G(x_{1})$ and $G(y_{2})\setminus G(x_{2})$.
Proof.
If Statement (1) is true, then the lemma is easy to verify. Consider the case where Statement (2) is true. If $G\cup f$ is an atom, then the lemma is proved. Otherwise, assume to the contrary that some cms $E$ of $G\cup f$ does not contain $f$, then $E$ is clearly a cms of $G$. Furthermore, by assumption, the endpoints of $f$ are contained in distinct $E$-components of $G$. Since the atom of graph of $G$ is a path, there are exactly two $E$-components of $G$. Hence, these facts show that there exists some path in $G\cup f$ between any two vertices in $(G\cup f)\setminus E$. This contradicts the fact that $E$ is a cms of $G\cup f$, and so $E$ must contain $f$. Thus, each cms of $G$ contains $f$, and so the proof is complete. ∎
Lemma 60 (How a cms partitions atoms and cms’s).
For any cms $C$ of a graph $G$, the set of all atoms of $G$ is equal to the disjoint union of the sets of all atoms of all $C$-components of $G$. Also, if no proper subgraph of any cms of $G$ is a cms of $G$, then set of all cms’s of $G$ is equal to the disjoint union of the set $\{C\}$ and the sets of all cms’s of all $C$-components of $G$.
Proof.
The first claim follows from the observations that any atom of $G$ is contained in a unique $C$-component of $G$ and the set of all atoms of $G$ is equal to the union of the sets of all atoms of all $C$-components of $G$. Next, assume that no proper subgraph of any cms of $G$ is a cms of $G$. First, note that $C$ is not a cms of any $C$-component of $G$. Second, any cms $E$ of a $C$-component of $G$ is clearly a cms of $G$. Also, since $E$ is not a subgraph of $C$, it is contained in a unique $C$-component of $G$. Third, by a similar argument, any cms $E$ of $G$ other than $C$ is contained in a unique $C$-component $H$ of $G$. Hence, it is easy to see that $E$ is a cms of $H$. Combining these facts completes the proof. ∎
Lemma 61 (How a cms partitions the vertices of a separator).
Let $S$ be a separator of a graph $G$ and $H$ be the union of $S$ and any subset connected components of $G\setminus S$. If $G$ is an atom, then, for any cms $E$ of $H$, each connected component of $H\setminus E$ contains at least one vertex of $S\setminus E$.
Proof.
If some cms $E$ of $H$ is such that some connected component of $H\setminus E$ does not contain any vertex of $S\setminus E$, then clearly $E$ is a cs of $G$, and thus $G$ is not as atom. ∎
Lemma 62 (Condition under which an atom graph is a tree).
[…]
Proof.
We proceed by induction on the number $n$ of cms’s of $G$. When $n=0$, the lemma is immediate, and so the base case is true. For any integers $k\geq 0$ and $0\leq j\leq k$, assume the lemma is true when $n=j$, and we will prove it when $n=k+1$. Since $k+1>0$, $G$ has a cms $C$. Consider the set $\{H_{i}\}$ of $C$-components of $G$. Lemma 60 shows that each graph $H_{i}$ has strictly fewer cms’s than $G$, and any cms $E$ of $H_{i}$ is a cms of $G$. Hence, no proper subgraph of $E$ is a cms of $H_{i}$. Therefore, the inductive hypothesis applies to each atom graph $\mathcal{G}(H_{i})$.
[…]
For the first claim, since $C$ is a cms of $G$ but no proper subgraph of $C$ is a cms of $G$, $C$ is contained in some atom of each $C$-component $H_{i}$. Note that the intersection of any two atoms of $G$ is either empty or a cms of $G$. Hence, if there exist distinct atoms of $H_{i}$ that both contain $C$, then $C$ must be a proper subgraph of some of $G$, which is a contradiction. Therefore, $C$ is contained in a unique atom $J_{i}$ of $H_{i}$.
[…]
Lemma 63 (Condition under which an atom graph is a path).
Let $G$ be a graph that is an atom and let $S$ be one of its subgraphs and separators such that, for any cms $C$ of $S$, there are exactly two $C$-components of $S$, both of which are atoms. If $H$ is the union of $S$ and any subset of connected components of $G\setminus S$, then the atom graph of $H$ is a path.
Proof.
If $H$ is an atom, then the lemma is proved. Otherwise, for any cms $E$ of $H$, we can use Lemma 61 and our assumption on $S$ to see that there are exactly two $E$-components of $H$. This implies that no proper subgraph of $E$ is a cms of $H$. Since $E$ was arbitrary, this shows that Lemma 62 applies. Hence, each vertex in $X(H)$ has degree $2$, and it suffices to show that each vertex in $Y(H)$ has degree $2$. If this is not the case, then there exist three leaf nodes $y_{1}$, $y_{2}$, and $y_{3}$ in $\mathcal{G}(H)$ whose neighbors are $x_{1}$, $x_{2}$, and $x_{3}$, respectively. We will show that this implies the existence of a cms $C$ of $S$ such that some $C$-component of $S$ is not as atom, which contradicts our assumption.
By the discussion above, $H(y_{1})$ is a $H(x_{1})$-component of $H$. Hence, by Lemma 61, the connected component $H(y_{1})\setminus H(x_{1})$ contains a vertex $z_{1}$ of $S$. Similarly, $H(y_{2})\setminus H(x_{2})$ and $H(y_{3})\setminus H(x_{3})$ contain the vertices $z_{2}$ and $z_{3}$ of $S$, respectively. Consider any cms $E$ of $H$ other than $H(x_{1})$. If $H(y_{1})\setminus H(x_{1})$ contains some vertex in $E$, then $H(y_{1})$ contains $E$. However, this implies that $y_{1}$ has degree $2$, which contradicts the fact that $y_{1}$ is a leaf node. Hence, $H(y_{1})\setminus H(x_{1})$ does not contain any vertex in $E$. Similarly, $H(y_{2})\setminus H(x_{2})$ and $H(y_{3})\setminus H(x_{3})$ do not contain any vertex in $E$. Since the intersection of any two atoms of $H$ is either empty or a cms of $H$, this implies that $z_{1}$, $z_{2}$, $z_{3}$ are distinct vertices. However, these facts imply that $H(x_{1})\cap S$ is a cms of $S$ such that some $H(x_{1})\cap S$-component of $S$ is not as atom, which is a contradiction. Thus, setting $C=H(x_{1})$ completes the proof. ∎
Lemma 64 (Gluing atoms to obtain an atom 1).
Let $G$ be a graph with with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $S=G_{1}\cap G_{2}$ contains some nonedge $e$, and $G\setminus S$ is disconnected. If $G_{1}$ is an atom and either $G_{2}$ or $G_{2}\cup e$ is an atom, then $G$ is an atom.
Proof.
Assume that $G_{1}$ and either $G_{2}$ or $G_{2}\cup e$ is an atom, and consider any clique subgraph $E$ of $G$. Since $S$ contains a nonedge, the graph $S\setminus E$ is non-empty. Also, since $G_{1}$ is an atom, $G_{1}\setminus E$ is connected. Lastly, since either $G_{2}$ or $G_{2}\cup e$ is an atom, it is easy to see that, for any vertex $x$ of $G_{2}\setminus E$, there is some path in $G_{2}\setminus E$ between $x$ and some vertex in $S\setminus E$. Combining the above facts clearly shows that $G\setminus E$ is connected. Thus, since $E$ was an arbitrary clique subgraph of $G$, the lemma is proved. ∎
Lemma 65 (Gluing atoms to obtain an atom 2).
Let $G$ be a graph with with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $S=G_{1}\cap G_{2}$ contains some nonedge $e$, and $G\setminus S$ is disconnected. Then, $G$ is an atom if the following statements are true:
  1. (1)
    $G_{2}\cup e$ is an atom,
  2. (2)
    either $G_{1}$ or $G_{1}\cup e$ is an atom,
  3. (3)
    $G_{1}\setminus A$ is connected for any clique subgraph $A$ of $S$, and
  4. (4)
    $A$ does not separate the endpoints of $e$ in $G_{2}$.
Proof.
Assume that Statements (1)-(4) are true. If $G_{1}$ is an atom, then the lemma follows from Lemma 64. Otherwise, we must show that $G\setminus E$ is connected for any clique subgraph $E$ of $G$. First, we show that $G_{1}\setminus E$ is connected if $E$ is contained in $G_{2}$, and $G_{2}\setminus E$ is connected otherwise. Note that $A=E\cap S$ is either empty or a clique. If $E$ is contained in $G_{2}$, then $G_{1}\setminus E=G_{1}\setminus A$, which is connected by Statement (3). Otherwise, $E$ must be contained in $G_{1}$ since $S$ is a separator of $G$. Hence, we have $G_{2}\setminus E=G_{2}\setminus A$. Statement (1) shows that $(G_{2}\cup e)\setminus A$ is connected. Therefore, if $G_{2}\setminus A$ is disconnected, then $A$ must separate the endpoints of $e$ in $G_{2}$. Combining this with Statement (4) shows that $G_{2}\setminus E$ is connected.
Next, assume that $G_{1}\setminus E$ is connected. Since $S$ contains the nonedge $e$, $S\setminus E$ is non-empty. Using Statement (1), it is easy to see that, for any vertex $x\in G_{2}\setminus E$, there exists a path in $G_{2}\setminus E$ between $x$ and some vertex in $S\setminus E$. Hence, since $G_{1}\setminus E$ contains $S\setminus E$, this implies that $G\setminus E$ is connected. A similar argument using Statement (2) instead of (1) shows that $G\setminus E$ is connected if $G_{2}\setminus E$ is connected. This completes the proof. ∎
Lemma 66 (Paths between vertices of atoms).
Given an atom $J$ of a graph $G$, if there exists a path in $G$ between two vertices in $J$, then some subsequence of the vertices in this path is a path in $J$ between the same two vertices.
Proof.
Let $x$ and $y$ be any two vertices of $J$ such that there exists a path $P$ in $G$ between them. If $P$ is contained in $J$, then we are done. Otherwise, moving along $P$ from $x$ to $y$, let $b$ be the first vertex not contained in $J$ and let $a$ be the vertex that immediately precedes $b$. Then, there must exist a cms of $G$ containing $a$ and some vertex $c$ after $b$ in $P$. Hence, $ac$ is an edge of $G$. Note that replacing the path in $P$ between $a$ and $c$ with the edge $ac$ yields a shorter path $P^{\prime}$ between $x$ and $y$, which is a subsequence of $P$. Repeating this argument for $P^{\prime}$, and so on, eventually yields a path with the desired properties. ∎